### v1

# StefaLand: An Efficient Geoscience Foundation Model That Improves Dynamic Land–Surface Predictions

StefaLand: An Efficient Geoscience Foundation Model That Improves Dynamic Land–Surface Predictions
Nicholas Kraabel
Department of Civil and Environmental Engineering
Pennsylvania State University
&Jiangtao Liu11footnotemark: 1
Department of Civil and Environmental Engineering
Pennsylvania State University
&Yuchen Bian
Amazon Web Services
&Daniel Kifer
Department of Computer Science & Engineering
Pennsylvania State University
&Chaopeng Shen
Department of Civil and Environmental Engineering
Pennsylvania State University
Equal contribution.Corresponding author. cshen@engr.psu.edu
Stewarding natural resources, mitigating floods, droughts, wildfires, and landslides, and meeting growing demands require models that can predict climate-driven land-surface responses and human feedback with high accuracy. Traditional impact models, whether process-based, statistical, or machine learning, struggle with spatial generalization due to limited observations and concept drift. Recently proposed vision foundation models trained on satellite imagery demand massive compute and are ill-suited for dynamic land-surface prediction. We introduce StefaLand, a generative spatiotemporal earth foundation model centered on landscape interactions. StefaLand improves predictions on three tasks and four datasets: streamflow, soil moisture, and soil composition, compared to prior state-of-the-art. Results highlight its ability to generalize across diverse, data-scarce regions and support broad land-surface applications. The model builds on a masked autoencoder backbone that learns deep joint representations of landscape attributes, with a location-aware architecture fusing static and time-series inputs, attribute-based representations that drastically reduce compute, and residual fine-tuning adapters that enhance transfer. While inspired by prior methods, their alignment with geoscience and integration in one model enables robust performance on dynamic land–surface tasks. StefaLand can be pretrained and finetuned on academic compute yet outperforms state-of-the-art baselines and even fine-tuned vision foundation models. To our knowledge, this is the first geoscience land-surface foundation model that demonstrably improves dynamic land-surface interaction predictions and supports diverse downstream applications.
1 Introduction
Climate change is ushering in strong and widespread changes on the land surface, including higher frequencies of floods, droughts, wildfires and other geohazards Ebi et al. (2021), IPCC (2021). To mitigate the impact of these disasters, there are urgent needs for models that can accurately predict land surface dynamics such as streamflow, soil moisture, soil composition, landslides, snow water equivalent, groundwater levels, and vegetation carbon content. Among these, soil moisture controls the partition of rainfall into infiltration and runoff, modulates flood generation and landslides, and critically influences land-atmosphere interactions Dorigo et al. (2013). Streamflow is the flow rate of water running in the rivers, the most accessible water resource to humans, and too high or too low streamflow can cause flooding or hydrologic drought, respectively. Soil composition (sand, silt, clay fractions) governs infiltration capacity and root-zone storage, while slope–soil-vegetation interactions directly influence landslide hazards. Here, we limit our scope to the predictions of dynamical or static land surface processes that represent the impacts of climate change. Predicting these dynamic variables is distinct from image-recognition tasks, as here we seek to predict what will happen in the near or distant future.
Traditionally, these tasks were undertaken by physically-based models that take atmospheric forcings (precipitation, temperature) as inputs and sequentially calculate the physical processes that eventually lead to the variables of interest Li et al. (2015). In recent years, there has been a proliferation of data-driven machine learning (ML) models Solomatine and Ostfeld (2008). These models are often set up to accept forcing (dynamic weather) and landscape characteristics (static) data as inputs, and are trained to directly predict the natural land surface variables given the weather inputs. However, up to now, most of the geoscientific ML models have been supervised ML approaches trained specifically for a narrow set of tasks. Foundation models, which are trained on broad tasks to grasp the joint data distribution, have just started to emerge in geosciences, e.g., Terramind, Prithvi and Aurora. However, they were built mainly on satellite images for landcover-identification tasks Jakubik et al. (2023; 2025), Schmude et al. (2024). There has been a notable absence of foundation models built for dynamical land surface modeling, and a lack of effort to leverage valuable temporal datasets and ground-based observations Xie et al. (2023).
A Grand challenge for geoscientific ML models is to improve their spatial generalization, because a frequent issue facing them is the sparsity and spatial imbalance of observational data. Due to the cost of installing instruments and varying policies on data sharing, data are often only available in high density in certain regions of some developed nations. For example, streamflow gauge data are available at high density in United States, Europe, Australia and Japan, but are scanty in Africa, South America, and the rest of the Asia Global Runoff Data Centre (2020). Similar distribution patterns are found for high-quality in-situ soil moisture probes and soil property measurements. Satellite data have too coarse resolution and too large uncertainty levels compared to in-situ measurements. As quantified in many studies Feng et al. (2023), a deep network trained using data from some regions can face substantial performance degradation when applied in data-scarce regions. This is partly because there are not enough sites in space to learn the true dependencies of the targets on static land surface characteristics, and partly because of systematic data discrepancies among different regions (concept drift). This means ML land surface predictions may be unusable in more than 70% of the land. This limitation is particularly acute in data-scarce regions that are often most vulnerable to climate impacts. Any method that can reliably improve spatial generalization can be a significant and rare advance to greatly democratizes information for stakeholders worldwide.
Related Work: In hydrologic and ecosystem predictions, long short-term memory (LSTM) networks Hochreiter and Schmidhuber (1997) remain the dominant architecture, in part because land surface processes often behave like Markov processes where LSTMs’ gating mechanisms handle noisy continuous inputs well. Attempts to adopt transformers, so successful in natural language processing, have generally underperformed in hydrologic regression tasks Xue et al. (2023), Liu et al. (2024), with evidence of overfitting on continuous signals Zeng et al. (2022). Nonetheless, recent studies show that with task-specific modifications and careful fine-tuning, transformers can achieve competitive results and even offer advantages in autoregressive forecasting and extreme event prediction Wen et al. (2023), precisely the areas where current hydrologic models struggle most with spatial generalization.
Traditional hydrologic research on "prediction in ungauged basins" (PUB) have examined regionalization and spatial interpolation approaches including clustering or classifying catchments and transferring parameters from donor catchments in the same class Hrachowitz et al. (2013), Yang et al. (2023). Such an expert-derived design represents a crude practice of unsupervised learning that indicate the importance of understanding the joint data distribution. However, modern weakly-supervised foundation models can, in general, much better grasp the joint data distribution than expert-driven approaches. Foundation models offer a promising approach to address these spatial generalization challenges. By pretraining on large-scale datasets to learn generalizable representations, these models can potentially transfer knowledge across regions and geoscientific domains Zhang et al. (2024).
Existing geoscience foundation models such as TerraMind, Prithvi, Aurora, and AlphaEarth have largely focused on satellite imagery Hsu et al. (2024), which, while providing global coverage, may not capture the temporal dynamics and physical processes most relevant to land surface predictions. For example, TerraMind is trained on 9 million globally distributed satellite imagery samples across various modalities (optical, radar, elevation, land use) to create a generative multimodal foundation model for Earth observation Jakubik et al. (2023). It has been claimed to serve as "any-to-any generative, multimodal foundation model for Earth observation" Jakubik et al. (2025). However, many critical variables regarding soil and the subsurface are not directly observable from space. In addition, satellite images are noisy and a large portion of the data is redundant and irrelevant across repeated revisits. Therefore, it is unclear whether they are of high value to land-surface dynamical prediction tasks.
Our contributions: We present Spatiotemporal Earth Foundation model with Attributes for the Land Surface (StefaLand), the first land-focused geoscientific foundation model designed for dynamic land–surface prediction. StefaLand improves predictions on streamflow, soil moisture, and soil composition compared to state-of-the-art baselines. StefaLand further demonstrates strong spatial generalization across diverse landscapes and data-scarce regions. StefaLand’s attribute-based rather than image-based design (with the potential to link to image-like inputs in the future) is intentionally aligned with physical processes, drastically reducing compute requirements while retaining global coverage, making it accessible to academic researchers with modest resources. Pretraining requires only about 720 GPU hours, which makes it feasible within academic budgets. The model builds on a masked autoencoder backbone, a location-aware fusion of static and time-series inputs, grouped masking to promote cross-attribute interactions, and residual fine-tuning adapters, into a coherent design guided by geoscientific knowledge. Taken together, these contributions establish StefaLand as an efficient and accessible complement to vision-based foundation models, purpose-built for hydrologic and land–surface applications.
2 Methods
Dynamic land–surface prediction requires integrating heterogeneous information: landscape attributes such as topography, soils, and geology, together with dynamic meteorological forcings like precipitation and temperature. Capturing the joint influence of these variables is essential for spatial generalization but challenging with standard architectures. StefaLand addresses this gap through an attribute-based transformer framework that unifies static and time-varying signals in a shared representation space.
2.1 Model Architecture Overview
StefaLand processes both dynamic time series and static attributes through a unified encoder-decoder architecture, and is trained as a attribute-group-based masked autoencoder (Figure D1). The model embeds multivariate inputs, applies strategic masking to force the learning of cross-domain dependencies, and reconstructs the masked portions to generate deep representations. Our location-aware transformer encoder captures spatial relationships while positional encodings maintain temporal context. The complete mathematical formulation of our model architecture is provided in Appendix A. As well as a visual representation of the model is provide in appendix D.
2.2 Pretraining Setup
Our pretraining approach uses Cross-Variable Group Masking (CVGM), which forces the model to capture interactions among groups of landscape variables (e.g., terrain, climate, soil, vegetation) rather than treating them independently. This encourages the learning of physical relationships between different environmental drivers. The objective is a reconstruction loss on the masked slice of the sequence, normalized by variable-wise standard deviations when available and with per-variable weighting via a learnable weight vector. During pretraining, the dataset is stored in shards rather than fully loaded into memory. At the start of each epoch, we randomly reindex these shards, which effectively reshuffles the available samples. This lazy-loading strategy avoids materializing the entire corpus in memory while still ensuring high sample diversity across epochs.
The pretraining dataset is a derived global attribute dataset spanning 4,229 locations over 40 years. A complete list of variables and their sources is provided in Appendix C. The accompanying code release will enable reconstruction of this dataset. Hyperparameter settings and configurations for pretraining are also listed in the appendix above.
2.3 Transformer-based Direct Prediction Models
We developed several finetuned transformer models that combine StefaLand’s base structure with new units to directly predict streamflow and soil moisture. StefaLand with residual connection (resConn) is our main proposed architecture, integrating pretrained StefaLand transformers with an LSTM decoder through a residual pathway (Figure D1). This architecture enables iterative integration of pretrained transformer features throughout the decoding process, combining frozen transformer embeddings with forcings via a convolutional neural network before merging with the decoder through skip connections. This enables effective integration of pretrained features with task-specific requirements. Detailed adapter architectures are described in Appendix A.
For comparisons, "StefaLand without resConn" utilizes transformer outputs only once through a simple adapter (Figure D2), and a comparison with this model highlights the value of the resConn structure. "StefaLand Transformer" represents the base StefaLand model trained from scratch with the finetuning dataset. "StefaLand Ablation - resConn" utilizes the same structure as "StefaLand - resConn", but randomize the weights of StefaLand instead of using the pretrained weights. Comparing with these two models demonstrates the value of pretraining. As discussed below, we also implemented Terramind with resConn, which uses IBM’s Terramind foundational model together with our residual connection architecture, to contrast the effectiveness of Terramind vs StefaLand pretraining. Of course, supervised LSTM serves as a widely employed and highly-competitive baseline.
2.4 Physics-Based Integration
To leverage domain knowledge, we also implemented physics-based configurations that combine our data-driven approach with process understanding. For the process-based backbone, we employed the Hydrologiska Byråns Vattenbalansavdelning (HBV) model, a conceptual hydrologic model with state variables like snow storage, soil water, and subsurface storage. In this setup, the neural network only provides parameters for the process-based backbone, which then outputs a range of interpretable variables, and they are trained together in an end-to-end fashion.
Our physics-based approaches include LSTM-HBV1.1, a baseline configuration where LSTM outputs parameters for the HBV model; StefaLand-resConn HBV1.1; and StefaLand-no resConn HBV1.1, which integrates transformer features without residual connections. These physics-based approaches maintain interpretability while leveraging improved representations from StefaLand. Detailed implementations are provided in Appendix B.
2.5 Foundation model comparisons
For completeness, we also evaluated two existing Earth-Observation-oriented foundation models, TerraMind and PrithviWxC Jakubik et al. (2025), Hsu et al. (2024), using the same fine-tuning heads and training protocols as StefaLand. TerraMind was tested on streamflow and soil moisture, and PrithviWxC on soil moisture. Inputs were harmonized to each model’s modality. It is noteworthy that either Terramind or PrithviWxC requires orders of magnitude larger storage space for the input data than StefaLand, as well as computational cost for finetuning. Due to intensive data requirements for PrithviWxC on surface-level variables mostly likely to demonstrate relevance to land surface interactions, along with all 14 static variables were used. The full atmospheric variables at differing pressure levels were excluded.
3 Experiments
For all experiments described below, forcing data and static features remained consistent across all model configurations, with hyperparameters held constant within each test type except for fundamental architectural differences inherent to transformer models. Complete details of hyperparameters, forcings, and static features are provided in the Appendix C.
In addition to the neural and process-based baselines reported below, we trained plain linear regression models for every task using the same features (and appropriate standardization). These consistently underperformed all learning baselines by large margins, so we omit them from tables for readability; full metrics are provided in Appendix E.
3.1 CAMELS Streamflow Prediction
To compare the scheme’s spatial generalization capability on a well benchmarked dataset, we follow the experimental setup in (Feng et al., 2021), testing prediction in ungauged basins (PUB) and ungauged regions (PUR). These correspond to randomized spatial K-fold and regional-specific K-fold regimes, reflecting real-world challenges in generalizing to nearby unmonitored basins and to data-scarce regions, respectively. We use the CAMELS dataset (Addor et al., 2017, Newman et al., 2014), restricted to the 531-basin subset with clear watershed boundaries (Newman et al., 2017). Basins were divided into 10 random groups for PUB and 7 contiguous regions for PUR, employing leave-one-out in both cases. To avoid leakage, all CAMELS-overlapping stations were removed during pretraining for PUB tests, and entire regions were excluded for PUR tests. We evaluated both direct prediction models and neural networks used to parameterize process-based models (see Methods). Direct models included StefaLand without finetuning (StefaLand-Transformer), pretrained variants with and without residual connections (resConn / noResConn), a from-scratch ablation trained only on task data, and a version using TerraMind as encoder. For process-based models, we tested LSTM HBV1.1, StefaLand-resConn HBV1.1, and StefaLand-noResConn HBV1.1.
Table 1: CAMELS Streamflow PUB and PUR Results
Model Random holdout (ungauged basins) Regional holdout (ungauged regions)
RMSE $\downarrow$ pbRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$ RMSE $\downarrow$ pbRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$
Direct Prediction Models
LSTM - SL 1.402 1.360 0.762 0.636 1.609 1.457 0.743 0.554
StefaLand - Transformer 1.882 1.849 0.538 0.395 1.982 1.949 0.230 0.201
StefaLand - resConn 1.111 1.068 0.869 0.717 1.344 1.334 0.801 0.635
StefaLand - no resConn 1.171 1.154 0.823 0.706 1.376 1.356 0.798 0.610
Terramind - resConn 1.332 1.301 0.777 0.637 1.420 1.398 0.763 0.551
StefaLand Ablation - resConn 1.355 1.332 0.801 0.661 1.516 1.378 0.771 0.560
Physics Based Models
LSTM - HBV1.1 1.325 1.298 0.857 0.672 1.561 1.521 0.746 0.578
StefaLand - resConn HBV1.1 1.234 1.216 0.863 0.714 1.345 1.332 0.842 0.643
StefaLand - no resConn HBV1.1 1.315 1.302 0.848 0.707 1.401 1.379 0.835 0.623
StefaLand Ablation - resConn HBV1.1 1.310 1.306 0.842 0.693 1.465 1.432 0.607 0.512
The original multi-basin LSTM studies on CAMELS defined the modern supervised learning benchmark (Kratzert et al., 2019), and subsequent large-scale comparisons confirmed that vanilla Transformers generally fail to outperform LSTMs on rainfall–runoff prediction (Liu et al., 2024, Table 1 therein). This is again reinforced in our results, where StefaLand-Transformer underperforms LSTM. The Nash-Sutcliffe model efficiency coefficient (NSE) values reported here for LSTM are highly similar to those reported in the domain literature Feng et al. (2021). Broader evaluations likewise continue to rank LSTM-family models among the best-performing approaches (Lees et al., 2021), and even Google’s global flood-forecasting system adopts an encoder–decoder LSTM backbone (Nearing et al., 2024). We therefore use LSTM-SL as the supervised SOTA comparator, complemented by a process-informed LSTM–HBV1.1 variant, and additionally test TerraMind to control for the role of EO-centric pretraining.
Against these strong baselines, the foundation model pretraining clearly provides a rare boost to generalization capability compared to supervised learning models (LSTM-SL and StefaLand-Transformer) across both PUB and PUR scenarios (Table 1). StefaLand-resConn’s RMSE is almost $20\%$ lower than that of supervised LSTM in the PUB and $17\%$ lower in the harder PUR test, while the correlation are noticebly higher. The value of pretraining is confirmed by StefaLand-Ablation-resConn, which appears to have only minor advantage compared to LSTM. The residual-connection architecture enables effective integration of transformer features with temporal dynamicsh, showing a modest benefit compared to StefaLand-no resConn, but both are clearly stronger than either LSTM or StefaLand-Transformer. Finally, Terramind-resConn has quite comparable metrics to LSTM. It appears the pretrained Terramind offers no generalization value.
The foundation model showed its versatility in supporting the parameterization of hybrid models and improving their generalization. We observe the same patterns — the pretrained StefaLand is helpful and resConn is a valuable construct. The physics-based experiments highlight that residual connections can combine pretrained representations with process-based constraints, producing more reliable spatial generalization than even the strongest LSTM configurations. We note that the hybrid model in fact showed a slightly better NSE than the direction prediction models for PUR for either LSTM-HBV1.1 or StefaLand-resConn-HBV1.1, because the physics-based equations serve as additional constraints to mitigate overfitting.
3.2 Global Streamflow
We designed a global-scale runoff prediction experiment to assess the robustness and generalization ability worldwide. We filtered global basin datasets based on data completeness and retained 3,434 basins for the global experiment. To manage computational costs efficiently, we employed random hold-out sampling combined with three-fold cross-validation. Additionally, we implemented a regionally hold-out continental scenario (RH-C), excluding all basins from North America, South America, and Europe, respectively, from the training set. The model trained from remaining continents was then evaluated on the excluded continent, simulating the realistic scenario of deploying the model to unseen geographical domains.
Table 2: Global streamflow prediction across 3,434 basins worldwide
Random holdout Regional holdout
(ungauged basins) (ungauged continents)
Model RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$ RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$
LSTM - SL 0.870 0.864 0.798 1.253 1.202 0.672
StefaLand - Transformer 0.749 0.751 0.843 1.075 1.048 0.697
Terramind + finetuning 1.156 1.111 0.580 1.234 1.158 0.670
Cross-continental transfer testing using three-fold validation. Detailed metric calculations in Appendix E.
Results indicate that StefaLand-Transformer outperformed both LSTM-SL and the fine-tuned vision-based Transformer model (TerraMind) across all evaluation metrics. Specifically, in the RH scenario, StefaLand achieved an RMSE of 0.749, approximately 14% lower than LSTM (0.87), with Corr improving to 0.843. Notably, StefaLand’s unbiased RMSE closely matched its RMSE, indicating errors were primarily random fluctuations rather than systematic bias. In contrast, TerraMind exhibited a higher RMSE of 1.156 and a low Corr of 0.580. We hypothesize that this is because the pretraining satellite image data for the vision foundation network TerraMind, did not have high relevance to hydrologic predictions.
Under the more challenging RH-C scenario, all models experienced increased errors, but StefaLand continued to demonstrate higher performance. Its RMSE was 1.075, approximately 14.1% lower than LSTM’s 1.253, while also maintaining the highest Corr (0.697). Moreover, StefaLand’s ubRMSE remained close to its RMSE, confirming its robust ability to correct regional-scale biases even under extreme out-of-domain conditions. Conversely, LSTM and TerraMind exhibited large gaps between ubRMSE and RMSE, highlighting their challenging in producing hydrologically-relevant features.
3.3 Global Soil Moisture
We evaluated finetuning StefaLand for soil moisture predictions following the experimental design from Liu et al. (2023a), using data from the International Soil Moisture Network (ISMN) Dorigo et al. (2011; 2013). Even though there is a globally covering satellite-based product for soil moisture, the data quality can hardly match that of in-situ moisture sensors. Thus the ability to generalize in-situ data will still be valuable. ISMN consists of 1,316 ground-based measurement stations. We randomly partitioned sites into five groups for spatial cross-validation in the random holdout test, training on four groups and testing on the fifth, rotating through all groups to obtain comprehensive spatial generalization performance. For the regional holdout test, we specifically evaluated model performance on Europe as an ungauged region, training on all other continents while excluding European sites entirely to assess cross-continental transferability, removing 129 sites for testing. We tested the following model configurations: StefaLand-Transformer (no finetuning, direct prediction), StefaLand with and without residual connections (resConn / noResConn), an ablation trained from scratch without pretraining, a baseline LSTM, and a version using IBM’s TerraMind encoder with resConn.
Table 3: Soil moisture prediction across 1,316 ISMN stations
Random location holdout Regional holdout
(random sites) (Europe)
Model RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$ RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$
LSTM - SL 0.073 0.055 0.764 0.112 0.053 0.510
StefaLand - Transformer 0.140 0.103 0.637 0.135 0.112 0.503
StefaLand - resConn 0.068 0.054 0.783 0.090 0.059 0.638
StefaLand - no resConn 0.075 0.057 0.741 0.095 0.058 0.545
Terramind - resConn 0.083 0.062 0.694 0.101 0.080 0.519
PrithviWcX - resConn 0.081 0.060 0.703 0.103 0.079 0.523
Ablation StefaLand - resConn 0.074 0.058 0.749 0.108 0.064 0.528
5-fold spatial validation and cross-continental validation on Europe (129 sites). Detailed metric calculations in Appendix E.
LSTM again serves here as the established state-of-the-art baseline. Comprehensive reviews have shown that LSTM variants consistently provide the most stable and effective performance among supervised learning methods for soil moisture prediction due to their temporal modeling strengths (Wang et al., 2024). On the ISMN-linked global dataset in particular, multitask LSTM models are widely adopted and competitive with newer architectures (Liu et al., 2023b). Given these findings, we benchmark against LSTM-SL as the SOTA/near-SOTA supervised model, and include transformer-only and vision-foundation encoders to disentangle temporal modeling from representation learning.
Against these strong baselines, the soil moisture experiments confirm the superiority of the StefaLand-resConn architecture (Table 3). StefaLand-resConn achieves the best performance, with RMSE of 0.068 and correlation of 0.783 for random holdout, and maintains strong performance even in cross-continental testing on Europe (RMSE = 0.090, Corr = 0.638). The direct StefaLand model performs poorly since transformers alone lack an effective mechanism for temporal dependencies, while the resConn architecture compensates for this limitation. Similar to the streamflow results, TerraMind and PrithviWcX both showed no performance benefit compared to LSTM despite architectural adaptations, confirming both the challenge of repurposing vision foundation models across domains and the importance of pretraining with geoscience-relevant variables. Notably, the regional holdout on Europe demonstrates StefaLand-resConn’s superior spatial generalization capabilities, achieving a 25% correlation improvement over the LSTM baseline in this highly challenging extrapolation scenario.
We want to emphasize that TerraMind and PrithviWxC were not designed for the hydrologic and land–surface prediction tasks studied here. They excel in their intended domains of EO imagery and atmospheric variables, but are out-of-domain for dynamic land–surface and hydrologic modeling. We include them only to explore whether such models transfer any useful signal in our setting. Their weaker results should not be interpreted as evidence against their capability in other applications, but rather as a reflection of the mismatch between their pretraining objectives and the land–surface tasks we target. Nevertheless, we believe these benchmarks are helpful for clarifying their respective strengths, since the AI community may not be familiar with these datasets and models that have been claimed "any-to-any generative foundation model for Earth observation".
3.4 Soil Property Prediction
There are different soil datasets, each collected with different protocols and data processing techniques, resulting in significant discrepancies between datasets. In this test, we finetuned StefaLand to predict in-situ soil profile data from another dataset (ISRIC). This application can produce a seamless dataset that is consistent with a set of in-situ data, improving data availability and addressing systematic biases. In addition, it helps us understand the noise associate with each dataset. StefaLand’s pretraining soils dataset is HWSD, which has some overlap but also quite extensive differences from ISRIC, which is larger and potentially noisier. We finetuned StefaLand to predict one soil texture property (e.g., clay percentage) in ISRIC while masking the corresponding complementary attributes (e.g., sand, silt) from the same profile to avoid information leakage. As a result, we are probing how easy it is to infer soil properties using other attributes such as climate, terrain and land cover. We compared StefaLand with and without pretraining against a supervised random forest baseline, which is prevalently used in the present geoscientific literature.
Table 4: In-situ soil property prediction using ISRIC WoSIS data.
Model WoSIS Property Corr $\uparrow$ $R^{2}$ $\uparrow$
StefaLand - Transformer Clay 0.20 0.04
StefaLand - Finetune Clay 0.51 0.26
Random Forest Clay 0.05 0.00
StefaLand - Transformer Sand 0.25 0.06
StefaLand - Finetune Sand 0.70 0.50
Random Forest Sand 0.26 0.07
StefaLand - Transformer Silt 0.41 0.17
StefaLand - Finetune Silt 0.63 0.40
Random Forest Silt 0.84 0.70
StefaLand with finetuning achieved markedly higher predictive power, especially for sand fraction ($R^{2}=0.50$), compared to both direct training and the random forest baseline. These results highlight StefaLand’s ability to reconcile disparate soil datasets and improve the utility of noisy in-situ observations. At the same time, the weaker performance on silt suggests limits of transfer: silt fractions are often residually defined (sand and clay measured directly, silt inferred as the remainder), and thus more sensitive to error propagation across datasets. This makes them inherently harder to learn, as any noise or systematic difference between HWSD and ISRIC is amplified. The fact that the random forest attains higher accuracy on silt may also indicate that simpler models can more directly exploit the compositional closure relationship, whereas StefaLand relies more on cross-variable and environmental cues, which provide less direct signal for silt.
4 Discussion
4.1 Key Findings and Contributions
As we review the literature, the methods to improve spatial generalization to data‐scarce regions are rare and rarely effective Beery et al. (2018), Gacu et al. (2025). In contrast, StefaLand, combined with lightweight fine‐tuning heads, achieves state-of-the-art or competitive performance across four broad problem classes: streamflow (both CAMELS and global), soil moisture, soil composition while also strengthening the parameterization of differentiable process-based models. Across tasks, the strongest gains come from architectures that fuse StefaLand embeddings with explicit temporal modeling via residual connections, indicating that pretraining on attribute-based spatiotemporal structure yields problem-relevant representations while temporal heads resolve sequence dynamics. These outcomes support the premise that foundation models can democratize prediction quality in data-scarce regions by improving out-of-domain transfer.
The streamflow experiments, spanning basin-scale CAMELS PUB/PUR and global cross-continental runoff, consistently show StefaLand’s improved spatial generalization. The soil moisture results confirm robustness in both random site holdouts and continental transfer (Europe). For soil composition, StefaLand reconciles noisy WoSIS in-situ measurements with improved predictive power over random forests, highlighting its ability to harmonize disparate static datasets.
An additional contribution is computational: our attribute-based approach is markedly more efficient than pixel-wise satellite transformers. The underlying transformer has far fewer parameters (roughly 12 million) and avoids the heavy data management requirements of image-centric pretraining. TerraMind’s larger configuration corresponds to about 7,680 GPU hours Jakubik et al. (2025), Aurora required roughly 14,592 GPU hours Bodnar et al. (2025), and AlphaEarth consumed on the order of 28,672 GPU hours Brown et al. (2025). These figures do not include the petabyte-scale data storage and transfer demands for satellite archives. By comparison, StefaLand’s attribute-based pretraining requires only about 720 GPU hours, making high-quality spatial generalization feasible on budgets accessible to academic groups.
Finally, StefaLand is useful both operationally and as a representation layer. For hydrology, it supports point-scale prediction and regional products, for soil and geomorphic tasks it supplies frozen features that enhance conventional models. This dual role helps bridge a persistent gap between data-rich and data-poor regions by providing transferable geoscience embeddings that downstream models can reliably exploit.
4.2 Limitations
While StefaLand demonstrates strong generalization across diverse tasks, several limitations remain. The pretraining dataset is intentionally narrow and high quality, covering a carefully curated set of forcings and attributes. This focus improves signal extraction but may limit the model’s usefulness for problems outside or only loosely related to those variables, such as deeper subsurface processes like groundwater dynamics. Validation is also biased toward regions with dense observational coverage, such as the United States and Europe; although regional holdout tests suggest good transfer, true performance in data-scarce regions of Africa, Asia, and South America is less certain.
4.3 Future Work
Several directions remain to further develop StefaLand. Expanding the range of targets to include variables such as evapotranspiration, snow water equivalent, and groundwater levels would broaden its applicability, while harmonized global inventories could support more comprehensive evaluation of soil and landslide predictions across differing lithologies and climates. Methodologically, advances such as sequence-aware pretraining, uncertainty-aware prediction heads, and tighter integration with process models offer promising avenues to improve calibration and interpretability while preserving efficiency. Establishing standardized train–test protocols for random and regional holdouts, released with accompanying data splits, will also be important for enabling consistent benchmarking of future geoscience foundation models.
Overall, StefaLand shows that attribute-centric pretraining combined with lightweight temporal or physics heads can deliver strong spatial generalization across geoscientific tasks while remaining computationally accessible. This points toward a practical path for high-quality predictions in regions where they are most needed but data are most limited.
5 Reproducibility Statement
The pretrained StefaLand model and all code for both pretraining and finetuning is released publicly at [https://anonymous.4open.science/r/StefaLand-9421/]. All datasets used in this work both pretraining dataset and all finetuneing benchmarks are fully public. A complete list of variables used for each task, along with their data sources, is provided in Appendix C. We also report all hyperparameters and model details in the same Appendix.
Appendix A Detailed Model Architecture
This appendix provides the complete mathematical formulation of the StefaLand model architecture.
A.1 Embedding Dynamic and Static Inputs
StefaLand independently embeds each dynamic and static variable into a latent space. Specifically, for each dynamic variable $c$ at each time step $t$, a two-step nonlinear embedding is applied individually:
$z_{t,c}=\text{GELU}(x_{t,c}W_{1,c}+b_{1,c})W_{2,c}+b_{2,c}$ (1)
where $W_{1,c}\in\mathbb{R}^{1\times 64}$ and $W_{2,c}\in\mathbb{R}^{64\times 256}$ are embedding parameters. After embedding all dynamic variables individually, embeddings are stacked and summed across the variable dimension, resulting in a single embedding vector per time step:
$z_{t}=\sum_{c=1}^{C}z_{t,c}$ (2)
Similarly, static attributes are embedded individually:
$z_{\text{static},i}=\text{GELU}(s_{i}W_{1,i}+b_{1,i})W_{2,i}+b_{2,i}$ (3)
where separate embedding layers are used for static features. These individual static embeddings are then concatenated with dynamic embeddings along the temporal dimension, resulting in a unified embedding tensor:
$Z=[z_{1};z_{2};\ldots;z_{T};z_{\text{static}}]$ (4)
This static embedding acts as a global learnable token, allowing the model to incorporate basin-specific context into temporal dynamics at any depth of the Transformer layers.
A.2 Transformer Encoder
The embeddings enriched by positional encoding are processed through an $N$-layer Transformer encoder, where each Transformer block successively applies Multi-Head Self-Attention (MHA) with $h$ attention heads, followed by a residual connection and Layer Normalization. Subsequently, a position-wise Feedforward Network (FFN) is applied, also followed by another residual connection and Layer Normalization:
$A^{(\ell)}$ $=\text{MHA}(H^{(\ell-1)})$ (5)
$\tilde{H}^{(\ell)}$ $=\text{LayerNorm}(H^{(\ell-1)}+A^{(\ell)})$ (6)
[…]
$H^{(\ell)}$ $=\text{LayerNorm}(\tilde{H}^{(\ell)}+F^{(\ell)})$ (8)
A.3 Reconstruction of Original Inputs
The final hidden states from the Transformer encoder, $H^{(N)}$, are linearly projected and passed through a single-layer bidirectional LSTM to capture the local temporal dependencies and continuity:
$U=\text{LSTM}(H^{(N)}W_{\text{enc-proj}}+b_{\text{enc-proj}})$ (9)
The outputs $U$ are then separated into dynamic and static components, $U_{t}$ and $U_{\text{static}}$, corresponding to the temporal sequence and static attributes:
$U_{t},U_{\text{static}}=U_{1:T},U_{T+1}$ (10)
Finally, both dynamic and static representations are individually projected back to their original dimensions through separate embedding layers, reconstructing the masked portions of the inputs. Dynamic variables are restored via:
$\hat{x}_{t}=\text{DynamicDecEmbedding}(U_{t})$ (11)
while static attributes are restored by:
$\hat{s}=\text{StaticDecEmbedding}(U_{\text{static}})$ (12)
The projections leverage the learned latent representations to reconstruct the original hydrologic inputs.
Appendix B Physics-Based Differentiable Modeling
To leverage domain knowledge and physical constraints inherent in hydrological systems, we implemented physics-based models that explicitly represent hydrological processes through mathematical formulations. These differentiable versions can be trained end-to-end within neural network frameworks, combining process understanding with machine learning flexibility Shen et al. (2023).
For the process-based backbone, we employed the Hydrologiska Byråns Vattenbalansavdelning (HBV) model Aghakouchak and Habib (2010), Beck et al. (2020), Bergström (1976; 1992), Seibert and Vis (2012), a relatively simple bucket-type conceptual hydrologic model. HBV has state variables like snow storage, soil water, and subsurface storage, and can simulate flux variables like evapotranspiration (ET), recharge, surface runoff, shallow subsurface flow, and groundwater flow. The parameters of HBV are learned from basin characteristics by a deep learning network.
We used an updated modern version, HBV1.1 Song et al. (2025), which includes modifications such as increased parallel storage components to represent heterogeneity within basins and dynamic parameterization capabilities.
For physics-based configurations, we tested: (1) a baseline LSTM-HBV1.1 configuration as a standard reference, (2) StefaLand HBV1.1 with resConn, which combines the physics-based approach with our residual connection architecture, and (3) StefaLand HBV1.1 without resConn. These physics-based approaches incorporate hydrological process understanding while maintaining the ability to learn from data.
Appendix C Experimental Details
Table C1: StefaLand Pretraining Variables and Sources
Variable Type Variable Name Source
Time Series Forcings Precipitation, Short-wave solar radiation downwards, Relative humidity, Maximum temperature, Minimum temperature, Potential evapotranspiration Multi-Source Weather (MSWX) and Multi-Source Weighted-Ensemble Precipitation (MSWEP)
Static Attributes Forest cover fraction, grassland cover fraction Climate Change Initiative (CCI) land cover dataset
Normalized Difference Vegetation Index (NDVI) Terra Moderate Resolution Imaging Spectroradiometer (MODIS) Vegetation Indices (MOD13A3)
Sand, silt, clay fractions Harmonized World Soil Database (HWSD)
Elevation, slope, aspect Global Multi-resolution Terrain Elevation Data (GMTED)
Soil depth Global 1-km Gridded Thickness of Soil, Regolith, and Sedimentary Deposit Layers
Carbonate sedimentary rock fraction Global Lithological Map (GLiM)
Rock porosity, permeability GLobal HYdrogeology MaPS (GLHYMPS)
Population density Gridded Population of the World (GPW) v4 dataset
GDP per capita; population density Gross Domestic Product and Human Development Index over 1990-2015
Forest intact fraction Intact Forest Landscapes Data
Outputs None (self-supervised pretraining)
Table C2: CAMELS Streamflow HBV Model Hyperparameters
Parameter Value
General Settings
Random seed 111111
Data sampler finetune_sampler
Training Configuration
Time period 1989/10/01–2008/09/30
Optimizer Adadelta
Batch size 64
Epochs 25
Neural Model Configuration
Sequence length 365
Hidden size 512
Dropout 0.2
Encoder layers 4
Decoder layers 2
Feed-forward dimension 512
Physical Model (HBV-1.1)
Model type HBV_1_1p
Number of runs (nmul) 16
Warm-up period 365 days
Warm-up states True
Dynamic dropout 0.0
Use routing True
Dynamic parameters parBETA, parK0, parBETAET
Near-zero threshold 1e-05
Loss Function
Type RmseLoss
Table C3: CAMELS Streamflow Variables and Sources
Variable Type Variable Name Source
Time Series Forcings Precipitation, Temperature, Potential evapotranspiration, Solar radiation, Vapor pressure Catchment Attributes and Meteorology for Large-sample Studies (CAMELS)
Static Attributes Elevation, slope, catchment area, forest cover, LAI, GVF, soil depth, porosity, conductivity, sand, silt, clay fractions, carbonate fraction, permeability, aridity, snow fraction, precipitation extremes CAMELS
Outputs Streamflow CAMELS gauge records
Table C4: Soil Moisture Model Configuration
Parameter Value
General Settings
Mode train_test
Random seed 111111
Data loader onlylstm_loader
Data sampler finetuning_noHBV
Training Configuration
Time period 2015/04/01–2020/12/31
Target soil_moisture
Optimizer Adadelta
Batch size 128
Epochs 50
Save frequency Every 25 epochs
Neural Network Configuration
Hidden size 128
Dropout 0.3
Learning rate 1.2
Encoder layers 16
Decoder layers 12
Feed-forward dimension 512
Rho 365
Loss Function
Type RmseLoss
Table C5: Soil Moisture Variables and Sources
Variable Type Variable Name Source
Time Series Forcings Albedo (BSA, WSA) Moderate Resolution Imaging Spectroradiometer (MODIS) MCD43A3 Version 6
LST (Day, Night) MODIS Land Surface Temperature/Emissivity Daily (MYD11A1) Version 6.1
Precipitation Global Precipitation Measurement (GPM), Multi-Source WeightedEnsemble Precipitation (MSWEP)
Forecast albedo, LAI (high/low vegetation), soil temperature (layer 1), surface pressure, solar radiation, 2m temperature, evaporation, precipitation, U/V wind (10m) ERA5-Land Hourly - ECMWF Climate Reanalysis
Static Attributes elevation, slope, aspect, roughness, curvature Global 1,5,10,100-km Topography database
Sand, clay, silt, bulk density Harmonized world soil database (HWSD)
Land cover; urban; open water; snow/ice ESA CCI land cover 2018
NDVI Vegetation Indices Monthly L3 Global 0.05Deg CMG
Outputs Soil moisture ESA CCI SM / in-situ merged
Table C6: Global Streamflow Variables and Sources
Variable Type Variable Name Source
Time Series Forcings Precipitation, Temperature, Potential evapotranspiration, Radiation, Humidity MSWX and MSWEP
Static Attributes Forest cover fraction, grassland cover fraction Climate Change Initiative (CCI) land cover dataset
Normalized Difference Vegetation Index (NDVI) Terra Moderate Resolution Imaging Spectroradiometer (MODIS) Vegetation Indices (MOD13A3)
Sand, silt, clay fractions Harmonized World Soil Database (HWSD)
Elevation, slope, aspect Global Multi-resolution Terrain Elevation Data (GMTED)
Soil depth Global 1-km Gridded Thickness of Soil, Regolith, and Sedimentary Deposit Layers
Carbonate sedimentary rock fraction Global Lithological Map (GLiM)
Rock porosity, permeability GLobal HYdrogeology MaPS (GLHYMPS)
Population density Gridded Population of the World (GPW) v4 dataset
GDP per capita; population density Gross Domestic Product and Human Development Index over 1990-2015
Forest intact fraction Intact Forest Landscapes Data
Outputs Streamflow Global gauge networks
Table C7: Soil Composition (ISRIC) Variables and Sources
Variable Type Variable Name Source
Time Series Forcings Same as Table 7
Static Attributes Same as Table 7
Outputs Soil property (clay; sand; silt) The World Soil Information Service (WoSIS)
Table C8: Computation Resources for StefaLand and Comparison Models
Model Seconds/Epoch #GPUs GPU Type Memory
StefaLand (Pretraining) 16,000 6 NVIDIA V100 240 GB
StefaLand with resConn 30 2 NVIDIA V100 80 GB
StefaLand without resConn 26 2 NVIDIA V100 80 GB
LSTM Baseline 12 2 NVIDIA V100 80 GB
LSTM-HBV1.1 280 2 NVIDIA V100 80 GB
StefaLand-resConn HBV1.1 320 2 NVIDIA V100 80 GB
StefaLand-no resConn HBV1.1 300 2 NVIDIA V100 80 GB
  • Note: All values except pretraining are for the CAMELS benchmark experiment. The relative differences in computational requirements are consistent across other experiments.
Appendix D Model Architectures
Figure D1: StefaLand model architecture with residual connections. The pretrained StefaLand encoder (frozen during fine-tuning) processes static attributes and generates embeddings that are combined with meteorological forcings through residual connections. This architecture enables iterative integration of transformer features with input data through the CNN module, allowing the model to effectively leverage pretrained representations while adapting to task-specific requirements.
Figure D2: StefaLand model architecture without residual connections. In this configuration, the frozen transformer embeddings are used only once through a standard adapter that combines them with forcings and static features. This represents a more conventional fine-tuning approach where transformer features are not integrated iteratively throughout the decoding process.
Figure D3: Different adapter architectures tested in our experiments. Left: Gated Adapter with a selector mechanism that controls information flow. Center: BottleNeck Adapter with compression and expansion phases. Right: Residual Adapter that adds transformer features through a skip connection.
Appendix E Metric Calculations
This appendix details the calculation of the evaluation metrics used in our experiments. All metrics presented in the main paper tables are the median values across test basins or stations, as computed using the following formulations.
E.1 Primary Evaluation Metrics
E.1.1 Root Mean Square Error (RMSE)
RMSE measures the average magnitude of prediction errors. Lower values indicate better performance.
$\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{\text{pred},i}-y_{\text{target},i})^{2}}$ (13)
E.1.2 Unbiased Root Mean Square Error (µbRMSE)
µbRMSE removes the bias component from the error calculation, focusing on the error’s random component. It is calculated by first computing anomalies from the mean for both predictions and targets.
$y^{\prime}_{\text{pred},i}$ $=y_{\text{pred},i}-\overline{y}_{\text{pred}}$ (14)
$y^{\prime}_{\text{target},i}$ $=y_{\text{target},i}-\overline{y}_{\text{target}}$ (15)
µbRMSE $=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y^{\prime}_{\text{pred},i}-y^{\prime}_{\text{target},i})^{2}}$ (16)
E.1.3 Correlation (Corr)
Correlation quantifies the linear relationship between predictions and targets. Values range from -1 to 1, with 1 indicating perfect positive correlation.
$\text{Corr}=\frac{\sum_{i=1}^{n}(y_{\text{pred},i}-\overline{y}_{\text{pred}})(y_{\text{target},i}-\overline{y}_{\text{target}})}{\sqrt{\sum_{i=1}^{n}(y_{\text{pred},i}-\overline{y}_{\text{pred}})^{2}\sum_{i=1}^{n}(y_{\text{target},i}-\overline{y}_{\text{target}})^{2}}}$ (17)
This is calculated using Pearson’s correlation coefficient between predicted and observed values.
E.2 Secondary Metrics
The following metrics are used in our comprehensive evaluation but may not appear directly in the main tables.
E.2.1 Nash-Sutcliffe Efficiency (NSE) / $R^{2}$
NSE evaluates the predictive skill relative to using the mean of observations as a predictor. Values range from -$\infty$ to 1, with 1 indicating perfect prediction.
$\text{NSE}=1-\frac{\sum_{i=1}^{n}(y_{\text{target},i}-y_{\text{pred},i})^{2}}{\sum_{i=1}^{n}(y_{\text{target},i}-\overline{y}_{\text{target}})^{2}}$ (18)
E.2.2 Mean Absolute Error (MAE)
MAE measures the average absolute difference between predictions and targets.
$\text{MAE}=\frac{1}{n}\sum_{i=1}^{n}|y_{\text{pred},i}-y_{\text{target},i}|$ (19)
E.2.3 Flow Duration Curve RMSE (RMSE_FDC)
RMSE_FDC evaluates errors in the statistical distribution of flows rather than in their timing.
$\text{RMSE\_FDC}=\sqrt{\frac{1}{100}\sum_{j=1}^{100}(FDC_{\text{pred},j}-FDC_{\text{target},j})^{2}}$ (20)
where $FDC_{j}$ represents the $j$-th percentile of the sorted flow values.
E.2.4 Flow Biases
Several flow-specific biases were computed to evaluate performance across different flow regimes:
  • FLVFLV (Low Flow Volume Bias): Percent bias in the lowest 30% of flows
  • FHVFHV (High Flow Volume Bias): Percent bias in the highest 2% of flows
  • PBIASPBIAS (Percent Bias): Overall percent bias across all flows
The general form for these biases is:
$\text{PBIAS}_{\text{regime}}=\frac{\sum(y_{\text{pred},\text{regime}}-y_{\text{target},\text{regime}})}{\sum y_{\text{target},\text{regime}}}\times 100\%$ (21)
E.2.5 Kling-Gupta Efficiency (KGE)
KGE combines correlation, bias, and variability components:
$\text{KGE}=1-\sqrt{(r-1)^{2}+\left(\frac{\sigma_{\text{pred}}}{\sigma_{\text{target}}}-1\right)^{2}+\left(\frac{\mu_{\text{pred}}}{\mu_{\text{target}}}-1\right)^{2}}$ (22)
where $r$ is the correlation coefficient, $\sigma$ represents standard deviation, and $\mu$ represents the mean.
E.3 Metric Aggregation
For each evaluation scenario (Random Holdout and Regional Holdout), metrics were calculated for each individual basin or station and then aggregated using median values to provide a robust measure of central tendency less sensitive to outliers. All metrics shown in tables throughout the paper represent these median values across the test set.
E.4 Implementation Details
All metrics were implemented in Python using NumPy for numerical computations and SciPy’s statistical functions for correlation coefficients. Special care was taken to handle missing values (NaNs) appropriately in all calculations. For time series with missing values, only timestamps where both predicted and target values were available were used in metric calculations.
Appendix F Additional Model results
Table F1: Additional experiments with linear regression baselines.
Experiment Random holdout Regional holdout
RMSE ↓ µbRMSE ↓ Corr ↑ RMSE ↓ µbRMSE ↓ Corr ↑
Camels Streamflow Linear Regression 2.190 2.180 0.500 2.260 2.250 0.500
Global Streamflow Linear Regression 1.823 1.746 0.252 1.816 1.721 0.248
Soil Moisture Linear Regression 0.120 0.101 0.188 0.121 0.103 0.187

### v2

@@
-StefaLand: An Efficient Geoscience Foundation Model That Improves Dynamic Land–Surface Predictions
-StefaLand: An Efficient Geoscience Foundation Model That Improves Dynamic Land–Surface Predictions
-Nicholas Kraabel
-Department of Civil and Environmental Engineering
-Pennsylvania State University
-&Jiangtao Liu11footnotemark: 1
-Department of Civil and Environmental Engineering
-Pennsylvania State University
-&Yuchen Bian
-Amazon Web Services
-&Daniel Kifer
-Department of Computer Science & Engineering
-Pennsylvania State University
-&Chaopeng Shen
-Department of Civil and Environmental Engineering
-Pennsylvania State University
-Equal contribution.Corresponding author. cshen@engr.psu.edu
+StefaLand: An Efficient Geoscience Foundation Model That Improves Dynamic Land-Surface Predictions
+Abstract
 Stewarding natural resources, mitigating floods, droughts, wildfires, and landslides, and meeting growing demands require models that can predict climate-driven land-surface responses and human feedback with high accuracy.
@@
 We introduce StefaLand, a generative spatiotemporal earth foundation model centered on landscape interactions.
-StefaLand improves predictions on three tasks and four datasets: streamflow, soil moisture, and soil composition, compared to prior state-of-the-art.
+StefaLand improves predictions on four tasks and five datasets: streamflow, soil moisture, soil composition, and landslide detection compared to prior state-of-the-art.
 Results highlight its ability to generalize across diverse, data-scarce regions and support broad land-surface applications.
@@
 1 Introduction
-Climate change is ushering in strong and widespread changes on the land surface, including higher frequencies of floods, droughts, wildfires and other geohazards Ebi et al.
-(2021), IPCC (2021).
+Climate change is ushering in strong and widespread changes on the land surface, including higher frequencies of floods, droughts, wildfires and other geohazards (Ebi et al., 2021, IPCC, 2021).
 To mitigate the impact of these disasters, there are urgent needs for models that can accurately predict land surface dynamics such as streamflow, soil moisture, soil composition, landslides, snow water equivalent, groundwater levels, and vegetation carbon content.
-Among these, soil moisture controls the partition of rainfall into infiltration and runoff, modulates flood generation and landslides, and critically influences land-atmosphere interactions Dorigo et al.
-(2013).
+Among these, soil moisture controls the partition of rainfall into infiltration and runoff, modulates flood generation and landslides, and critically influences land-atmosphere interactions (Dorigo et al., 2013a).
 Streamflow is the flow rate of water running in the rivers, the most accessible water resource to humans, and too high or too low streamflow can cause flooding or hydrologic drought, respectively.
@@
 Predicting these dynamic variables is distinct from image-recognition tasks, as here we seek to predict what will happen in the near or distant future.
-Traditionally, these tasks were undertaken by physically-based models that take atmospheric forcings (precipitation, temperature) as inputs and sequentially calculate the physical processes that eventually lead to the variables of interest Li et al.
-(2015).
-In recent years, there has been a proliferation of data-driven machine learning (ML) models Solomatine and Ostfeld (2008).
+Traditionally, these tasks were undertaken by physics-based models that take atmospheric forcings (precipitation, temperature) as inputs and sequentially calculate the physical processes that eventually lead to the variables of interest (Li et al., 2015).
+In recent years, there has been a proliferation of data-driven machine learning (ML) models (Solomatine and Ostfeld, 2008).
 These models are often set up to accept forcing (dynamic weather) and landscape characteristics (static) data as inputs, and are trained to directly predict the natural land surface variables given the weather inputs.
 However, up to now, most of the geoscientific ML models have been supervised ML approaches trained specifically for a narrow set of tasks.
-Foundation models, which are trained on broad tasks to grasp the joint data distribution, have just started to emerge in geosciences, e.g., Terramind, Prithvi and Aurora.
-However, they were built mainly on satellite images for landcover-identification tasks Jakubik et al.
-(2023; 2025), Schmude et al.
-(2024).
-There has been a notable absence of foundation models built for dynamical land surface modeling, and a lack of effort to leverage valuable temporal datasets and ground-based observations Xie et al.
-(2023).
+A particularly notable gap is that currently established geoscience foundation models have largely been trained on satellite imagery for landcover-identification tasks (Jakubik et al., 2023; 2025, Schmude et al., 2024), which limits their ability to capture dynamic land–atmosphere interactions.
+As a result, valuable temporal datasets and ground-based observations remain underutilized (Xie et al., 2023), and to our knowledge no foundation model has yet been developed with a primary focus on dynamical land surface modeling.
 A grand challenge for geoscientific ML models is to improve their spatial generalization, because a frequent issue facing them is the sparsity and spatial imbalance of observational data.
-Due to the cost of installing instruments and varying policies on data sharing, data are often only available in high density in certain regions of some developed nations.
-For example, streamflow gauge data are available at high density in United States, Europe, Australia and Japan, but are scanty in Africa, South America, and the rest of the Asia Global Runoff Data Centre (2020).
+Satellite data are often coarse in resolution and uncertain compared to in-situ measurements.
+SMAP, for example, provides global soil moisture observations at  9–36 km resolution every 2–3 days, which are useful for regional climate and hydrologic research but far less valuable than in-situ probes for operational field tasks such as irrigation scheduling or crop stress monitoring (Entekhabi et al., 2010).
+However, in-situ data, due to the cost of installing instruments and varying policies on data sharing, is only available in high density in certain regions of some developed nations.
+For example, streamflow gauge data are abundant in the United States, Europe, Australia and Japan, but remain sparse in Africa, South America, and much of Asia (Global Runoff Data Centre, 2020).
 Similar distribution patterns are found for high-quality in-situ soil moisture probes and soil property measurements.
-Satellite data have too coarse resolution and too large uncertainty levels compared to in-situ measurements.
-As quantified in many studies Feng et al.
-(2023), a deep network trained using data from some regions can face substantial performance degradation when applied in data-scarce regions.
-This is partly because there are not enough sites in space to learn the true dependencies of the targets on static land surface characteristics, and partly because of systematic data discrepancies among different regions (concept drift).
-This means ML land surface predictions may be unusable in more than 70% of the land.
-This limitation is particularly acute in data-scarce regions that are often most vulnerable to climate impacts.
-Any method that can reliably improve spatial generalization can be a significant and rare advance to greatly democratizes information for stakeholders worldwide.
-Related Work: In hydrologic and ecosystem predictions, long short-term memory (LSTM) networks Hochreiter and Schmidhuber (1997) remain the dominant architecture, in part because land surface processes often behave like Markov processes where LSTMs’ gating mechanisms handle noisy continuous inputs well.
-Attempts to adopt transformers, so successful in natural language processing, have generally underperformed in hydrologic regression tasks Xue et al.
-(2023), Liu et al.
-(2024), with evidence of overfitting on continuous signals Zeng et al.
-(2022).
-Nonetheless, recent studies show that with task-specific modifications and careful fine-tuning, transformers can achieve competitive results and even offer advantages in autoregressive forecasting and extreme event prediction Wen et al.
-(2023), precisely the areas where current hydrologic models struggle most with spatial generalization.
-Traditional hydrologic research on "prediction in ungauged basins" (PUB) have examined regionalization and spatial interpolation approaches including clustering or classifying catchments and transferring parameters from donor catchments in the same class Hrachowitz et al.
-(2013), Yang et al.
-(2023).
-Such an expert-derived design represents a crude practice of unsupervised learning that indicate the importance of understanding the joint data distribution.
+As quantified in many studies (Feng et al., 2023), a deep network trained on data from some regions can face substantial performance degradation when applied in data-scarce regions.
+This occurs partly because there are not enough sites to learn the true dependencies of the targets on static land surface characteristics, and partly because of systematic data discrepancies across regions (concept drift).
+While such limitations hinder traditional supervised ML models, foundation models offer a potential path forward: by jointly learning from broad, heterogeneous datasets (including temporal records and ground observations where available), they may transfer useful representations to data-scarce regions where task-specific training data are limited.
+Related Work: In hydrologic and ecosystem predictions, supervised long short-term memory (LSTM) networks (Hochreiter and Schmidhuber, 1997) remain a highly popular architecture, in part because land surface processes often behave like Markov processes where LSTMs’ gating mechanisms handle noisy continuous inputs well(Kratzert et al., 2018).
+Attempts to adopt transformers, so successful in natural language processing, have generally found it difficult to noticeably surpass LSTM in time series regression tasks (Xue et al., 2023, Liu et al., 2024), with evidence of overfitting on continuous signals (Zeng et al., 2022).
+Nonetheless, recent studies show that with task-specific modifications and careful fine-tuning, transformers can achieve competitive results in extreme event prediction (Wen et al., 2023), precisely the areas where current hydrologic models struggle most with spatial generalization.
+Traditional hydrologic research on "prediction in ungauged basins" (PUB) have examined regionalization and spatial interpolation approaches including clustering or classifying catchments and transferring parameters from donor catchments in the same class (Hrachowitz et al., 2013, Yang et al., 2023).
+Such an expert-derived design represents a crude practice of unsupervised learning that indicates the importance of understanding the joint data distribution.
 However, modern weakly-supervised foundation models can, in general, much better grasp the joint data distribution than expert-driven approaches.
 Foundation models offer a promising approach to address these spatial generalization challenges.
-By pretraining on large-scale datasets to learn generalizable representations, these models can potentially transfer knowledge across regions and geoscientific domains Zhang et al.
-(2024).
-Existing geoscience foundation models such as TerraMind, Prithvi, Aurora, and AlphaEarth have largely focused on satellite imagery Hsu et al.
-(2024), which, while providing global coverage, may not capture the temporal dynamics and physical processes most relevant to land surface predictions.
-For example, TerraMind is trained on 9 million globally distributed satellite imagery samples across various modalities (optical, radar, elevation, land use) to create a generative multimodal foundation model for Earth observation Jakubik et al.
-(2023).
-It has been claimed to serve as "any-to-any generative, multimodal foundation model for Earth observation" Jakubik et al.
-(2025).
+By pretraining on large-scale datasets to learn generalizable representations, these models can potentially transfer knowledge across regions and geoscientific domains (Zhang et al., 2024).
+Existing geoscience foundation models, e.g., TerraMind (Jakubik et al., 2025) and Prithvi (Hsu et al., 2024) from IBM, and Aurora (Bodnar et al., 2025) from Microsoft, have largely focused on satellite imagery, which may not capture the temporal dynamics and physical processes most relevant to land surface predictions.
+For example, TerraMind is trained on 9 million globally distributed satellite imagery samples across various modalities (optical, radar, elevation, land use) to create a generative multimodal foundation model for Earth observation (Jakubik et al., 2023).
+It has been claimed to serve as "any-to-any generative, multimodal foundation model for Earth observation" (Jakubik et al., 2025).
 However, many critical variables regarding soil and the subsurface are not directly observable from space.
-In addition, satellite images are noisy and a large portion of the data is redundant and irrelevant across repeated revisits.
-Therefore, it is unclear whether they are of high value to land-surface dynamical prediction tasks.
-Our contributions: We present Spatiotemporal Earth Foundation model with Attributes for the Land Surface (StefaLand), the first land-focused geoscientific foundation model designed for dynamic land–surface prediction.
-StefaLand improves predictions on streamflow, soil moisture, and soil composition compared to state-of-the-art baselines.
-StefaLand further demonstrates strong spatial generalization across diverse landscapes and data-scarce regions.
-StefaLand’s attribute-based rather than image-based design (with the potential to link to image-like inputs in the future) is intentionally aligned with physical processes, drastically reducing compute requirements while retaining global coverage, making it accessible to academic researchers with modest resources.
-Pretraining requires only about 720 GPU hours, which makes it feasible within academic budgets.
+In addition, satellite images are noisy and a large portion of the data is redundant across repeated revisits or irrelevant for hydrologic processes.
+Therefore, it is unclear whether they are of high relevance to land-surface dynamical prediction tasks.
+Our contributions: We present Spatial-Temporal Earth Foundation model with Attributes for the Land Surface (StefaLand), the first land-focused geoscientific foundation model designed for dynamic land–surface prediction.
+StefaLand improves predictions on streamflow, soil moisture, soil composition, and landslide susceptibility compared to state-of-the-art supervised learning baselines as well as finetuned satellite-image-trained earth foundation models.
+Especially, it shows strong spatial generalization across diverse landscapes and data-scarce regions for a wide variety of tasks.
+StefaLand’s attribute-based rather than image-based design (with the potential to link to image-like inputs in the future) incorporates a variety of ground-based measurement data, emphasizes relevant land-surface physical processes, drastically reduces compute requirements while retaining global coverage, making it accessible to researchers with modest resources.
+Pretraining our model required only about 720 V100 GPU hours (could be shorter with more advanced GPUs).
 The model builds on a masked autoencoder backbone, a location-aware fusion of static and time-series inputs, grouped masking to promote cross-attribute interactions, and residual fine-tuning adapters, into a coherent design guided by geoscientific knowledge.
-Taken together, these contributions establish StefaLand as an efficient and accessible complement to vision-based foundation models, purpose-built for hydrologic and land–surface applications.
+Taken together, these contributions establish StefaLand as an efficient and accessible complement to vision-based foundation models.
 2 Methods
-Dynamic land–surface prediction requires integrating heterogeneous information: landscape attributes such as topography, soils, and geology, together with dynamic meteorological forcings like precipitation and temperature.
-Capturing the joint influence of these variables is essential for spatial generalization but challenging with standard architectures.
-StefaLand addresses this gap through an attribute-based transformer framework that unifies static and time-varying signals in a shared representation space.
-2.1 Model Architecture Overview
-StefaLand processes both dynamic time series and static attributes through a unified encoder-decoder architecture, and is trained as a attribute-group-based masked autoencoder (Figure D1).
-The model embeds multivariate inputs, applies strategic masking to force the learning of cross-domain dependencies, and reconstructs the masked portions to generate deep representations.
-Our location-aware transformer encoder captures spatial relationships while positional encodings maintain temporal context.
-The complete mathematical formulation of our model architecture is provided in Appendix A.
-As well as a visual representation of the model is provide in appendix D.
-2.2 Pretraining Setup
-Our pretraining approach uses Cross-Variable Group Masking (CVGM), which forces the model to capture interactions among groups of landscape variables (e.g., terrain, climate, soil, vegetation) rather than treating them independently.
-This encourages the learning of physical relationships between different environmental drivers.
-The objective is a reconstruction loss on the masked slice of the sequence, normalized by variable-wise standard deviations when available and with per-variable weighting via a learnable weight vector.
-During pretraining, the dataset is stored in shards rather than fully loaded into memory.
-At the start of each epoch, we randomly reindex these shards, which effectively reshuffles the available samples.
-This lazy-loading strategy avoids materializing the entire corpus in memory while still ensuring high sample diversity across epochs.
-The pretraining dataset is a derived global attribute dataset spanning 4,229 locations over 40 years.
-A complete list of variables and their sources is provided in Appendix C.
-The accompanying code release will enable reconstruction of this dataset.
-Hyperparameter settings and configurations for pretraining are also listed in the appendix above.
-2.3 Transformer-based Direct Prediction Models
-We developed several finetuned transformer models that combine StefaLand’s base structure with new units to directly predict streamflow and soil moisture.
-StefaLand with residual connection (resConn) is our main proposed architecture, integrating pretrained StefaLand transformers with an LSTM decoder through a residual pathway (Figure D1).
-This architecture enables iterative integration of pretrained transformer features throughout the decoding process, combining frozen transformer embeddings with forcings via a convolutional neural network before merging with the decoder through skip connections.
-This enables effective integration of pretrained features with task-specific requirements.
-Detailed adapter architectures are described in Appendix A.
-For comparisons, "StefaLand without resConn" utilizes transformer outputs only once through a simple adapter (Figure D2), and a comparison with this model highlights the value of the resConn structure.
-"StefaLand Transformer" represents the base StefaLand model trained from scratch with the finetuning dataset.
-"StefaLand Ablation - resConn" utilizes the same structure as "StefaLand - resConn", but randomize the weights of StefaLand instead of using the pretrained weights.
-Comparing with these two models demonstrates the value of pretraining.
-As discussed below, we also implemented Terramind with resConn, which uses IBM’s Terramind foundational model together with our residual connection architecture, to contrast the effectiveness of Terramind vs StefaLand pretraining.
-Of course, supervised LSTM serves as a widely employed and highly-competitive baseline.
-2.4 Physics-Based Integration
-To leverage domain knowledge, we also implemented physics-based configurations that combine our data-driven approach with process understanding.
-For the process-based backbone, we employed the Hydrologiska Byråns Vattenbalansavdelning (HBV) model, a conceptual hydrologic model with state variables like snow storage, soil water, and subsurface storage.
-In this setup, the neural network only provides parameters for the process-based backbone, which then outputs a range of interpretable variables, and they are trained together in an end-to-end fashion.
-Our physics-based approaches include LSTM-HBV1.1, a baseline configuration where LSTM outputs parameters for the HBV model; StefaLand-resConn HBV1.1; and StefaLand-no resConn HBV1.1, which integrates transformer features without residual connections.
-These physics-based approaches maintain interpretability while leveraging improved representations from StefaLand.
-Detailed implementations are provided in Appendix B.
+Dynamic land–surface prediction requires combining heterogeneous information: static landscape attributes such as topography, soils, vegetation, and geology, together with dynamic forcings such as precipitation and temperature.
+StefaLand addresses this challenge with a transformer-based masked autoencoder that jointly embeds static and dynamic variables, pretrained with a cross-variable masking strategy, and then adapted for specific prediction tasks with task-specific heads.
+2.1 StefaLand Transformer Architecture
+Embedding.
+Each dynamic input variable $c$ at time $t$ is independently embedded with a nonlinear two-layer projection:
+$z_{t,c}=\text{GELU}(x_{t,c}W_{1,c}+b_{1,c})W_{2,c}+b_{2,c},$ (1)
+where $W_{1,c}\in\mathbb{R}^{1\times 64}$ and $W_{2,c}\in\mathbb{R}^{64\times 256}$. Summing across all $C$ dynamic variables yields a per-step embedding:
+$z_{t}=\sum_{c=1}^{C}z_{t,c}.$ (2)
+Static attributes $s_{i}$ are embedded similarly:
+$z_{\text{static},i}=\text{GELU}(s_{i}W_{1,i}+b_{1,i})W_{2,i}+b_{2,i},$ (3)
+and concatenated as a special static token, producing the joint sequence:
+$Z=[z_{1};z_{2};\ldots;z_{T};z_{\text{static}}].$ (4)
+Cross-Variable Group Masking (CVGM).
+Let $G=\{g_{1},\ldots,g_{K}\}$ be groups of related variables. During pretraining, a temporal window $\tau,\tau+\ell)$ is sampled, and entire groups are stochastically masked:
+$m_{k}\sim\text{Bernoulli}(p_{\text{mask}}).$ (5)
+For all $c\in g_{k}$ with $m_{k}=1$, the embedding is replaced by a learned mask vector $\mathbf{m}_{c}\in\mathbb{R}^{256}$.
+This forces reconstructions to rely on cross-variable dependencies rather than treating drivers independently.
+Transformer encoder.
+After adding learnable positional encodings $P$, we obtain:
+$\tilde{Z}=Z+P.$ (6)
+This sequence is processed by $N$ Transformer blocks:
+$A^{(\ell)}$ $=\text{MHA}(H^{(\ell-1)}),$ (7)
+$\tilde{H}^{(\ell)}$ $=\text{LayerNorm}(H^{(\ell-1)}+A^{(\ell)}),$ (8)
+[…]
+$H^{(\ell)}$ $=\text{LayerNorm}(\tilde{H}^{(\ell)}+F^{(\ell)}).$ (10)
+Decoder and loss.
+The final hidden states are passed through a bidirectional LSTM to refine local temporal continuity:
+$U=\text{LSTM}(H^{(N)}W_{\text{proj}}+b_{\text{proj}}).$ (11)
+Outputs corresponding to masked groups are projected back to the original variable dimensions. The reconstruction loss is
+$\mathcal{L}=\sum_{c\in\mathcal{M}}w_{c}\frac{\|\hat{x}_{c}-x_{c}\|_{2}^{2}}{\sigma_{c}^{2}},$ (12)
+where $\mathcal{M}$ is the set of masked variables, $\sigma_{c}$ is the standard deviation of variable $c$, and $w_{c}$ is a learnable weight.
+2.2 Pretraining Details
+The pretraining dataset is a derived global attribute dataset spanning $\sim$8,000 locations (basins) over 40 years.
+Variables were chosen to represent the key controls on fluxes of water, energy, momentum, sediment, and nutrients.
+A complete list of variables, their group assignments, and their sources is provided in Appendix C.
+The CVGM is designed so variables with reciprocal or bidirectional causality are masked together, preventing them from acting as predictors for each other.
+Most groupings are straightforward, such as masking silt and clay percentages together.
+One exception is soil depth, which we classify as a terrain attribute since it is strongly tied to topographic derivations.
+By masking and reconstructing at the group level, the model is encouraged to capture cross-domain interactions, such as the coupling between soil texture and climate seasonality.
+While the grouped-masking strategy has sometimes been employed in multimodality AI models, we did not find it applied in earth foundation models like Terramind, Prithvi, or Aurora, which instead use random dropout, spatial patches, or future periods for masking.
+Our objective is a reconstruction loss on the masked slice of the sequence, normalized by variable-wise standard deviations when available and with per-variable weighting via a learnable weight vector.
+2.3 Finetuning for Prediction Tasks
+Our primary finetuning model, StefaLand-resConn, integrates pretrained embeddings with raw forcings through a residual pathway (Figure D1).
+Let $E_{t}$ denote the transformer embedding at time $t$, and $x_{t}$ the raw forcings.
+We compute:
+$r_{t}$ $=f_{\text{conv+linear}}(x_{t}),$ (13)
+$h_{t}$ $=\text{LSTM}(E_{t}+r_{t}),$ (14)
+$\hat{y}_{t}$ $=W_{o}h_{t}+b_{o}.$ (15)
+Skip connections propagate $E_{t}$ and $r_{t}$ into intermediate layers, allowing general knowledge (from $E_{t}$) to be iteratively refined with task-specific signals ($r_{t}$).
+This design strengthens spatial generalization while adapting to local variability.
+To see a visual representation of this D1.
 2.4 Foundation model comparisons
-For completeness, we also evaluated two existing Earth-Observation-oriented foundation models, TerraMind and PrithviWxC Jakubik et al.
-(2025), Hsu et al.
-(2024), using the same fine-tuning heads and training protocols as StefaLand.
-TerraMind was tested on streamflow and soil moisture, and PrithviWxC on soil moisture.
-Inputs were harmonized to each model’s modality.
-It is noteworthy that either Terramind or PrithviWxC requires orders of magnitude larger storage space for the input data than StefaLand, as well as computational cost for finetuning.
-Due to intensive data requirements for PrithviWxC on surface-level variables mostly likely to demonstrate relevance to land surface interactions, along with all 14 static variables were used.
+For completeness, we also evaluated two existing Earth-Observation-oriented foundation models, TerraMind and PrithviWxC (Jakubik et al., 2025, Hsu et al., 2024), using the same fine-tuning heads and training protocols as StefaLand.
+We coupled them together with our residual connection architecture and finetuned the added units while keeping the foundation models’ weights frozen.
+Because both TerraMind and PrithviWxC require orders of magnitude larger storage space, data transfer, and fine-tuning costs than StefaLand, we chose comparison tasks selectively, focusing on cases where their pretraining data was potentially most relevant.
+Accordingly, we compared both models on soil moisture, while streamflow was included only in a more exploratory capacity (only TerraMind).
+For a rough comparison, StefaLand, TerraMind and PrithviWxC used respectively $\approx 2$, $\approx 11$ and $\approx 27$ TB data during pretraining.
+Due to the intensive data requirements for PrithviWxC, only surface-level variables likely relevant to land surface interactions, along with all 14 static variables, were used.
 The full atmospheric variables at differing pressure levels were excluded.
 3 Experiments
-For all experiments described below, forcing data and static features remained consistent across all model configurations, with hyperparameters held constant within each test type except for fundamental architectural differences inherent to transformer models.
-Complete details of hyperparameters, forcings, and static features are provided in the Appendix C.
-In addition to the neural and process-based baselines reported below, we trained plain linear regression models for every task using the same features (and appropriate standardization).
-These consistently underperformed all learning baselines by large margins, so we omit them from tables for readability; full metrics are provided in Appendix E.
+We tested the value of foundation model pretraining on 5 datasets and 6 experiments, including, streamflow on the CAMELS dataset on USA, CAMELS streamflow prediction with hybrid model, global streamflow, global in-situ soil moisture, global soil properties, and landslide susceptibility in Oregon, USA.
+For all experiments, hyperparameters were tuned with Ray Tune and kept consistent across model configurations within each experimental case (e.g., CAMELS streamflow, soil moisture, etc).
+Because we compare spatial generalization, we used temporal validation splits for hyperparameter optimization.
+Complete details of hyperparameters, forcings, and static features for all four experiments are in Appendix C.
+Model variants used.
+Unless noted otherwise, pretrained encoders are frozen and only task heads are trained.
+We evaluate: (i) LSTM-SL: supervised LSTM baseline; (ii) StefaLand-direct: the StefaLand encoder together with its original decoder trained directly on the task data, without pretraining; (iii) StefaLand-resConn: pretrained encoder with residual pathway + LSTM decoder, our proposed architecture; (iv) StefaLand-noResConn: pretrained encoder with a simple adapter (no residual/LSTM); (v) StefaLand-scratch: same resConn head but StefaLand initialized randomly (no pretraining) and unfrozen (ablation test); (vi) TerraMind-resConn and PrithviWxC-resConn: EO/atmospheric foundation encoders frozen with the same residual head.
 3.1 CAMELS Streamflow Prediction
-To compare the scheme’s spatial generalization capability on a well benchmarked dataset, we follow the experimental setup in (Feng et al., 2021), testing prediction in ungauged basins (PUB) and ungauged regions (PUR).
-These correspond to randomized spatial K-fold and regional-specific K-fold regimes, reflecting real-world challenges in generalizing to nearby unmonitored basins and to data-scarce regions, respectively.
-We use the CAMELS dataset (Addor et al., 2017, Newman et al., 2014), restricted to the 531-basin subset with clear watershed boundaries (Newman et al., 2017).
+To compare spatial generalization on a well benchmarked dataset, we follow (Feng et al., 2021), testing prediction in ungauged basins (PUB) and ungauged regions (PUR).
+These correspond to randomized spatial K-fold and regional-specific K-fold regimes, respectively.
+We use CAMELS (Addor et al., 2017, Newman et al., 2014), restricted to the 531-basin subset with clear watershed boundaries (Newman et al., 2017).
 Basins were divided into 10 random groups for PUB and 7 contiguous regions for PUR, employing leave-one-out in both cases.
-To avoid leakage, all CAMELS-overlapping stations were removed during pretraining for PUB tests, and entire regions were excluded for PUR tests.
-We evaluated both direct prediction models and neural networks used to parameterize process-based models (see Methods).
-Direct models included StefaLand without finetuning (StefaLand-Transformer), pretrained variants with and without residual connections (resConn / noResConn), a from-scratch ablation trained only on task data, and a version using TerraMind as encoder.
-For process-based models, we tested LSTM HBV1.1, StefaLand-resConn HBV1.1, and StefaLand-noResConn HBV1.1.
+To avoid leakage, all CAMELS-overlapping stations were removed during pretraining for PUB, and entire regions were excluded for PUR.
 Table 1: CAMELS Streamflow PUB and PUR Results
 Model Random holdout (ungauged basins) Regional holdout (ungauged regions)
-RMSE $\downarrow$ pbRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$ RMSE $\downarrow$ pbRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$
-Direct Prediction Models
+RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$ RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$
 LSTM - SL 1.402 1.360 0.762 0.636 1.609 1.457 0.743 0.554
-StefaLand - Transformer 1.882 1.849 0.538 0.395 1.982 1.949 0.230 0.201
+StefaLand - direct 1.882 1.849 0.538 0.395 1.982 1.949 0.230 0.201
 StefaLand - resConn 1.111 1.068 0.869 0.717 1.344 1.334 0.801 0.635
 StefaLand - no resConn 1.171 1.154 0.823 0.706 1.376 1.356 0.798 0.610
+StefaLand - scratch 1.355 1.332 0.801 0.661 1.516 1.378 0.771 0.560
 TerraMind - resConn 1.332 1.301 0.777 0.637 1.420 1.398 0.763 0.551
-StefaLand Ablation - resConn 1.355 1.332 0.801 0.661 1.516 1.378 0.771 0.560
-Physics Based Models
-LSTM - HBV1.1 1.325 1.298 0.857 0.672 1.561 1.521 0.746 0.578
-StefaLand - resConn HBV1.1 1.234 1.216 0.863 0.714 1.345 1.332 0.842 0.643
-StefaLand - no resConn HBV1.1 1.315 1.302 0.848 0.707 1.401 1.379 0.835 0.623
-StefaLand Ablation - resConn HBV1.1 1.310 1.306 0.842 0.693 1.465 1.432 0.607 0.512
-The original multi-basin LSTM studies on CAMELS defined the modern supervised learning benchmark (Kratzert et al., 2019), and subsequent large-scale comparisons confirmed that vanilla Transformers generally fail to outperform LSTMs on rainfall–runoff prediction (Liu et al., 2024, Table 1 therein).
-This is again reinforced in our results, where StefaLand-Transformer underperforms LSTM.
-The Nash-Sutcliffe model efficiency coefficient (NSE) values reported here for LSTM are highly similar to those reported in the domain literature Feng et al.
-(2021).
+The foundation model pretraining clearly provides a rare boost to generalization capability compared to supervised learning models (LSTM-SL and StefaLand-direct) across both PUB and PUR (Table 1).
+StefaLand-resConn’s RMSE is almost $20\%$ lower than that of supervised LSTM in PUB and $17\%$ lower in PUR, while the correlation are noticeably higher.
+The value of pretraining is confirmed by StefaLand-scratch, which appears to have only minor advantage compared to LSTM.
+The residual-connection architecture enables effective integration of transformer features with temporal dynamics, showing a modest benefit compared to StefaLand-no resConn, but both are clearly stronger than either LSTM or StefaLand-direct.
+Finally, TerraMind-resConn has quite comparable metrics to LSTM; the pretrained TerraMind offers no extra generalization value in this case.
+We ran additional experiments that hybridize StefaLand with the HBV1.1 physics backbone on the same PUB/PUR splits, testing its ability to parameterize physics-based models.
+These hybrids achieved up to a 13% RMSE reduction and a 10% correlation gain compared to the LSTM–HBV1.1 baseline, demonstrating that pretraining strengthens hybridization with process models.
+By constraining predictions with physics while leveraging StefaLand features, these hybrids further improve upon the general results above and highlight the versatility of the approach.
+Full results are provided in Appendix B, Table B1.
+On a related note, the supervised LSTM is not an easy benchmark to surpass.
+The original multi-basin LSTM and subsequent large-scale comparisons (Kratzert et al., 2019, Feng et al., 2021) showed that vanilla Transformers generally fail to outperform LSTMs on rainfall–runoff prediction (Liu et al., 2024, Table 1 therein).
+The LSTM NSE values reported here are very similar to those in the domain literature (Feng et al., 2021).
 Broader evaluations likewise continue to rank LSTM-family models among the best-performing approaches (Lees et al., 2021), and even Google’s global flood-forecasting system adopts an encoder–decoder LSTM backbone (Nearing et al., 2024).
-We therefore use LSTM-SL as the supervised SOTA comparator, complemented by a process-informed LSTM–HBV1.1 variant, and additionally test TerraMind to control for the role of EO-centric pretraining.
-Against these strong baselines, the foundation model pretraining clearly provides a rare boost to generalization capability compared to supervised learning models (LSTM-SL and StefaLand-Transformer) across both PUB and PUR scenarios (Table 1).
-StefaLand-resConn’s RMSE is almost $20\%$ lower than that of supervised LSTM in the PUB and $17\%$ lower in the harder PUR test, while the correlation are noticebly higher.
-The value of pretraining is confirmed by StefaLand-Ablation-resConn, which appears to have only minor advantage compared to LSTM.
-The residual-connection architecture enables effective integration of transformer features with temporal dynamicsh, showing a modest benefit compared to StefaLand-no resConn, but both are clearly stronger than either LSTM or StefaLand-Transformer.
-Finally, Terramind-resConn has quite comparable metrics to LSTM.
-It appears the pretrained Terramind offers no generalization value.
-The foundation model showed its versatility in supporting the parameterization of hybrid models and improving their generalization.
-We observe the same patterns — the pretrained StefaLand is helpful and resConn is a valuable construct.
-The physics-based experiments highlight that residual connections can combine pretrained representations with process-based constraints, producing more reliable spatial generalization than even the strongest LSTM configurations.
-We note that the hybrid model in fact showed a slightly better NSE than the direction prediction models for PUR for either LSTM-HBV1.1 or StefaLand-resConn-HBV1.1, because the physics-based equations serve as additional constraints to mitigate overfitting.
 3.2 Global Streamflow
-We designed a global-scale runoff prediction experiment to assess the robustness and generalization ability worldwide.
-We filtered global basin datasets based on data completeness and retained 3,434 basins for the global experiment.
-To manage computational costs efficiently, we employed random hold-out sampling combined with three-fold cross-validation.
-Additionally, we implemented a regionally hold-out continental scenario (RH-C), excluding all basins from North America, South America, and Europe, respectively, from the training set.
-The model trained from remaining continents was then evaluated on the excluded continent, simulating the realistic scenario of deploying the model to unseen geographical domains.
+We designed a global-scale runoff prediction experiment to assess robustness and generalization worldwide.
+We filtered global basin datasets based on data completeness and retained 3,434 basins.
+To manage computational cost, we employed random hold-out sampling with three-fold cross-validation.
+Additionally, we implemented a regionally hold-out continental scenario (RH-C), excluding all basins from North America, South America, and Europe, respectively, from training, then evaluating on the excluded continent.
 Table 2: Global streamflow prediction across 3,434 basins worldwide
@@
 LSTM - SL 0.870 0.864 0.798 1.253 1.202 0.672
-StefaLand - Transformer 0.749 0.751 0.843 1.075 1.048 0.697
-Terramind + finetuning 1.156 1.111 0.580 1.234 1.158 0.670
+StefaLand - direct 0.749 0.751 0.843 1.075 1.048 0.697
+TerraMind - resConn 1.156 1.111 0.580 1.234 1.158 0.670
 Cross-continental transfer testing using three-fold validation. Detailed metric calculations in Appendix E.
-Results indicate that StefaLand-Transformer outperformed both LSTM-SL and the fine-tuned vision-based Transformer model (TerraMind) across all evaluation metrics.
-Specifically, in the RH scenario, StefaLand achieved an RMSE of 0.749, approximately 14% lower than LSTM (0.87), with Corr improving to 0.843.
+Results indicate that StefaLand-direct outperformed both LSTM-SL and the fine-tuned vision-based Transformer model (TerraMind) across all metrics for either random or regional holdout (Table 2).
+Specifically, in the random holdout scenario, StefaLand achieved an RMSE of 0.749, approximately 14% lower than LSTM (0.87), with Corr improving to 0.843.
 Notably, StefaLand’s unbiased RMSE closely matched its RMSE, indicating errors were primarily random fluctuations rather than systematic bias.
 In contrast, TerraMind exhibited a higher RMSE of 1.156 and a low Corr of 0.580.
-We hypothesize that this is because the pretraining satellite image data for the vision foundation network TerraMind, did not have high relevance to hydrologic predictions.
-Under the more challenging RH-C scenario, all models experienced increased errors, but StefaLand continued to demonstrate higher performance.
+We hypothesize this is because the pretraining satellite image data for the vision foundation network TerraMind did not have high relevance to hydrologic predictions like streamflow.
+We note that using TerraMind here was exploratory, as it was not specifically designed or pretrained for this use case.
+Under the more challenging RH-C scenario, all models exhibited larger errors, but StefaLand continued to perform best.
 Its RMSE was 1.075, approximately 14.1% lower than LSTM’s 1.253, while also maintaining the highest Corr (0.697).
-Moreover, StefaLand’s ubRMSE remained close to its RMSE, confirming its robust ability to correct regional-scale biases even under extreme out-of-domain conditions.
-Conversely, LSTM and TerraMind exhibited large gaps between ubRMSE and RMSE, highlighting their challenging in producing hydrologically-relevant features.
+Moreover, StefaLand’s ubRMSE remained close to its RMSE, confirming robust correction of regional-scale biases even under extreme out-of-domain conditions.
+Conversely, LSTM and TerraMind showed larger gaps between ubRMSE and RMSE, highlighting challenges in producing hydrologically-relevant features.
 3.3 Global Soil Moisture
-We evaluated finetuning StefaLand for soil moisture predictions following the experimental design from Liu et al.
-(2023a), using data from the International Soil Moisture Network (ISMN) Dorigo et al.
-(2011; 2013).
-Even though there is a globally covering satellite-based product for soil moisture, the data quality can hardly match that of in-situ moisture sensors.
-Thus the ability to generalize in-situ data will still be valuable.
-ISMN consists of 1,316 ground-based measurement stations.
-We randomly partitioned sites into five groups for spatial cross-validation in the random holdout test, training on four groups and testing on the fifth, rotating through all groups to obtain comprehensive spatial generalization performance.
-For the regional holdout test, we specifically evaluated model performance on Europe as an ungauged region, training on all other continents while excluding European sites entirely to assess cross-continental transferability, removing 129 sites for testing.
-We tested the following model configurations: StefaLand-Transformer (no finetuning, direct prediction), StefaLand with and without residual connections (resConn / noResConn), an ablation trained from scratch without pretraining, a baseline LSTM, and a version using IBM’s TerraMind encoder with resConn.
+We evaluated finetuning StefaLand for soil moisture predictions following Liu et al.
+(2023a), using ISMN (Dorigo et al., 2011; 2013a).
+Even though there is a globally covering satellite-based product for soil moisture, the data quality can hardly match that of in-situ moisture sensors; thus the ability to generalize in-situ data is valuable.
+ISMN consists of 1,316 ground-based stations.
+We performed five-fold spatial cross-validation for random holdout and a regional holdout on Europe, training on all other continents while excluding European sites (129) for testing.
+We tested StefaLand-direct, StefaLand with and without residual connections (resConn / noResConn), a from-scratch ablation, an LSTM baseline, and a version using IBM’s TerraMind encoder with resConn.
+LSTM again serves as the established state-of-the-art baseline (Wang et al., 2024, Liu et al., 2023b).
 Table 3: Soil moisture prediction across 1,316 ISMN stations
@@
 LSTM - SL 0.073 0.055 0.764 0.112 0.053 0.510
-StefaLand - Transformer 0.140 0.103 0.637 0.135 0.112 0.503
+StefaLand - direct 0.140 0.103 0.637 0.135 0.112 0.503
 StefaLand - resConn 0.068 0.054 0.783 0.090 0.059 0.638
 StefaLand - no resConn 0.075 0.057 0.741 0.095 0.058 0.545
+StefaLand - scratch 0.074 0.058 0.749 0.108 0.064 0.528
 TerraMind - resConn 0.083 0.062 0.694 0.101 0.080 0.519
-PrithviWcX - resConn 0.081 0.060 0.703 0.103 0.079 0.523
-Ablation StefaLand - resConn 0.074 0.058 0.749 0.108 0.064 0.528
+PrithviWxC - resConn 0.081 0.060 0.703 0.103 0.079 0.523
 5-fold spatial validation and cross-continental validation on Europe (129 sites). Detailed metric calculations in Appendix E.
-LSTM again serves here as the established state-of-the-art baseline.
-Comprehensive reviews have shown that LSTM variants consistently provide the most stable and effective performance among supervised learning methods for soil moisture prediction due to their temporal modeling strengths (Wang et al., 2024).
-On the ISMN-linked global dataset in particular, multitask LSTM models are widely adopted and competitive with newer architectures (Liu et al., 2023b).
-Given these findings, we benchmark against LSTM-SL as the SOTA/near-SOTA supervised model, and include transformer-only and vision-foundation encoders to disentangle temporal modeling from representation learning.
 Against these strong baselines, the soil moisture experiments confirm the superiority of the StefaLand-resConn architecture (Table 3).
-StefaLand-resConn achieves the best performance, with RMSE of 0.068 and correlation of 0.783 for random holdout, and maintains strong performance even in cross-continental testing on Europe (RMSE = 0.090, Corr = 0.638).
-The direct StefaLand model performs poorly since transformers alone lack an effective mechanism for temporal dependencies, while the resConn architecture compensates for this limitation.
-Similar to the streamflow results, TerraMind and PrithviWcX both showed no performance benefit compared to LSTM despite architectural adaptations, confirming both the challenge of repurposing vision foundation models across domains and the importance of pretraining with geoscience-relevant variables.
-Notably, the regional holdout on Europe demonstrates StefaLand-resConn’s superior spatial generalization capabilities, achieving a 25% correlation improvement over the LSTM baseline in this highly challenging extrapolation scenario.
+It achieves the best performance, with RMSE of 0.068 and correlation of 0.783 for random holdout, and maintains strong performance even in cross-continental testing on Europe (RMSE = 0.090, Corr = 0.638).
+The direct StefaLand model performs poorly because it lacks an effective mechanism to capture task-specific temporal dependencies.
+Similar to the streamflow cases above, TerraMind and PrithviWxC showed no performance benefit compared to LSTM, underscoring the challenge of repurposing vision foundation models across domains and the importance of pretraining with geoscience-relevant variables.
+The regional holdout on Europe further demonstrates StefaLand-resConn’s superior spatial generalization, achieving a 25% correlation improvement over the LSTM baseline in this difficult extrapolation scenario.
 We want to emphasize that TerraMind and PrithviWxC were not designed for the hydrologic and land–surface prediction tasks studied here.
-They excel in their intended domains of EO imagery and atmospheric variables, but are out-of-domain for dynamic land–surface and hydrologic modeling.
+They excel in their intended domains of Earth observation imagery and atmospheric variables, but are out-of-domain for dynamic land–surface and hydrologic modeling.
 We include them only to explore whether such models transfer any useful signal in our setting.
-Their weaker results should not be interpreted as evidence against their capability in other applications, but rather as a reflection of the mismatch between their pretraining objectives and the land–surface tasks we target.
-Nevertheless, we believe these benchmarks are helpful for clarifying their respective strengths, since the AI community may not be familiar with these datasets and models that have been claimed "any-to-any generative foundation model for Earth observation".
+These results should not be interpreted as evidence against their capability in other applications, but rather as a reflection of the mismatch between their pretraining objectives and the land–surface tasks we target.
+Nevertheless, we believe these benchmarks are helpful for clarifying their respective strengths, since the general AI community may not be familiar with these datasets and models that have been marketed as ‘any-to-any’ generative EO foundation models.
+Ultimately we view StefaLand as complementary to satellite-image foundation models, targeting a different but relevant data regime.
+We also could not identify other foundation models for land surface.
 3.4 Soil Property Prediction
-There are different soil datasets, each collected with different protocols and data processing techniques, resulting in significant discrepancies between datasets.
+There are different soil datasets, each collected with different protocols and data processing techniques, resulting in significant discrepancies.
 In this test, we finetuned StefaLand to predict in-situ soil profile data from another dataset (ISRIC).
 This application can produce a seamless dataset that is consistent with a set of in-situ data, improving data availability and addressing systematic biases.
-In addition, it helps us understand the noise associate with each dataset.
-StefaLand’s pretraining soils dataset is HWSD, which has some overlap but also quite extensive differences from ISRIC, which is larger and potentially noisier.
-We finetuned StefaLand to predict one soil texture property (e.g., clay percentage) in ISRIC while masking the corresponding complementary attributes (e.g., sand, silt) from the same profile to avoid information leakage.
-As a result, we are probing how easy it is to infer soil properties using other attributes such as climate, terrain and land cover.
-We compared StefaLand with and without pretraining against a supervised random forest baseline, which is prevalently used in the present geoscientific literature.
+In addition, it helps us understand the noise associated with each dataset.
+StefaLand’s pretraining soils dataset is HWSD, which has some overlap but also extensive differences from ISRIC, which is larger and potentially noisier.
+We finetuned StefaLand to predict one soil texture property (e.g., clay percentage) in ISRIC while masking the corresponding complementary attribute (e.g., sand) from the same profile to avoid information leakage, probing how easy it is to infer soil properties using other attributes such as climate, terrain and land cover.
+We compared StefaLand with and without pretraining against a supervised random forest baseline.
+Train-test splits can be found here C.2
 Table 4: In-situ soil property prediction using ISRIC WoSIS data.
 Model WoSIS Property Corr $\uparrow$ $R^{2}$ $\uparrow$
-StefaLand - Transformer Clay 0.20 0.04
-StefaLand - Finetune Clay 0.51 0.26
-Random Forest Clay 0.05 0.00
-StefaLand - Transformer Sand 0.25 0.06
-StefaLand - Finetune Sand 0.70 0.50
-Random Forest Sand 0.26 0.07
-StefaLand - Transformer Silt 0.41 0.17
-StefaLand - Finetune Silt 0.63 0.40
-Random Forest Silt 0.84 0.70
-StefaLand with finetuning achieved markedly higher predictive power, especially for sand fraction ($R^{2}=0.50$), compared to both direct training and the random forest baseline.
-These results highlight StefaLand’s ability to reconcile disparate soil datasets and improve the utility of noisy in-situ observations.
-At the same time, the weaker performance on silt suggests limits of transfer: silt fractions are often residually defined (sand and clay measured directly, silt inferred as the remainder), and thus more sensitive to error propagation across datasets.
-This makes them inherently harder to learn, as any noise or systematic difference between HWSD and ISRIC is amplified.
-The fact that the random forest attains higher accuracy on silt may also indicate that simpler models can more directly exploit the compositional closure relationship, whereas StefaLand relies more on cross-variable and environmental cues, which provide less direct signal for silt.
+StefaLand - direct Clay 0.197 0.038
+StefaLand - finetune Clay 0.509 0.259
+Random Forest Clay 0.456 0.207
+Linear Regression Clay 0.138 0.019
+StefaLand - direct Sand 0.253 0.064
+StefaLand - finetune Sand 0.704 0.495
+Random Forest Sand 0.585 0.342
+Linear Regression Sand 0.347 0.120
+StefaLand with finetuning achieved markedly higher predictive power for both clay and sand fractions, substantially outperforming direct training, random forest, and linear regression baselines.
+The other models struggle in this task, with random forest (RF) scoring noticeably lower.
+This task is not dynamical prediction, but a test of the models ability to build deep representations of the landscape that can be adapted to infer certain attributes using cross-variable group connections.
+StefaLand-finetune’s advantage against random forest suggests there are indeed deep representations uncovered by pretraining but not exploited by random forest.
+Finetuning StefaLand is also apparently superior to training from scratch as supervised learning (StefaLand - direct), again highlighting the value of pretraining.
+These results highlight StefaLand’s ability to reconcile disparate attribute datasets and improve the utility of noisy in-situ observations.
+3.5 Landslide Susceptibility Prediction
+Landslide is a geohazard that kill thousands each year.
+We next evaluated StefaLand for landslide susceptibility prediction using the SLIDO dataset from the State of Oregon, which provides detailed landslide occurrence records.
+Following Liu et al.
+(2025), this is a binary classification task indicating the presence or absence of landslides in a 30m by 30m patch.
+We finetuned StefaLand by extracting frozen hidden features and concatenating them with a 2D CNN, then retrained the CNN classifier to assess StefaLand’s ability to provide generalizable geoscience features.
+Table 5: Landslide susceptibility prediction results on the Oregon SLIDO dataset.
+Model Accuracy $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ ROC AUC $\uparrow$
+Logistic Regression 0.742 0.707 0.792 0.47 0.819
+Random Forest 0.751 0.703 0.834 0.763 0.839
+CNN2D 0.778 0.760 0.787 0.773 0.863
+StefaLand + CNN2D 0.799 0.791 0.793 0.792 0.875
+Note: All baseline results (Logistic Regression, Random Forest, CNN2D) are taken from previously published 30m-resolution experiments in Liu et al.
+(2025), except for StefaLand + CNN2D, which represents our proposed method.
+Results show that StefaLand’s pretrained features improved the CNN’s generalization, yielding modest gains across all metrics except Recall.
+This is a particularly difficult baseline to improve so even modest gains are rare.
+ROC AUC increased from 0.863 to 0.875, and precision rose from 0.760 to 0.791, reflecting fewer false positives.
+While random forest finds a high Recall, this is a particular realization with a large tradeoff on Precision.
+StefaLand produces overall well-rounded predictions, with both better Precision and Recall than CNN2D.
 4 Discussion
 4.1 Key Findings and Contributions
-As we review the literature, the methods to improve spatial generalization to data‐scarce regions are rare and rarely effective Beery et al.
-(2018), Gacu et al.
-(2025).
-In contrast, StefaLand, combined with lightweight fine‐tuning heads, achieves state-of-the-art or competitive performance across four broad problem classes: streamflow (both CAMELS and global), soil moisture, soil composition while also strengthening the parameterization of differentiable process-based models.
+As we review the literature, the methods to improve spatial generalization to data‐scarce regions are rare and rarely effective (Beery et al., 2018, Gacu et al., 2025).
+In contrast, StefaLand, combined with lightweight fine‐tuning heads, achieves state-of-the-art or competitive performance across four broad problem classes: streamflow (both CAMELS and global), soil moisture, soil composition, and landslide susceptibility, while also strengthening the parameterization of differentiable process-based models.
 Across tasks, the strongest gains come from architectures that fuse StefaLand embeddings with explicit temporal modeling via residual connections, indicating that pretraining on attribute-based spatiotemporal structure yields problem-relevant representations while temporal heads resolve sequence dynamics.
 These outcomes support the premise that foundation models can democratize prediction quality in data-scarce regions by improving out-of-domain transfer.
-The streamflow experiments, spanning basin-scale CAMELS PUB/PUR and global cross-continental runoff, consistently show StefaLand’s improved spatial generalization.
-The soil moisture results confirm robustness in both random site holdouts and continental transfer (Europe).
-For soil composition, StefaLand reconciles noisy WoSIS in-situ measurements with improved predictive power over random forests, highlighting its ability to harmonize disparate static datasets.
-An additional contribution is computational: our attribute-based approach is markedly more efficient than pixel-wise satellite transformers.
-The underlying transformer has far fewer parameters (roughly 12 million) and avoids the heavy data management requirements of image-centric pretraining.
-TerraMind’s larger configuration corresponds to about 7,680 GPU hours Jakubik et al.
-(2025), Aurora required roughly 14,592 GPU hours Bodnar et al.
-(2025), and AlphaEarth consumed on the order of 28,672 GPU hours Brown et al.
-(2025).
-These figures do not include the petabyte-scale data storage and transfer demands for satellite archives.
-By comparison, StefaLand’s attribute-based pretraining requires only about 720 GPU hours, making high-quality spatial generalization feasible on budgets accessible to academic groups.
-Finally, StefaLand is useful both operationally and as a representation layer.
-For hydrology, it supports point-scale prediction and regional products, for soil and geomorphic tasks it supplies frozen features that enhance conventional models.
-This dual role helps bridge a persistent gap between data-rich and data-poor regions by providing transferable geoscience embeddings that downstream models can reliably exploit.
-4.2 Limitations
-While StefaLand demonstrates strong generalization across diverse tasks, several limitations remain.
-The pretraining dataset is intentionally narrow and high quality, covering a carefully curated set of forcings and attributes.
-This focus improves signal extraction but may limit the model’s usefulness for problems outside or only loosely related to those variables, such as deeper subsurface processes like groundwater dynamics.
-Validation is also biased toward regions with dense observational coverage, such as the United States and Europe; although regional holdout tests suggest good transfer, true performance in data-scarce regions of Africa, Asia, and South America is less certain.
-4.3 Future Work
-Several directions remain to further develop StefaLand.
-Expanding the range of targets to include variables such as evapotranspiration, snow water equivalent, and groundwater levels would broaden its applicability, while harmonized global inventories could support more comprehensive evaluation of soil and landslide predictions across differing lithologies and climates.
-Methodologically, advances such as sequence-aware pretraining, uncertainty-aware prediction heads, and tighter integration with process models offer promising avenues to improve calibration and interpretability while preserving efficiency.
-Establishing standardized train–test protocols for random and regional holdouts, released with accompanying data splits, will also be important for enabling consistent benchmarking of future geoscience foundation models.
+The five dynamic prediction cases, along with benchmark models that reproduce state-of-the-art results in the literature, together paint a clear picture.
+The pretraining of StefaLand on attributes builds deep landscape representations.
+These features are highly relevant to hydrologic prediction tasks and can avoid overfitting compared to those built in task-specific supervised learning (LSTM or StefaLand-direct), improving model spatial generalization.
+They are also decidedly more relevant to such tasks than those obtained from existing Earth Foundation models expensively trained on massive amounts of satellite images.
+It shows that the importance of larger pretraining data may not necessarily exceed that of problem relevance, and image-like data may not be the optimal data representation for such tasks.
+We stress that our approach is complementary to satellite-based foundation models: where they exploit large-scale visual patterns, StefaLand focuses on problem-relevant attributes, offering an efficient and domain-specific alternative.
+We have tested a large number of adapter formulations.
+While we can certainly test more finetuning options, especially internal foundation model layers, the computational and data-storage demands already start to be limiting.
+Our attribute-based approach is at least an order of magnitude more efficient than pixel-wise satellite transformers.
+Our underlying transformer has far fewer parameters (roughly 12 million) and avoids the heavy data management requirements of image-centric pretraining.
+TerraMind’s larger configuration corresponds to about 7,680 GPU hours (Jakubik et al., 2025), PrithviWxC require roughly 23,040 GPU hours (Schmude et al., 2024) and Aurora required roughly 14,592 GPU hours (Bodnar et al., 2025).
+In contrast, StefaLand’s attribute-based pretraining requires only about 720 GPU hours, making high-quality spatial generalization feasible on limited budgets.
+These figures do not include the large gaps in data-storage and transfer demands during pretraining (StefaLand, Terramind and PrithviWxC used $\approx 2$, $\approx 11$ and $\approx 27$ data as mentioned earlier).
+In fact, it is costly and resource-straining to run the satellite-imagery-focused foundation models, which impedes us from comparing with them in every case.
+Finally, to the best of our effort, we could not identify other earth foundation models that are designed for land-surface predictions.
+4.2 Limitations and Future Work
+Several limitations remain.
+The selection of geological- or ecologically-focused attributes is limited and more can be added to further characterize the subsurface.
+Two-dimensional (or image-like) data like elevation map can be selectively incorporated using vision transformer heads in the future.
+Expanding the range of targets to include variables such as evapotranspiration, snow water equivalent, and groundwater levels would broaden its applicability.
+Methodologically, advances such as uncertainty-aware prediction heads, and tighter integration with additional process models offer promising avenues to improve calibration and interpretability while preserving efficiency.
 Overall, StefaLand shows that attribute-centric pretraining combined with lightweight temporal or physics heads can deliver strong spatial generalization across geoscientific tasks while remaining computationally accessible.
 This points toward a practical path for high-quality predictions in regions where they are most needed but data are most limited.
-5 Reproducibility Statement
-The pretrained StefaLand model and all code for both pretraining and finetuning is released publicly at [https://anonymous.4open.science/r/StefaLand-9421/].
-All datasets used in this work both pretraining dataset and all finetuneing benchmarks are fully public.
-A complete list of variables used for each task, along with their data sources, is provided in Appendix C.
-We also report all hyperparameters and model details in the same Appendix.
 Appendix A Detailed Model Architecture
-This appendix provides the complete mathematical formulation of the StefaLand model architecture.
+This appendix provides the complete mathematical formulation of the StefaLand model architecture more detailed then 2.
 A.1 Embedding Dynamic and Static Inputs
@@
 This static embedding acts as a global learnable token, allowing the model to incorporate basin-specific context into temporal dynamics at any depth of the Transformer layers.
-A.2 Transformer Encoder
+A.2 Location-aware Cross-Variable Group Masking
+StefaLand introduces Cross-Variable Group Masking (CVGM), a masking strategy that forces the model to capture interactions among correlated hydrologic variables rather than treating them independently.
+Given a predefinition of hydrological variables into groups $G=\{g_{1},g_{2},\ldots,g_{k}\}$, masking occurs as follows:
+  1. 1.
+    A temporal masking window $\tau,\tau+\ell)$ is randomly sampled with length $\ell\sim\mathcal{U}(L_{\text{min}},L_{\text{max}})$.
+  2. 2.
+    For each variable group $g_{k}$, a Bernoulli mask indicator $m_{k}\sim\text{Bernoulli}(p_{\text{mask}})$ determines whether the group is masked.
+  3. 3.
+    For each time step $t$ within the masked temporal window and each variable $c$ belonging to masked groups, the embedded feature vector is replaced by a learned mask vector $\mathbf{m}_{c}$.
+Each hydrologic variable $c$ has its own trainable mask embedding vector $\mathbf{m}_{c}\in\mathbb{R}^{256}$.
+This CVGM procedure creates reconstruction targets that require modeling cross-variable dependencies and physical interactions.
+A.3 Learnable Positional Encoding
+To provide positional information, StefaLand employs learnable positional encoding.
+Each position $i$, corresponding to each time step and the appended static embedding, is assigned a trainable embedding vector $\mathbf{p}_{i}$.
+The encoded embedding becomes:
+$\tilde{Z}=Z+P$ (20)
+where $P=[\mathbf{p}_{1};\ldots;\mathbf{p}_{T+1}]$.
+A.4 Transformer Encoder
 The embeddings enriched by positional encoding are processed through an $N$-layer Transformer encoder, where each Transformer block successively applies Multi-Head Self-Attention (MHA) with $h$ attention heads, followed by a residual connection and Layer Normalization.
@@
 $H^{(\ell)}$ $=\text{LayerNorm}(\tilde{H}^{(\ell)}+F^{(\ell)})$ (24)
-A.3 Reconstruction of Original Inputs
+A.5 Reconstruction of Original Inputs
 The final hidden states from the Transformer encoder, $H^{(N)}$, are linearly projected and passed through a single-layer bidirectional LSTM to capture the local temporal dependencies and continuity:
@@
 The projections leverage the learned latent representations to reconstruct the original hydrologic inputs.
-Appendix B Physics-Based Differentiable Modeling
+Appendix B Additional Experiments
+B.1 Physics-Based Differentiable Modeling
 To leverage domain knowledge and physical constraints inherent in hydrological systems, we implemented physics-based models that explicitly represent hydrological processes through mathematical formulations.
-These differentiable versions can be trained end-to-end within neural network frameworks, combining process understanding with machine learning flexibility Shen et al.
-(2023).
-For the process-based backbone, we employed the Hydrologiska Byråns Vattenbalansavdelning (HBV) model Aghakouchak and Habib (2010), Beck et al.
-(2020), Bergström (1976; 1992), Seibert and Vis (2012), a relatively simple bucket-type conceptual hydrologic model.
-HBV has state variables like snow storage, soil water, and subsurface storage, and can simulate flux variables like evapotranspiration (ET), recharge, surface runoff, shallow subsurface flow, and groundwater flow.
-The parameters of HBV are learned from basin characteristics by a deep learning network.
-We used an updated modern version, HBV1.1 Song et al.
-(2025), which includes modifications such as increased parallel storage components to represent heterogeneity within basins and dynamic parameterization capabilities.
-For physics-based configurations, we tested: (1) a baseline LSTM-HBV1.1 configuration as a standard reference, (2) StefaLand HBV1.1 with resConn, which combines the physics-based approach with our residual connection architecture, and (3) StefaLand HBV1.1 without resConn.
-These physics-based approaches incorporate hydrological process understanding while maintaining the ability to learn from data.
+These differentiable versions can be trained end-to-end within neural network frameworks, combining process understanding with machine learning flexibility (Shen et al., 2023).
+For the process-based backbone, we employed the Hydrologiska Byråns Vattenbalansavdelning (HBV) model (Aghakouchak and Habib, 2010, Beck et al., 2020, Bergström, 1976; 1992, Seibert and Vis, 2012), a relatively simple bucket-type conceptual hydrologic model.
+HBV has state variables like snow storage, soil water, and subsurface storage, and can simulate flux variables such as evapotranspiration (ET), recharge, surface runoff, shallow subsurface flow, and groundwater flow.
+We used an updated modern version, HBV1.1 (Song et al., 2025), which includes modifications such as increased parallel storage components to represent heterogeneity within basins and dynamic parameterization capabilities.
+The hybrid model employs a differentiable parameter learning (dPL) framework where neural networks generate parameters for HBV1.1, and errors are backpropagated through the entire system.
+A machine learning network takes basin attributes and meteorological forcings as inputs and outputs HBV parameters—both static (e.g., recession coefficients) and dynamic parameters that vary daily.
+Because HBV1.1 supports automatic differentiation, it serves as the physical backbone: during training, loss is calculated between simulated and observed streamflow, gradients are backpropagated through HBV equations, and neural network weights are updated.
+This differs from traditional calibration because parameters are learned regionally across all basins simultaneously rather than individually, allowing the network to capture generalizable relationships between basin characteristics and optimal parameters while maintaining mass balance constraints.
+The system uses 16 parallel response units for spatial heterogeneity and outputs diagnostic variables (e.g., evapotranspiration, soil moisture, baseflow) not directly trained on, providing interpretability with competitive performance.
+For physics-based configurations, we tested: (1) a baseline LSTM–HBV1.1 configuration as a standard reference, (2) StefaLand HBV1.1 with resConn, which combines the physics-based approach with our residual connection architecture, and (3) StefaLand HBV1.1 without resConn.
+These physics-based approaches incorporate hydrological process understanding while maintaining the ability to learn from data..
+Table B1: CAMELS Streamflow PUB and PUR Results (Physics-Based Models)
+Model Random holdout (ungauged basins) Regional holdout (ungauged regions)
+RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$ RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$
+LSTM - HBV1.1 1.325 1.298 0.857 0.672 1.561 1.521 0.746 0.578
+StefaLand - resConn HBV1.1 1.234 1.216 0.863 0.714 1.345 1.332 0.842 0.643
+StefaLand - no resConn HBV1.1 1.315 1.302 0.848 0.707 1.401 1.379 0.835 0.623
+StefaLand Ablation - resConn HBV1.1 1.310 1.306 0.842 0.693 1.465 1.432 0.607 0.512
+B.2 Linear Regression baselines
+To justify the use of complex neural networks over traditional methods, we have conducted baseline comparisons using linear regression models.
+As shown in the table below, linear regression performs poorly across all tasks by a fair margin when compared to our neural network approaches.
+Table B2: Additional experiments with linear regression baselines.
+Experiment Random holdout Regional holdout
+RMSE ↓ µbRMSE ↓ Corr ↑ RMSE ↓ µbRMSE ↓ Corr ↑
+Camels Streamflow Linear Regression 2.190 2.180 0.500 2.260 2.250 0.500
+Global Streamflow Linear Regression 1.823 1.746 0.252 1.816 1.721 0.248
+Soil Moisture Linear Regression 0.120 0.101 0.188 0.121 0.103 0.187
 Appendix C Experimental Details
+Table C1: StefaLand Pretraining Configuration
+Parameter Value
+General Settings
+Task pretrain
+Model MFFormer_dec_LSTM
+Random seed 111
+Time Period 1980/1/1–2018/12/31
+Sequence Configuration
+Sequence length 365
+Label length 365
+Prediction length 365
+Sampling stride 1
+Minimum window size 30
+Maximum window size 90
+Model Architecture
+Input dimension (enc_in) 32
+Decoder input (dec_in) 6
+Output dimension (c_out) 6
+Model dimension 256
+Number of heads 4
+Encoder layers 4
+Decoder layers 2
+Feed-forward dimension 512
+Dropout 0.1
+Activation gelu
+Training Configuration
+Optimizer AdamW
+Loss criterion MaskedNSE
+Epochs 25
+Batch size 256
+Learning rate 0.0001
+Weight decay 0.0
+Patience 30
+Gradient clipping 5.0
+Number of workers 10
+Loss Weights
+Time series loss ratio 1.0
+Static loss ratio 0.5
 Table C2: StefaLand Pretraining Variables and Sources
 Variable Type Variable Name Source
-Time Series Forcings Precipitation, Short-wave solar radiation downwards, Relative humidity, Maximum temperature, Minimum temperature, Potential evapotranspiration Multi-Source Weather (MSWX) and Multi-Source Weighted-Ensemble Precipitation (MSWEP)
-Static Attributes Forest cover fraction, grassland cover fraction Climate Change Initiative (CCI) land cover dataset
-Normalized Difference Vegetation Index (NDVI) Terra Moderate Resolution Imaging Spectroradiometer (MODIS) Vegetation Indices (MOD13A3)
-Sand, silt, clay fractions Harmonized World Soil Database (HWSD)
-Elevation, slope, aspect Global Multi-resolution Terrain Elevation Data (GMTED)
-Soil depth Global 1-km Gridded Thickness of Soil, Regolith, and Sedimentary Deposit Layers
-Carbonate sedimentary rock fraction Global Lithological Map (GLiM)
-Rock porosity, permeability GLobal HYdrogeology MaPS (GLHYMPS)
-Population density Gridded Population of the World (GPW) v4 dataset
-GDP per capita; population density Gross Domestic Product and Human Development Index over 1990-2015
-Forest intact fraction Intact Forest Landscapes Data
+Time Series Forcings Precipitation, Short-wave solar radiation downwards, Relative humidity, Maximum temperature, Minimum temperature, Potential evapotranspiration Multi-Source Weather (MSWX) and Multi-Source Weighted-Ensemble Precipitation (MSWEP) (Beck et al., 2022; 2019)
+Static Attributes Forest cover fraction, grassland cover fraction Climate Change Initiative (CCI) land cover dataset (ESA, 2017)
+Normalized Difference Vegetation Index (NDVI) Terra Moderate Resolution Imaging Spectroradiometer (MODIS) Vegetation Indices (MOD13A3) (Didan, 2015a)
+Sand, silt, clay fractions Harmonized World Soil Database (HWSD) (FAO et al., 2012)
+Elevation, slope, aspect Global Multi-resolution Terrain Elevation Data (GMTED) (Danielson and Gesch, 2011, Ramcharan et al., 2018)
+Soil depth Global 1-km Gridded Thickness of Soil, Regolith, and Sedimentary Deposit Layers (Pelletier et al., 2016)
+Carbonate sedimentary rock fraction Global Lithological Map (GLiM) (Hartmann and Moosdorf, 2012)
+Rock porosity, permeability GLobal HYdrogeology MaPS (GLHYMPS) (Gleeson et al., 2014)
+Population density Gridded Population of the World (GPW) v4 dataset (CIESIN, 2016)
+GDP per capita; population density Gross Domestic Product and Human Development Index over 1990-2015 (Kummu et al., 2018)
+Forest intact fraction Intact Forest Landscapes Data (Potapov et al., 2017)
 Outputs None (self-supervised pretraining)
+Table C3: Attribute Groups Used in Group Masking Pretraining
+Group Variables
+Topography meanelevation, meanslope
+Soil HWSD_clay, HWSD_sand, HWSD_silt, HWSD_gravel, SoilGrids1km_sand, SoilGrids1km_clay, SoilGrids1km_silt
+Geology permeability, Porosity, glaciers, permafrost
+Vegetation NDVI, FW
+Climate aridity, meanP, ETPOT_Hargr, meanTa, seasonality_P, seasonality_PET, snow_fraction, snowfall_fraction
 Table C4: CAMELS Streamflow HBV Model Hyperparameters
@@
 Variable Type Variable Name Source
-Time Series Forcings Precipitation, Temperature, Potential evapotranspiration, Solar radiation, Vapor pressure Catchment Attributes and Meteorology for Large-sample Studies (CAMELS)
+Time Series Forcings Precipitation, Temperature, Potential evapotranspiration, Solar radiation, Vapor pressure Catchment Attributes and Meteorology for Large-sample Studies (CAMELS) (Addor et al., 2017, Newman et al., 2014)
 Static Attributes Elevation, slope, catchment area, forest cover, LAI, GVF, soil depth, porosity, conductivity, sand, silt, clay fractions, carbonate fraction, permeability, aridity, snow fraction, precipitation extremes CAMELS
@@
 Variable Type Variable Name Source
-Time Series Forcings Albedo (BSA, WSA) Moderate Resolution Imaging Spectroradiometer (MODIS) MCD43A3 Version 6
-LST (Day, Night) MODIS Land Surface Temperature/Emissivity Daily (MYD11A1) Version 6.1
-Precipitation Global Precipitation Measurement (GPM), Multi-Source WeightedEnsemble Precipitation (MSWEP)
-Forecast albedo, LAI (high/low vegetation), soil temperature (layer 1), surface pressure, solar radiation, 2m temperature, evaporation, precipitation, U/V wind (10m) ERA5-Land Hourly - ECMWF Climate Reanalysis
-Static Attributes elevation, slope, aspect, roughness, curvature Global 1,5,10,100-km Topography database
-Sand, clay, silt, bulk density Harmonized world soil database (HWSD)
-Land cover; urban; open water; snow/ice ESA CCI land cover 2018
-NDVI Vegetation Indices Monthly L3 Global 0.05Deg CMG
-Outputs Soil moisture ESA CCI SM / in-situ merged
+Time Series Forcings Albedo (BSA, WSA) Moderate Resolution Imaging Spectroradiometer (MODIS) MCD43A3 version 6 (Schaaf, Crystal and Wang, Zhuosen, 2021)
+LST (Day, Night) MODIS Land Surface Temperature/Emissivity Daily (MYD11A1) Version 6.1 (Wan et al., 2021)
+Precipitation Global Precipitation Measurement (GPM), & MSWEP &ERA5 precipitation (Huffman et al., 2019, Beck et al., 2019, Muñoz Sabater, 2019)
+Forecast albedo, LAI (high/low vegetation), soil temperature (layer 1), surface pressure, solar radiation, 2 m temperature, evaporation, precipitation, U/V wind (10 m) ECMWF Reanalysis v5 (ERA5) (Muñoz Sabater, 2019)
+Static Attributes elevation, slope, aspect, roughness, curvature Global 1/5/10/100-km topography derivatives (Amatulli et al., 2018)
+Sand, clay, silt, bulk density HWSD v1.2 (FAO et al., 2012)
+Land cover; urban; open water; snow/ice ESA CCI Land Cover (ESA, 2017)
+NDVI Vegetation Indices Monthly L3 Global 0.05Deg CMG (Didan et al., 2015)
+Outputs Soil moisture International Soil Moisture Network (ISMN) (Dorigo et al., 2013b; 2011)
 Table C8: Global Streamflow Variables and Sources
 Variable Type Variable Name Source
-Time Series Forcings Precipitation, Temperature, Potential evapotranspiration, Radiation, Humidity MSWX and MSWEP
-Static Attributes Forest cover fraction, grassland cover fraction Climate Change Initiative (CCI) land cover dataset
-Normalized Difference Vegetation Index (NDVI) Terra Moderate Resolution Imaging Spectroradiometer (MODIS) Vegetation Indices (MOD13A3)
-Sand, silt, clay fractions Harmonized World Soil Database (HWSD)
-Elevation, slope, aspect Global Multi-resolution Terrain Elevation Data (GMTED)
-Soil depth Global 1-km Gridded Thickness of Soil, Regolith, and Sedimentary Deposit Layers
-Carbonate sedimentary rock fraction Global Lithological Map (GLiM)
-Rock porosity, permeability GLobal HYdrogeology MaPS (GLHYMPS)
-Population density Gridded Population of the World (GPW) v4 dataset
-GDP per capita; population density Gross Domestic Product and Human Development Index over 1990-2015
-Forest intact fraction Intact Forest Landscapes Data
-Outputs Streamflow Global gauge networks
+Time Series Forcings Precipitation, Temperature, Potential evapotranspiration, Radiation, Humidity MSWX and MSWEP (Beck et al., 2022; 2019)
+Static Attributes Forest cover fraction, grassland cover fraction Climate Change Initiative (CCI) land cover dataset (ESA, 2017)
+Normalized Difference Vegetation Index (NDVI) Terra Moderate Resolution Imaging Spectroradiometer (MODIS) Vegetation Indices (MOD13A3) (Didan, 2015a)
+Sand, silt, clay fractions Harmonized World Soil Database (HWSD) (FAO et al., 2012)
+Elevation, slope, aspect Global Multi-resolution Terrain Elevation Data (GMTED) (Danielson and Gesch, 2011, Ramcharan et al., 2018)
+Soil depth Global 1-km Gridded Thickness of Soil, Regolith, and Sedimentary Deposit Layers (Pelletier et al., 2016)
+Carbonate sedimentary rock fraction Global Lithological Map (GLiM) (Hartmann and Moosdorf, 2012)
+Rock porosity, permeability GLobal HYdrogeology MaPS (GLHYMPS) (Gleeson et al., 2014)
+Population density Gridded Population of the World (GPW) v4 dataset (CIESIN, 2016)
+GDP per capita; population density Gross Domestic Product and Human Development Index over 1990-2015 (Kummu et al., 2018)
+Forest intact fraction Intact Forest Landscapes Data (Potapov et al., 2017)
+Outputs Streamflow Global runoff data (GRDC, 2024)
+Table C9: Landslide (SLIDO, Oregon) Variables and Sources
+Variable Type Variable Name Source
+Input data Elevation National Elevation Dataset (NED) (Gesch et al., 2018)
+Soil sand, silt, clay, bulk density, saturated hydraulic conductivity Probabilistic Remapping of SSURGO (POLARIS) (Chaney et al., 2019)
+Lithology Global Lithological Map (GLiM) (har)
+Rainfall PRISM (PRISM Climate Group, 2014)
+NDVI Moderate Resolution Imaging Spectroradiometer (MODIS) Vegetation Indices Monthly L3 (Didan, 2015b)
+Landcover National Land Cover Database (NLCD) 2016 (Dewitz, 2019)
+Soil moisture SMAP-HydroBlocks (SMAP-HB) (Vergopolan et al., 2021)
+slope, aspect, curvature, TWI, SPI DEM-derived
+Outputs Landslide occurrence (binary) Statewide Landslide Information Database for Oregon (SLIDO) (Franczyk, J. J et al., 2020)
 Table C10: Soil Composition (ISRIC) Variables and Sources
@@
 Static Attributes Same as Table 7
-Outputs Soil property (clay; sand; silt) The World Soil Information Service (WoSIS)
+Outputs Soil property (clay; sand; silt) World Soil Information Service (WoSIS) (Batjes et al., 2020)
+Table C11: PrithviWxC Surface Variables and Sources (MERRA-2)
+Variable Type Variable Name Source
+Time Series Forcings Precipitation, fluxes, winds, soil moisture/temperature, LAI, runoff, etc. NASA MERRA-2 reanalysis (Gelaro et al., 2017)
+Static Attributes Land/ocean/ice fractions, surface geopotential, subgrid orography NASA MERRA-2 constants (Gelaro et al., 2017)
 Table C12: Computation Resources for StefaLand and Comparison Models
@@
   • Note: All values except pretraining are for the CAMELS benchmark experiment. The relative differences in computational requirements are consistent across other experiments.
+C.1 Pretraining Data Handling
+To handle large volumes efficiently, data were stored in shards rather than fully loaded into memory.
+At the start of each epoch, shards were randomly reindexed, providing diverse samples while avoiding costly materialization of the full dataset.
+This strategy ensures both reproducibility and wide sample coverage.
+The accompanying code release in the reproducibility statement enables reconstruction of this dataset and reproduces the full pretraining pipeline.
+C.2 Dataset Splitting
+For the WoSIS soil dataset, we collected soil property data from 106,503 locations.
+After removing low-quality records (e.g., sand values greater than 1 or negative values), we randomly sampled 5,000 soil points to reduce computational cost.
+We then applied 5-fold cross-validation (k=5) on this subset.
+For the landslide dataset at 30 m resolution, we used 14,604 historical landslide points. We split the dataset into 70% for training, 20% for validation, and 10% for testing.
 Appendix D Model Architectures
@@
 For time series with missing values, only timestamps where both predicted and target values were available were used in metric calculations.
-Appendix F Additional Model results
-Table F1: Additional experiments with linear regression baselines.
-Experiment Random holdout Regional holdout
-RMSE ↓ µbRMSE ↓ Corr ↑ RMSE ↓ µbRMSE ↓ Corr ↑
-Camels Streamflow Linear Regression 2.190 2.180 0.500 2.260 2.250 0.500
-Global Streamflow Linear Regression 1.823 1.746 0.252 1.816 1.721 0.248
-Soil Moisture Linear Regression 0.120 0.101 0.188 0.121 0.103 0.187

### v3

@@
 Abstract
-Stewarding natural resources, mitigating floods, droughts, wildfires, and landslides, and meeting growing demands require models that can predict climate-driven land-surface responses and human feedback with high accuracy.
-Traditional impact models, whether process-based, statistical, or machine learning, struggle with spatial generalization due to limited observations and concept drift.
-Recently proposed vision foundation models trained on satellite imagery demand massive compute and are ill-suited for dynamic land-surface prediction.
-We introduce StefaLand, a generative spatiotemporal earth foundation model centered on landscape interactions.
-StefaLand improves predictions on four tasks and five datasets: streamflow, soil moisture, soil composition, and landslide detection compared to prior state-of-the-art.
-Results highlight its ability to generalize across diverse, data-scarce regions and support broad land-surface applications.
-The model builds on a masked autoencoder backbone that learns deep joint representations of landscape attributes, with a location-aware architecture fusing static and time-series inputs, attribute-based representations that drastically reduce compute, and residual fine-tuning adapters that enhance transfer.
-While inspired by prior methods, their alignment with geoscience and integration in one model enables robust performance on dynamic land–surface tasks.
-StefaLand can be pretrained and finetuned on academic compute yet outperforms state-of-the-art baselines and even fine-tuned vision foundation models.
-To our knowledge, this is the first geoscience land-surface foundation model that demonstrably improves dynamic land-surface interaction predictions and supports diverse downstream applications.
+Managing natural resources and mitigating risks from floods, droughts, wildfires, and landslides require models that can accurately predict climate-driven land–surface responses.
+Traditional models often struggle with spatial generalization because they are trained/calibrated on limited observations and can degrade under concept drift.
+Recently proposed vision foundation models trained on satellite imagery demand massive compute, and they are not designed for dynamic land surface prediction tasks.
+We introduce StefaLand, a generative spatiotemporal Earth representation learning model centered on learning cross-domain interactions to suppress overfitting.
+StefaLand demonstrates especially strong spatial generalization on five datasets across four important tasks: streamflow, soil moisture, soil composition and landslides, compared to previous state-of-the-art methods.
+The domain-inspired design choices include a location-aware masked autoencoder that fuses static and time-series inputs, an attribute-based rather than image-based representation that drastically reduces compute demands, and residual fine-tuning adapters that strengthen knowledge transfer across tasks.
+StefaLand can be pretrained and finetuned on commonly-available academic compute resources, yet consistently outperforms state-of-the-art supervised learning baselines, fine-tuned vision foundation models and commercially-available embeddings, highlighting the previously overlooked value of cross-domain interactions and providing assistance to data-poor regions of the world.
+Machine Learning, Earth System Science, Hydrology, Land Surface Modeling
 1 Introduction
-Climate change is ushering in strong and widespread changes on the land surface, including higher frequencies of floods, droughts, wildfires and other geohazards (Ebi et al., 2021, IPCC, 2021).
+Climate change is ushering in strong and widespread changes on the land surface, including higher frequencies of floods, droughts, wildfires and other geohazards (Ebi et al., 2021; IPCC, 2021).
 To mitigate the impact of these disasters, there are urgent needs for models that can accurately predict land surface dynamics such as streamflow, soil moisture, soil composition, landslides, snow water equivalent, groundwater levels, and vegetation carbon content.
-Among these, soil moisture controls the partition of rainfall into infiltration and runoff, modulates flood generation and landslides, and critically influences land-atmosphere interactions (Dorigo et al., 2013a).
+Among these, soil moisture controls ecosystem health and influences land-atmosphere interactions (Dorigo et al., 2013a).
 Streamflow is the flow rate of water running in the rivers, the most accessible water resource to humans, and too high or too low streamflow can cause flooding or hydrologic drought, respectively.
@@
 Here, we limit our scope to the predictions of dynamical or static land surface processes that represent the impacts of climate change.
-Predicting these dynamic variables is distinct from image-recognition tasks, as here we seek to predict what will happen in the near or distant future.
 Traditionally, these tasks were undertaken by physics-based models that take atmospheric forcings (precipitation, temperature) as inputs and sequentially calculate the physical processes that eventually lead to the variables of interest (Li et al., 2015).
@@
 However, up to now, most of the geoscientific ML models have been supervised ML approaches trained specifically for a narrow set of tasks.
-A particularly notable gap is that currently established geoscience foundation models have largely been trained on satellite imagery for landcover-identification tasks (Jakubik et al., 2023; 2025, Schmude et al., 2024), which limits their ability to capture dynamic land–atmosphere interactions.
-As a result, valuable temporal datasets and ground-based observations remain underutilized (Xie et al., 2023), and to our knowledge no foundation model has yet been developed with a primary focus on dynamical land surface modeling.
+Large vision foundation models have been trained on satellite imagery of earth to facilitate mapping (Tseng et al., 2025), weather prediction (Schmude et al., 2024), air quality (Bodnar et al., 2025) and other geoscientific tasks (Jakubik et al., 2023, 2025).
+A particularly notable observation is that they have not focused on the interactions among landscape domains (climate, soil, vegetation, terrain, geology).
+The catchment coevolution hypothesis (Troch et al., 2015) states that terrain, soil, vegetation and climate all coevolve, shaping the landscape we have today.
+Knowing parts of the landscape domains often allow us to predict the others.
+This implies that their joint distributions can greatly inform latent processes relevant to the transport of water and materials in the catchments.
+However, valuable temporal datasets and ground-based observations remain underutilized (Xie et al., 2023), and to our knowledge, no foundation model has yet been developed with a primary focus on dynamical land surface modeling.
 A Grand Challenge for geoscientific ML models is to improve their spatial generalization, because a frequent issue facing them is the sparsity and spatial imbalance of observational data.
 Satellite data are often coarse in resolution and uncertain compared to in-situ measurements.
-SMAP, for example, provides global soil moisture observations at  9–36 km resolution every 2–3 days, which are useful for regional climate and hydrologic research but far less valuable than in-situ probes for operational field tasks such as irrigation scheduling or crop stress monitoring (Entekhabi et al., 2010).
+SMAP, for example, provides global soil moisture observations at  9–36 km resolution every 2–3 days, which are useful for regional climate and hydrologic research but far less valuable than in-situ probes for operational field tasks such as irrigation scheduling or crop stress monitoring (Entekhabi et al., 2010; Liu et al., 2022, 2023b).
 However, in-situ data, due to the cost of installing instruments and varying policies on data sharing, is only available in high density in certain regions of some developed nations.
 For example, streamflow gauge data are abundant in the United States, Europe, Australia and Japan, but remain sparse in Africa, South America, and much of Asia (Global Runoff Data Centre, 2020).
-Similar distribution patterns are found for high-quality in-situ soil moisture probes and soil property measurements.
 As quantified in many studies (Feng et al., 2023), a deep network trained on data from some regions can face substantial performance degradation when applied in data-scarce regions.
 This occurs partly because there are not enough sites to learn the true dependencies of the targets on static land surface characteristics, and partly because of systematic data discrepancies across regions (concept drift).
-While such limitations hinder traditional supervised ML models, foundation models offer a potential path forward: by jointly learning from broad, heterogeneous datasets (including temporal records and ground observations where available), they may transfer useful representations to data-scarce regions where task-specific training data are limited.
+While such limitations hinder traditional supervised ML models, foundation or representation learning models offer a potential path forward: by jointly learning from broad, heterogeneous datasets (including temporal records and ground observations where available), they may transfer useful representations to data-scarce regions where task-specific training data are limited.
 Related Work: In hydrologic and ecosystem predictions, supervised long short-term memory (LSTM) networks (Hochreiter and Schmidhuber, 1997) remain a highly popular architecture, in part because land surface processes often behave like Markov processes where LSTMs’ gating mechanisms handle noisy continuous inputs well (Kratzert et al., 2018).
-Attempts to adopt transformers, so successful in natural language processing, have generally found it difficult to noticeably surpass LSTM in time series regression tasks (Xue et al., 2023, Liu et al., 2024), with evidence of overfitting on continuous signals (Zeng et al., 2022).
-Nonetheless, recent studies show that with task-specific modifications and careful fine-tuning, transformers can achieve competitive results in extreme event prediction (Wen et al., 2023), precisely the areas where current hydrologic models struggle most with spatial generalization.
-Traditional hydrologic research on "prediction in ungauged basins" (PUB) have examined regionalization and spatial interpolation approaches including clustering or classifying catchments and transferring parameters from donor catchments in the same class (Hrachowitz et al., 2013, Yang et al., 2023).
+Attempts to adopt transformers, so successful in natural language processing, have generally found it difficult to noticeably surpass LSTM in time series regression tasks (Xue et al., 2023; Liu et al., 2024, 2025b), with evidence of overfitting on continuous signals (Zeng et al., 2022).
+Nonetheless, recent studies show that with task-specific modifications and careful fine-tuning, transformers can achieve competitive results in extreme event prediction (Wen et al., 2023), precisely the areas where current hydrologic models struggle most with.
+Traditional hydrologic research on ”prediction in ungauged basins” (PUB) have examined regionalization and spatial interpolation approaches including clustering or classifying catchments and transferring parameters from donor catchments in the same class (Hrachowitz et al., 2013; Yang et al., 2023).
 Such an expert-derived design represents a crude practice of unsupervised learning that indicates the importance of understanding the joint data distribution.
 However, modern weakly-supervised foundation models can, in general, much better grasp the joint data distribution than expert-driven approaches.
-Foundation models offer a promising approach to address these spatial generalization challenges.
+Representation learning models offer a promising approach to address these spatial generalization challenges.
 By pretraining on large-scale datasets to learn generalizable representations, these models can potentially transfer knowledge across regions and geoscientific domains (Zhang et al., 2024).
-Existing geoscience foundation models, e.g., TerraMind (Jakubik et al., 2025) and Prithvi (Hsu et al., 2024) from IBM, and Aurora (Bodnar et al., 2025) from Microsoft, have largely focused on satellite imagery, which may not capture the temporal dynamics and physical processes most relevant to land surface predictions.
-For example, TerraMind is trained on 9 million globally distributed satellite imagery samples across various modalities (optical, radar, elevation, land use) to create a generative multimodal foundation model for Earth observation (Jakubik et al., 2023).
-It has been claimed to serve as "any-to-any generative, multimodal foundation model for Earth observation" (Jakubik et al., 2025).
-However, many critical variables regarding soil and the subsurface are not directly observable from space.
-In addition, satellite images are noisy and a large portion of the data is redundant across repeated revisits or irrelevant for hydrologic processes.
-Therefore, it is unclear whether they are of high relevance to land-surface dynamical prediction tasks.
-Our contributions: We present Spatial-Temporal Earth Foundation model with Attributes for the Land Surface (StefaLand), the first land-focused geoscientific foundation model designed for dynamic land–surface prediction.
-StefaLand improves predictions on streamflow, soil moisture, soil composition, and landslide susceptibility compared to state-of-the-art supervised learning baselines as well as finetuned satellite-image-trained earth foundation models.
-Especially, it shows strong spatial generalization across diverse landscapes and data-scarce regions for a wide variety of tasks.
+Recent progress in geoscientific foundation models has been driven by the increasing availability of global Earth system datasets and advances in self-supervised learning, enabling pretrained representations that transfer across tasks and regions (Bommasani et al., 2022; Lacoste et al., 2023).
+Beyond vision-based approaches, recent work on structured data has shown that foundation models such as TabPFN can learn transferable priors over tabular regression tasks without task-specific training.
+(Hollmann et al., 2025).
+Most existing Earth observation foundation models, including TerraMind (Jakubik et al., 2025), Prithvi (Hsu et al., 2024), Aurora (Bodnar et al., 2025), AlphaEarth (Brown et al., 2025), and Galileo (Tseng et al., 2025), are pretrained primarily on remote sensing imagery and related EO products.
+These models are effective at capturing surface appearance and spatial patterns, and their representations often act as static or slowly varying landscape descriptors.
+In particular, AlphaEarth learns global, high-resolution spatial embeddings from large-scale Earth observation data that capture consistent landscape structure across climate and biome gradients, and these embeddings have been shown to transfer effectively as fixed spatial features for a wide range of downstream Earth system prediction tasks, including hydrology.
+Because AlphaEarth embeddings are globally available and independent of task-specific labels, they provide a strong and widely applicable baseline for evaluating spatial generalization in data-limited settings.
+However, many variables central to land-surface and hydrologic dynamics, including subsurface properties, storage processes, and long-term temporal interactions, are not directly observable from space and therefore cannot be reliably inferred from imagery alone.
+Moreover, EO foundation models are typically optimized for surface-level semantic consistency and invariance to transient atmospheric or observational effects, which can further suppress signals that are informative for hydrologic processes, such as moisture-related variability or persistence.
+Our contributions: We present the Spatial-Temporal Earth Foundation model with Attributes for the Land Surface (StefaLand), a land-focused geoscientific representation learning model for dynamic land–surface prediction.
+StefaLand is evaluated on streamflow, soil moisture, soil composition, and landslide susceptibility under spatial holdout regimes.
+It shows strong spatial generalization across diverse landscapes and data-scarce regions for a wide variety of tasks.
 StefaLand’s attribute-based rather than image-based design (with the potential to link to image-like inputs in the future) incorporates a variety of ground-based measurement data, emphasizes relevant land-surface physical processes, drastically reduces compute requirements while retaining global coverage, making it accessible to researchers with modest resources.
 Pretraining our model required only about 720 V100 GPU hours (could be shorter with more advanced GPUs).
-The model builds on a masked autoencoder backbone, a location-aware fusion of static and time-series inputs, grouped masking to promote cross-attribute interactions, and residual fine-tuning adapters, into a coherent design guided by geoscientific knowledge.
+The model builds on a masked autoencoder backbone, a location-aware fusion of static and time-series inputs, grouped masking to promote cross-domain interactions, and residual fine-tuning adapters, into a coherent design guided by geoscientific knowledge.
 Taken together, these contributions establish StefaLand as an efficient and accessible complement to vision-based foundation models.
+Figure 1: Conceptual overview of the StefaLand Structure.
+Static landscape attributes and dynamic forcings are jointly embedded using a transformer-based masked autoencoder with cross-variable group masking.
+With relevant dimensionality included.
 2 Methods
@@
 StefaLand addresses this challenge with a transformer-based masked autoencoder that jointly embeds static and dynamic variables, pretrained with a cross-variable masking strategy, and then adapted for specific prediction tasks with task-specific heads.
-2.1 StefaLand Transformer Architecture
-Embedding.
-Each dynamic input variable $c$ at time $t$ is independently embedded with a nonlinear two-layer projection:
-$z_{t,c}=\text{GELU}(x_{t,c}W_{1,c}+b_{1,c})W_{2,c}+b_{2,c},$ (1)
-where $W_{1,c}\in\mathbb{R}^{1\times 64}$ and $W_{2,c}\in\mathbb{R}^{64\times 256}$. Summing across all $C$ dynamic variables yields a per-step embedding:
-$z_{t}=\sum_{c=1}^{C}z_{t,c}.$ (2)
-Static attributes $s_{i}$ are embedded similarly:
-$z_{\text{static},i}=\text{GELU}(s_{i}W_{1,i}+b_{1,i})W_{2,i}+b_{2,i},$ (3)
-and concatenated as a special static token, producing the joint sequence:
-$Z=[z_{1};z_{2};\ldots;z_{T};z_{\text{static}}].$ (4)
-Cross-Variable Group Masking (CVGM).
-Let $G=\{g_{1},\ldots,g_{K}\}$ be groups of related variables. During pretraining, a temporal window $\tau,\tau+\ell)$ is sampled, and entire groups are stochastically masked:
-$m_{k}\sim\text{Bernoulli}(p_{\text{mask}}).$ (5)
-For all $c\in g_{k}$ with $m_{k}=1$, the embedding is replaced by a learned mask vector $\mathbf{m}_{c}\in\mathbb{R}^{256}$.
-This forces reconstructions to rely on cross-variable dependencies rather than treating drivers independently.
-Transformer encoder.
-After adding learnable positional encodings $P$, we obtain:
-$\tilde{Z}=Z+P.$ (6)
-This sequence is processed by $N$ Transformer blocks:
-$A^{(\ell)}$ $=\text{MHA}(H^{(\ell-1)}),$ (7)
-$\tilde{H}^{(\ell)}$ $=\text{LayerNorm}(H^{(\ell-1)}+A^{(\ell)}),$ (8)
-[…]
-$H^{(\ell)}$ $=\text{LayerNorm}(\tilde{H}^{(\ell)}+F^{(\ell)}).$ (10)
-Decoder and loss.
-The final hidden states are passed through a bidirectional LSTM to refine local temporal continuity:
-$U=\text{LSTM}(H^{(N)}W_{\text{proj}}+b_{\text{proj}}).$ (11)
-Outputs corresponding to masked groups are projected back to the original variable dimensions. The reconstruction loss is
-$\mathcal{L}=\sum_{c\in\mathcal{M}}w_{c}\frac{\|\hat{x}_{c}-x_{c}\|_{2}^{2}}{\sigma_{c}^{2}},$ (12)
-where $\mathcal{M}$ is the set of masked variables, $\sigma_{c}$ is the standard deviation of variable $c$, and $w_{c}$ is a learnable weight.
+2.1 Stefaland Structure
+StefaLand is a transformer-based masked autoencoder inspired by bidirectional language models such as BERT (Devlin et al., 2018), designed to jointly embed static landscape attributes and dynamic time-varying forcings within a unified representation.
+During pretraining, StefaLand randomly masks subsets of the input variables and learns to reconstruct the masked components using the remaining unmasked information.
+Masking may affect static attributes, temporal variables, or entire variable groups, encouraging the model to learn the joint distribution over heterogeneous land-surface controls rather than relying on any single data source.
+A key design choice in StefaLand is cross-variable group masking, in which physically or statistically related variables are masked together rather than independently.
+This prevents trivial reconstruction through correlated inputs and encourages the model to capture cross-domain linkages, such as interactions between soil texture and climate seasonality or between topography and hydrologic response.
+Grouped masking strategies have previously been explored in multimodal learning settings, including Presto (Tseng et al., 2023), and here we adapt this idea using a reconstruction loss applied to masked inputs, normalized by variable-wise standard deviations when available.
+The attention mechanism naturally accommodates missing inputs by suppressing attention weights for masked tokens, allowing StefaLand to flexibly reason over arbitrarily incomplete input configurations.
+Although more recent masked-model formulations exist, we adopt a BERT-style design for its simplicity, interpretability, and robustness across heterogeneous inputs.
+Following pretraining, the encoder produces contextualized embeddings for both static attributes and dynamic forcings, which are reused across downstream tasks through lightweight task-specific heads and residual adaptation pathways, as illustrated in Figure 1.
 2.2 Pretraining Details
-The pretraining dataset is a derived global attribute dataset spanning $\sim$8,000 locations (basins) over 40 years.
+The pretraining dataset is a derived global attribute dataset spanning $\sim$8,634 locations (basins) over 40 years.
 Variables were chosen to represent the key controls on fluxes of water, energy, momentum, sediment, and nutrients.
 A complete list of variables, their group assignments, and their sources is provided in Appendix  C.
-The CVGM is designed so variables with reciprocal or bidirectional causality are masked together, preventing them from acting as predictors for each other.
-Most groupings are straightforward, such as masking silt and clay percentages together.
-One exception is soil depth, which we classify as a terrain attribute since it is strongly tied to topographic derivations.
-By masking and reconstructing at the group level, the model is encouraged to capture cross-domain interactions, such as the coupling between soil texture and climate seasonality.
-While the grouped-masking strategy has sometimes been employed in multimodality AI models, we did not find it applied in earth foundation models like Terramind, Prithvi, or Aurora, which instead use random dropout, spatial patches, or future periods for masking.
-Our objective is a reconstruction loss on the masked slice of the sequence, normalized by variable-wise standard deviations when available and with per-variable weighting via a learnable weight vector.
+A visualization of the pretraining dataset coverage is provided in Appendix C Figure  4.
+The cross-variable group masking (CVGM) scheme is designed so that variables with reciprocal or bidirectional relationships are masked together, preventing them from acting as direct predictors for one another.
+Most groupings are straightforward and common sense, such as masking silt and clay fractions jointly, while a few variables require domain-specific treatment; for example, soil depth is grouped with terrain attributes due to its strong dependence on topographic derivations.
+The complete set of variables and their group assignments is provided in Table  16.
+By masking and reconstructing variables at the group level, the model is encouraged to capture cross-domain interactions, such as couplings between soil texture and climate seasonality.
 2.3 Finetuning for Prediction Tasks
-Our primary finetuning model, StefaLand-resConn, integrates pretrained embeddings with raw forcings through a residual pathway (Figure D1).
-Let $E_{t}$ denote the transformer embedding at time $t$, and $x_{t}$ the raw forcings.
+We adapt the pretrained StefaLand encoder to downstream prediction tasks using lightweight task-specific heads while keeping the main model weights frozen.
+This strategy, illustrated in Figure 1, preserves the general-purpose representations learned during pretraining and reduces the risk of overfitting when finetuning on limited task-specific data.
+Our primary finetuning configuration, StefaLand-resConn, integrates pretrained embeddings with raw meteorological forcings through a residual adaptation pathway via additive fusion prior to the recurrent decoder.
+Let $E_{t}$ denote the StefaLand embedding at time $t$, and let $x_{t}$ denote the corresponding raw forcing inputs.
 We compute:
@@
 $h_{t}$ $=\text{LSTM}(E_{t}+r_{t}),$ (2)
-$\hat{y}_{t}$ $=W_{o}h_{t}+b_{o}.$ (15)
-Skip connections propagate $E_{t}$ and $r_{t}$ into intermediate layers, allowing general knowledge (from $E_{t}$) to be iteratively refined with task-specific signals ($r_{t}$).
-This design strengthens spatial generalization while adapting to local variability.
-To see a visual representation of this D1.
-2.4 Foundation model comparisons
-For completeness, we also evaluated two existing Earth-Observation-oriented foundation models, TerraMind and PrithviWxC (Jakubik et al., 2025, Hsu et al., 2024), using the same fine-tuning heads and training protocols as StefaLand.
-We coupled them together with our residual connection architecture and finetuned the added units while keeping the foundation models’ weights frozen.
-Because both TerraMind and PrithviWxC require orders of magnitude larger storage space, data transfer, and fine-tuning costs than StefaLand, we chose comparison tasks selectively, focusing on cases where their pretraining data was potentially most relevant.
-Accordingly, we compared both models on soil moisture, while streamflow was included only in a more exploratory capacity (only TerraMind).
-For a rough comparison, StefaLand, TerraMind and PrithviWxC used respectively $\approx 2$, $\approx 11$ and $\approx 27$ TB data during pretraining.
-Due to the intensive data requirements for PrithviWxC, only surface-level variables likely relevant to land surface interactions, along with all 14 static variables, were used.
-The full atmospheric variables at differing pressure levels were excluded.
+$\hat{y}_{t}$ $=W_{o}h_{t}+b_{o},$ (3)
+where $f_{\text{conv+linear}}(\cdot)$ denotes a shallow convolutional and linear projection block.
+Residual connections propagate both the pretrained embedding $E_{t}$ and the task-specific signal $r_{t}$ into the recurrent decoder, allowing general spatial knowledge captured during pretraining to be iteratively refined using task-specific temporal information.
+This design strengthens spatial generalization while retaining flexibility to adapt to local dynamics.
+We employ an LSTM decoder for tasks with explicit temporal structure, such as streamflow and soil moisture prediction.
+For non-temporal or spatially structured tasks, including soil property inference and landslide susceptibility mapping, the temporal decoder is replaced with a task-appropriate adapter head, such as a multilayer perceptron or two-dimensional convolutional neural network (CNN2D).
+In these cases, frozen StefaLand embeddings are used as contextual features.
+Across all tasks, only the adapter modules and task heads are updated during finetuning, while the pretrained StefaLand encoder remains frozen, ensuring computational efficiency and stable transfer from the representation learning model.
 3 Experiments
-We tested the value of foundation model pretraining on 5 datasets and 6 experiments, including, streamflow on the CAMELS dataset on USA, CAMELS streamflow prediction with hybrid model, global streamflow, global in-situ soil moisture, global soil properties, and landslide susceptibility in Oregon, USA.
+We tested the value of foundation model pretraining on 5 datasets and 6 experiments, including, streamflow on the CAMELS dataset on USA (a widely used benchmark dataset in hydrology), CAMELS streamflow prediction with hybrid model, global streamflow on Caravan, global in-situ soil moisture, global soil properties, and landslide susceptibility in Oregon, USA.
 For all experiments, hyperparameters were tuned with Ray Tune and kept consistent across model configurations within each experimental case (e.g., CAMELS streamflow, soil moisture, etc).
@@
 Complete details of hyperparameters, forcings, and static features for all four experiments are in Appendix C.
-Model variants used.
-Unless noted otherwise, pretrained encoders are frozen and only task heads are trained.
-We evaluate: (i) LSTM-SL: supervised LSTM baseline; (ii) StefaLand-direct: the StefaLand encoder together with its original decoder trained directly on the task data, without pretraining; (iii) StefaLand-resConn: pretrained encoder with residual pathway + LSTM decoder, our proposed architecture; (iv) StefaLand-noResConn: pretrained encoder with a simple adapter (no residual/LSTM); (v) StefaLand-scratch: same resConn head but StefaLand initialized randomly (no pretraining) and unfrozen (ablation test); (vi) TerraMind-resConn and PrithviWxC-resConn: EO/atmospheric foundation encoders frozen with the same residual head.
+Model variants and baselines.
+We evaluate StefaLand alongside established baselines commonly used in hydrologic and geophysical time series modeling.
+As supervised baselines, we include LSTM-SL, which remains the dominant architecture for streamflow prediction and follows the same core model class used in prior large scale hydrology studies (Kratzert et al., 2019; Feng et al., 2021; Sabzipour et al., 2023).
+We additionally evaluate Informer (Zhou et al., 2021), Reformer (Kitaev et al., 2020), and DLinear (Zeng et al., 2022), which, although originally proposed for forecasting, are applied here in a sequence to sequence setting and are commonly included in recent hydrologic and geophysical modeling studies as representative transformer-based baselines.
+To assess pretrained Earth representations, we include AlphaEarth LSTM and AlphaEarth ResConn, which feed pretrained AlphaEarth embeddings into an LSTM decoder with and without a residual task adapter, respectively (Brown et al., 2025).
+Because our task periods do not fully overlap AlphaEarth’s 2017–2024 temporal coverage, AlphaEarth embeddings are spatially aggregated over each basin and collapsed into a fixed set of 64 static features per site.
+For CAMELS streamflow and soil moisture, we additionally evaluate TabPFN, a tabular foundation model (Hollmann et al., 2025).
+Since TabPFN is non sequential and constrained by a fixed context window, we provide it with compact statistical summaries of each historical window while preserving identical data splits and evaluation protocols.
+We also conducted an exploratory evaluation of additional Earth observation and atmospheric foundation models to assess their viability for land–surface prediction under matched adaptation strategies; these limited-scope results are reported in B.3
+Our proposed method, StefaLand resConn (residual connection), combines a pretrained encoder with a residual adapter and LSTM decoder.
 3.1 CAMELS Streamflow Prediction
-To compare spatial generalization on a well benchmarked dataset, we follow (Feng et al., 2021), testing prediction in ungauged basins (PUB) and ungauged regions (PUR).
-These correspond to randomized spatial K-fold and regional-specific K-fold regimes, respectively.
-We use CAMELS (Addor et al., 2017, Newman et al., 2014), restricted to the 531-basin subset with clear watershed boundaries (Newman et al., 2017).
+To compare spatial generalization on a well benchmarked dataset, we follow (Feng et al., 2021), testing prediction in ungauged basins (PUB, random spatial K-fold) and ungauged regions (PUR, regional spatial K-fold).
+We use CAMELS (Addor et al., 2017; Newman et al., 2014), restricted to the 531-basin subset with clear watershed boundaries (Newman et al., 2017).
 Basins were divided into 10 random groups for PUB and 7 contiguous regions for PUR, employing leave-one-out in both cases.
 To avoid leakage, all CAMELS-overlapping stations were removed during pretraining for PUB, and entire regions were excluded for PUR.
-Table 1: CAMELS Streamflow PUB and PUR Results
-Model Random holdout (ungauged basins) Regional holdout (ungauged regions)
-RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$ RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$
-LSTM - SL 1.402 1.360 0.762 0.636 1.609 1.457 0.743 0.554
-StefaLand - direct 1.882 1.849 0.538 0.395 1.982 1.949 0.230 0.201
-StefaLand - resConn 1.111 1.068 0.869 0.717 1.344 1.334 0.801 0.635
-StefaLand - no resConn 1.171 1.154 0.823 0.706 1.376 1.356 0.798 0.610
-StefaLand - scratch 1.355 1.332 0.801 0.661 1.516 1.378 0.771 0.560
-TerraMind - resConn 1.332 1.301 0.777 0.637 1.420 1.398 0.763 0.551
-The foundation model pretraining clearly provides a rare boost to generalization capability compared to supervised learning models (LSTM-SL and StefaLand-direct) across both PUB and PUR (Table 1).
-StefaLand-resConn’s RMSE is almost $20\%$ lower than that of supervised LSTM in PUB and $17\%$ lower in PUR, while the correlation are noticeably higher.
-The value of pretraining is confirmed by StefaLand-scratch, which appears to have only minor advantage compared to LSTM.
-The residual-connection architecture enables effective integration of transformer features with temporal dynamics, showing a modest benefit compared to StefaLand-no resConn, but both are clearly stronger than either LSTM or StefaLand-direct.
-Finally, TerraMind-resConn has quite comparable metrics to LSTM; the pretrained TerraMind offers no extra generalization value in this case.
+The success at these tasks would mean better flood forecasting information for populations in the world who do not have as many gauging stations around them.
+Table 1: CAMELS streamflow performance under PUB evaluation (random spatial holdout; ungauged basins). Values are median $\pm$ standard error across folds.
+Model RMSE $\downarrow$ ubRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$
+LSTM SL 1.402 $\pm$ 0.04 1.360 $\pm$ 0.04 0.762 $\pm$ 0.01 0.636 $\pm$ 0.04
+DLinear 2.012 $\pm$ 0.06 2.000 $\pm$ 0.06 0.598 $\pm$ 0.01 0.302 $\pm$ 0.48
+Informer 2.262 $\pm$ 0.07 2.237 $\pm$ 0.08 0.521 $\pm$ 0.01 0.104 $\pm$ 0.08
+Reformer 1.908 $\pm$ 0.09 1.871 $\pm$ 0.09 0.718 $\pm$ 0.00 0.270 $\pm$ 0.18
+TabPFN 2.727 $\pm$ 0.09 2.725 $\pm$ 0.09 0.718 $\pm$ 0.01 0.509 $\pm$ 0.014
+AlphaEarth LSTM 1.409 $\pm$ 0.10 1.393 $\pm$ 0.10 0.837 $\pm$ 0.11 0.618 $\pm$ 0.05
+AlphaEarth resConn 1.361 $\pm$ 0.09 1.345 $\pm$ 0.09 0.859 $\pm$ 0.01 0.647 $\pm$ 0.04
+StefaLand resConn 1.111 $\pm$ 0.04 1.068 $\pm$ 0.04 0.869 $\pm$ 0.01 0.717 $\pm$ 0.16
+Table 2: CAMELS streamflow performance (PUR: regional spatial holdout; ungauged regions). Values are median $\pm$ standard error across folds.
+Model RMSE $\downarrow$ ubRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$
+LSTM SL 1.609 $\pm$ 0.24 1.457 $\pm$ 0.22 0.743 $\pm$ 0.02 0.554 $\pm$ 0.13
+DLinear 2.019 $\pm$ 0.35 1.983 $\pm$ 0.34 0.597 $\pm$ 0.03 0.290 $\pm$ 0.99
+Informer 2.332 $\pm$ 0.41 2.295 $\pm$ 0.37 0.497 $\pm$ 0.01 0.046 $\pm$ 0.27
+Reformer 2.074 $\pm$ 0.33 1.978 $\pm$ 0.31 0.686 $\pm$ 0.02 0.257 $\pm$ 1.17
+TabPFN 2.709 $\pm$ 1.08 2.501 $\pm$ 0.89 0.576 $\pm$ 0.11 0.328 $\pm$ 0.12
+AlphaEarth LSTM 1.724 $\pm$ 0.85 1.685 $\pm$ 0.82 0.780 $\pm$ 0.07 0.456 $\pm$ 0.15
+AlphaEarth resConn 1.727 $\pm$ 0.74 1.684 $\pm$ 0.71 0.790 $\pm$ 0.06 0.520 $\pm$ 0.10
+StefaLand resConn 1.344 $\pm$ 0.21 1.334 $\pm$ 0.19 0.801 $\pm$ 0.02 0.635 $\pm$ 0.25
+The results in Tables 1 and 2 demonstrate that foundation-model pretraining substantially improves spatial generalization relative to purely supervised approaches under both PUB and PUR evaluation.
+StefaLand-resConn achieves the strongest performance across all reported metrics, reducing RMSE by approximately 20% relative to the supervised LSTM baseline under PUB and by about 16 to 17% under PUR, while also yielding consistently higher correlation and NSE.
+Among the baselines, the supervised LSTM remains competitive and outperforms linear and transformer-based sequence models, while TabPFN performs poorly with substantially higher RMSE and lower NSE despite moderate correlation, indicating that a tabular formulation fails to capture key temporal structure in rainfall runoff dynamics.
+AlphaEarth-based variants provide notable improvements over the supervised LSTM, particularly in correlation, but remain well below StefaLand-resConn across all metrics, even with residual feature reuse.
 We ran additional experiments that hybridize StefaLand with the HBV1.1 physics backbone on the same PUB/PUR splits, testing its ability to parameterize physics-based models.
-These hybrids achieved up to a 13% RMSE reduction and a 10% correlation gain compared to the LSTM–HBV1.1 baseline, demonstrating that pretraining strengthens hybridization with process models.
+These hybrids achieved up to a 13% RMSE reduction and a 10% correlation gain compared to the LSTM–HBV1.1 baseline, showing StefaLand’s .
 By constraining predictions with physics while leveraging StefaLand features, these hybrids further improve upon the general results above and highlight the versatility of the approach.
-Full results are provided in Appendix B, Table B1.
+Full results are provided in Appendix B, Table 9.
 On a related note, the supervised LSTM is not an easy benchmark to surpass.
-The original multi-basin LSTM and subsequent large-scale comparisons (Kratzert et al., 2019, Feng et al., 2021) showed that vanilla Transformers generally fail to outperform LSTMs on rainfall–runoff prediction (Liu et al., 2024, Table 1 therein).
+The original multi-basin LSTM and subsequent large-scale comparisons (Kratzert et al., 2019; Feng et al., 2021) showed that vanilla Transformers generally failed to outperform LSTMs on rainfall–runoff prediction and modified transforms essentially tied LSTM (Liu et al., 2024, Table 1 therein).
 The LSTM NSE values reported here are very similar to those in the domain literature (Feng et al., 2021).
-Broader evaluations likewise continue to rank LSTM-family models among the best-performing approaches (Lees et al., 2021), and even Google’s global flood-forecasting system adopts an encoder–decoder LSTM backbone (Nearing et al., 2024).
-3.2 Global Streamflow
-We designed a global-scale runoff prediction experiment to assess robustness and generalization worldwide.
-We filtered global basin datasets based on data completeness and retained 3,434 basins.
-To manage computational cost, we employed random hold-out sampling with three-fold cross-validation.
-Additionally, we implemented a regionally hold-out continental scenario (RH-C), excluding all basins from North America, South America, and Europe, respectively, from training, then evaluating on the excluded continent.
-Table 2: Global streamflow prediction across 3,434 basins worldwide
-Random holdout Regional holdout
-(ungauged basins) (ungauged continents)
-Model RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$ RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$
-LSTM - SL 0.870 0.864 0.798 1.253 1.202 0.672
-StefaLand - direct 0.749 0.751 0.843 1.075 1.048 0.697
-TerraMind - resConn 1.156 1.111 0.580 1.234 1.158 0.670
-Cross-continental transfer testing using three-fold validation. Detailed metric calculations in Appendix E.
-Results indicate that StefaLand-direct outperformed both LSTM-SL and the fine-tuned vision-based Transformer model (TerraMind) across all metrics for either random or regional holdout (Table 2).
-Specifically, in the random holdout scenario, StefaLand achieved an RMSE of 0.749, approximately 14% lower than LSTM (0.87), with Corr improving to 0.843.
-Notably, StefaLand’s unbiased RMSE closely matched its RMSE, indicating errors were primarily random fluctuations rather than systematic bias.
-In contrast, TerraMind exhibited a higher RMSE of 1.156 and a low Corr of 0.580.
-We hypothesize this is because the pretraining satellite image data for the vision foundation network TerraMind did not have high relevance to hydrologic predictions like streamflow.
-We note that using TerraMind here was exploratory, as it was not specifically designed or pretrained for this use case.
-Under the more challenging RH-C scenario, all models exhibited larger errors, but StefaLand continued to perform best.
-Its RMSE was 1.075, approximately 14.1% lower than LSTM’s 1.253, while also maintaining the highest Corr (0.697).
-Moreover, StefaLand’s ubRMSE remained close to its RMSE, confirming robust correction of regional-scale biases even under extreme out-of-domain conditions.
-Conversely, LSTM and TerraMind showed larger gaps between ubRMSE and RMSE, highlighting challenges in producing hydrologically-relevant features.
+Google’s global flood-forecasting system adopts an encoder–decoder LSTM backbone (Nearing et al., 2024).
+The gains demonstrated are also substantial: to provide some context, when LSTM raised NSE from 0.64 to 0.73 (without ensemble), it was considered a generational change in predictive performance (Nearing et al., 2021; Feng et al., 2021), and no method other than ensembling more models cleanly surpassed default LSTM on the PUB test.
+3.2 Caravan Global Streamflow
+We designed a global-scale runoff prediction experiment to assess spatial robustness and generalization across diverse hydroclimatic regimes.
+We use the open-source Caravan dataset (Kratzert et al., 2023), a global community dataset that integrates meteorological forcings, catchment attributes, and observed streamflow for river basins worldwide.
+From the full Caravan archive (16,300 basins), we restricted our evaluation to the official CAMELS-family datasets, yielding 3,278 basins with relatively fuller streamflow records.
+From this set, we constructed a curated subset of 3,026 basins by excluding those with more than 75% missing streamflow observations during the 10-year evaluation period from 2010 to 2020, ensuring sufficient data quality for model training and testing.
+This filtering removes sparsely observed basins while preserving substantial geographic, climatic, and physiographic diversity.
+We employed a five-fold random spatial holdout protocol to evaluate generalization to unseen basins at the global scale.
+Table 3: Caravan streamflow performance under random spatial holdout. Values are median $\pm$ standard error across folds
+Model RMSE $\downarrow$ ubRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$
+LSTM SL 1.623 $\pm$ 0.87 1.589 $\pm$ 0.84 0.496 $\pm$ 0.08 0.398 $\pm$ 0.74
+DLinear 1.605 $\pm$ 0.99 1.453 $\pm$ 0.99 0.512 $\pm$ 0.07 0.412 $\pm$ 0.81
+Informer 1.562 $\pm$ 1.04 1.471 $\pm$ 0.98 0.435 $\pm$ 0.07 0.256 $\pm$ 0.92
+Reformer 1.574 $\pm$ 1.06 1.507 $\pm$ 0.97 0.451 $\pm$ 0.08 0.378 $\pm$ 0.87
+AlphaEarth LSTM 1.533 $\pm$ 0.97 1.495 $\pm$ 0.89 0.523 $\pm$ 0.09 0.489 $\pm$ 0.89
+AlphaEarth resConn 1.424 $\pm$ 1.03 1.406 $\pm$ 1.00 0.541 $\pm$ 0.07 0.514 $\pm$ 0.79
+StefaLand resConn 1.457 $\pm$ 0.87 1.401 $\pm$ 0.83 0.589 $\pm$ 0.06 0.533 $\pm$ 0.70
+Across the global Caravan benchmark (Table 3), StefaLand-resConn achieves the highest correlation and NSE among all evaluated models, indicating improved agreement with observed runoff dynamics under substantial spatial heterogeneity.
+While error-based metrics such as RMSE and ubRMSE vary considerably across basins, StefaLand-resConn consistently improves correlation relative to both supervised sequence models and alternative pretrained representations.
+AlphaEarth-based variants yield noticeable gains over purely supervised baselines, but remain below StefaLand-resConn across most metrics.
+The Caravan is a more noisy dataset due to global inconsistencies in data collection.
+These results suggest that attribute-centric pretraining combined with residual temporal adaptation provides a foundation to improve global hydrologic services.
 3.3 Global Soil Moisture
-We evaluated finetuning StefaLand for soil moisture predictions following Liu et al.
-(2023a), using ISMN (Dorigo et al., 2011; 2013a).
+We evaluated finetuning StefaLand for soil moisture predictions following (Liu et al., 2023a), using ISMN (Dorigo et al., 2011, 2013a).
 Even though there is a globally covering satellite-based product for soil moisture, the data quality can hardly match that of in-situ moisture sensors; thus the ability to generalize in-situ data is valuable.
@@
 We performed five-fold spatial cross-validation for random holdout and a regional holdout on Europe, training on all other continents while excluding European sites (129) for testing.
-We tested StefaLand-direct, StefaLand with and without residual connections (resConn / noResConn), a from-scratch ablation, an LSTM baseline, and a version using IBM’s TerraMind encoder with resConn.
-LSTM again serves as the established state-of-the-art baseline (Wang et al., 2024, Liu et al., 2023b).
-Table 3: Soil moisture prediction across 1,316 ISMN stations
-Random location holdout Regional holdout
-(random sites) (Europe)
-Model RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$ RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$
-LSTM - SL 0.073 0.055 0.764 0.112 0.053 0.510
-StefaLand - direct 0.140 0.103 0.637 0.135 0.112 0.503
-StefaLand - resConn 0.068 0.054 0.783 0.090 0.059 0.638
-StefaLand - no resConn 0.075 0.057 0.741 0.095 0.058 0.545
-StefaLand - scratch 0.074 0.058 0.749 0.108 0.064 0.528
-TerraMind - resConn 0.083 0.062 0.694 0.101 0.080 0.519
-PrithviWxC - resConn 0.081 0.060 0.703 0.103 0.079 0.523
-5-fold spatial validation and cross-continental validation on Europe (129 sites). Detailed metric calculations in Appendix E.
-Against these strong baselines, the soil moisture experiments confirm the superiority of the StefaLand-resConn architecture (Table 3).
-It achieves the best performance, with RMSE of 0.068 and correlation of 0.783 for random holdout, and maintains strong performance even in cross-continental testing on Europe (RMSE = 0.090, Corr = 0.638).
-The direct StefaLand model performs poorly because it lacks an effective mechanism to capture task-specific temporal dependencies.
-Similar to the streamflow cases above, TerraMind and PrithviWxC showed no performance benefit compared to LSTM, underscoring the challenge of repurposing vision foundation models across domains and the importance of pretraining with geoscience-relevant variables.
-The regional holdout on Europe further demonstrates StefaLand-resConn’s superior spatial generalization, achieving a 25% correlation improvement over the LSTM baseline in this difficult extrapolation scenario.
-We want to emphasize that TerraMind and PrithviWxC were not designed for the hydrologic and land–surface prediction tasks studied here.
-They excel in their intended domains of Earth observation imagery and atmospheric variables, but are out-of-domain for dynamic land–surface and hydrologic modeling.
-We include them only to explore whether such models transfer any useful signal in our setting.
-These results should not be interpreted as evidence against their capability in other applications, but rather as a reflection of the mismatch between their pretraining objectives and the land–surface tasks we target.
-Nevertheless, we believe these benchmarks are helpful for clarifying their respective strengths, since the general AI community may not be familiar with these datasets and models that have been marketed as ‘any-to-any’ generative EO foundation models.
-Ultimately we view StefaLand as complementary to satellite-image foundation models, targeting a different but relevant data regime.
-We also could not identify other foundation models for land surface.
+LSTM again serves as the established state-of-the-art baseline (Wang and others, 2024; Liu et al., 2023b).
+Table 4: Soil moisture prediction across ISMN (random location holdout). Values are median $\pm$ standard error across folds.
+Model RMSE $\downarrow$ ubRMSE $\downarrow$ Corr $\uparrow$
+LSTM SL 0.073 $\pm$ 0.002 0.055 $\pm$ 0.001 0.764 $\pm$ 0.006
+DLinear 0.088 $\pm$ 0.001 0.065 $\pm$ 0.001 0.612 $\pm$ 0.007
+Informer 0.102 $\pm$ 0.002 0.082 $\pm$ 0.001 0.232 $\pm$ 0.012
+Reformer 0.090 $\pm$ 0.002 0.071 $\pm$ 0.001 0.568 $\pm$ 0.015
+TabPFN 0.068 $\pm$ 0.004 0.057 $\pm$ 0.003 0.404 $\pm$ 0.016
+AlphaEarth LSTM 0.075 $\pm$ 0.001 0.062 $\pm$ 0.001 0.427 $\pm$ 0.019
+AlphaEarth resConn 0.082 $\pm$ 0.001 0.067 $\pm$ 0.001 0.406 $\pm$ 0.012
+StefaLand resConn 0.068 $\pm$ 0.001 0.054 $\pm$ 0.001 0.783 $\pm$ 0.005
+Table 5: Soil moisture prediction across ISMN cross-continental validation on Europe (129 sites).
+Model RMSE $\downarrow$ ubRMSE $\downarrow$ Corr $\uparrow$
+LSTM SL 0.112 0.053 0.510
+DLinear 0.093 0.051 0.623
+Informer 0.088 0.063 0.358
+Reformer 0.087 0.058 0.553
+TabPFN 0.081 0.054 0.401
+AlphaEarth LSTM 0.087 0.068 0.313
+AlphaEarth resConn 0.090 0.071 0.308
+StefaLand resConn 0.090 0.059 0.638
+Against these baselines, the soil moisture experiments demonstrate advantages of the StefaLand-resConn architecture (Tables 4 and 5).
+For random spatial holdout, StefaLand-resConn attains the highest correlation (0.783) while also achieving the lowest ubRMSE (0.054) and matching the best RMSE (0.068), outperforming the LSTM baseline across all metrics.
+Although TabPFN achieves comparable RMSE, its substantially lower correlation indicates weaker agreement with temporal variability relative to sequence-based models.
+The regional holdout on Europe represents a more challenging cross-continental generalization setting.
+In this case, StefaLand-resConn again achieves the highest correlation (0.638), exceeding all baselines including LSTM and linear or transformer-based models, while maintaining competitive RMSE and ubRMSE.
+In contrast, AlphaEarth-based variants show limited transfer performance, with both LSTM and residual-connection formulations yielding lower correlations despite similar error magnitudes.
+Overall, these results highlight that residual adaptation on top of geoscience-pretrained representations is critical for robust soil moisture prediction, particularly under strong spatial distribution shift.
 3.4 Soil Property Prediction
 There are different soil datasets, each collected with different protocols and data processing techniques, resulting in significant discrepancies.
-In this test, we finetuned StefaLand to predict in-situ soil profile data from another dataset (ISRIC).
-This application can produce a seamless dataset that is consistent with a set of in-situ data, improving data availability and addressing systematic biases.
+In this test, we finetuned StefaLand to predict in situ soil profile data from another dataset (ISRIC).
+This application can produce a seamless dataset that is consistent with a set of in situ data, improving data availability and addressing systematic biases.
 In addition, it helps us understand the noise associated with each dataset.
 StefaLand’s pretraining soils dataset is HWSD, which has some overlap but also extensive differences from ISRIC, which is larger and potentially noisier.
-We finetuned StefaLand to predict one soil texture property (e.g., clay percentage) in ISRIC while masking the corresponding complementary attribute (e.g., sand) from the same profile to avoid information leakage, probing how easy it is to infer soil properties using other attributes such as climate, terrain and land cover.
-We compared StefaLand with and without pretraining against a supervised random forest baseline.
-Train-test splits can be found here C.2
-Table 4: In-situ soil property prediction using ISRIC WoSIS data.
-Model WoSIS Property Corr $\uparrow$ $R^{2}$ $\uparrow$
-StefaLand - direct Clay 0.197 0.038
-StefaLand - finetune Clay 0.509 0.259
-Random Forest Clay 0.456 0.207
-Linear Regression Clay 0.138 0.019
-StefaLand - direct Sand 0.253 0.064
-StefaLand - finetune Sand 0.704 0.495
-Random Forest Sand 0.585 0.342
-Linear Regression Sand 0.347 0.120
-StefaLand with finetuning achieved markedly higher predictive power for both clay and sand fractions, substantially outperforming direct training, random forest, and linear regression baselines.
-The other models struggle in this task, with random forest (RF) scoring noticeably lower.
-This task is not dynamical prediction, but a test of the models ability to build deep representations of the landscape that can be adapted to infer certain attributes using cross-variable group connections.
-StefaLand-finetune’s advantage against random forest suggests there are indeed deep representations uncovered by pretraining but not exploited by random forest.
-Finetuning StefaLand is also apparently superior to training from scratch as supervised learning (StefaLand - direct), again highlighting the value of pretraining.
-These results highlight StefaLand’s ability to reconcile disparate attribute datasets and improve the utility of noisy in-situ observations.
+We finetuned StefaLand to predict one soil texture property (e.g., clay percentage) in ISRIC while masking the corresponding complementary attribute (e.g., sand) from the same profile to avoid information leakage, probing how easy it is to infer soil properties using other attributes such as climate, terrain, and land cover.
+We compared StefaLand finetuning against AlphaEarth features incorporated into a random forest model, as well as supervised random forest and linear regression baselines.
+Train test splits can be found here C.5.
+Table 6: Soil composition prediction (clay). Values are median $\pm$ standard error across folds.
+Model Corr $\uparrow$ $R^{2}$ $\uparrow$
+Random Forest 0.456 $\pm$ 0.02 0.207 $\pm$ 0.01
+Linear Regression 0.138 $\pm$ 0.01 0.019 $\pm$ 0.01
+AlphaEarth RF 0.462 $\pm$ 0.03 0.205 $\pm$ 0.02
+StefaLand finetuned 0.509 $\pm$ 0.01 0.259 $\pm$ 0.01
+Table 7: Soil composition prediction (sand). Values are median $\pm$ standard error across folds.
+Model Corr $\uparrow$ $R^{2}$ $\uparrow$
+Random Forest 0.585 $\pm$ 0.01 0.342 $\pm$ 0.01
+Linear Regression 0.347 $\pm$ 0.01 0.120 $\pm$ 0.01
+AlphaEarth RF 0.615 $\pm$ 0.01 0.373 $\pm$ 0.01
+StefaLand finetuned 0.704 $\pm$ 0.01 0.495 $\pm$ 0.01
+StefaLand finetuning achieves the strongest performance for both clay and sand prediction, outperforming linear regression, random forest, and AlphaEarth RF baselines in terms of both correlation and $R^{2}$.
+While AlphaEarth RF improves over standard random forest baselines, it remains consistently below StefaLand finetuned across both soil properties.
+These results indicate that StefaLand’s pretrained representations are particularly effective when adapted to infer soil texture attributes from complementary variables and static environmental context.
 3.5 Landslide Susceptibility Prediction
@@
 We next evaluated StefaLand for landslide susceptibility prediction using the SLIDO dataset from the State of Oregon, which provides detailed landslide occurrence records.
-Following Liu et al.
-(2025), this is a binary classification task indicating the presence or absence of landslides in a 30m by 30m patch.
-We finetuned StefaLand by extracting frozen hidden features and concatenating them with a 2D CNN, then retrained the CNN classifier to assess StefaLand’s ability to provide generalizable geoscience features.
+Following (Liu et al., 2025a), this is a binary classification task indicating the presence or absence of landslides in a 30m by 30m patch.
 Table 8: Landslide susceptibility prediction results on the Oregon SLIDO dataset.
 Model Accuracy $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ ROC AUC $\uparrow$
-Logistic Regression 0.742 0.707 0.792 0.47 0.819
-Random Forest 0.751 0.703 0.834 0.763 0.839
-CNN2D 0.778 0.760 0.787 0.773 0.863
-StefaLand + CNN2D 0.799 0.791 0.793 0.792 0.875
-Note: All baseline results (Logistic Regression, Random Forest, CNN2D) are taken from previously published 30m-resolution experiments in Liu et al.
-(2025), except for StefaLand + CNN2D, which represents our proposed method.
-Results show that StefaLand’s pretrained features improved the CNN’s generalization, yielding modest gains across all metrics except Recall.
-This is a particularly difficult baseline to improve so even modest gains are rare.
-ROC AUC increased from 0.863 to 0.875, and precision rose from 0.760 to 0.791, reflecting fewer false positives.
-While random forest finds a high Recall, this is a particular realization with a large tradeoff on Precision.
-StefaLand produces overall well-rounded predictions, with both better Precision and Recall than CNN2D.
+Logistic Regression 2D 0.744 0.720 0.795 0.756 0.823
+Random Forest 2D 0.765 0.737 0.822 0.777 0.849
+CNN2D 0.880 0.896 0.858 0.877 0.954
+StefaLand + CNN2D 0.903 0.859 0.963 0.908 0.911
+Note: All baseline results (Logistic Regression 2D, Random Forest 2D, CNN2D) are taken from previously published 10m-resolution experiments in (Liu et al., 2025a), except for StefaLand + CNN2D, which represents our proposed method.
+We finetuned StefaLand by extracting frozen hidden features and concatenating them with a 2D CNN, then retrained the CNN classifier to assess StefaLand’s ability to provide generalizable geoscience features.
+Results show that StefaLand’s pretrained features improved the CNN’s generalization, yielding modest gains across most metrics.
+This is a particularly difficult baseline to improve, so even modest gains are rare.
+Recall increased from 0.858 to 0.963, and accuracy rose from 0.880 to 0.903, reflecting fewer misclassifications overall.
+While the random forest achieves high precision in this particular run, it does so at the cost of lower recall.
+Overall, StefaLand produces well-rounded predictions, with both higher accuracy and recall than CNN2D.
+3.6 Ablations
+Figure 2: Ablation impact matrix across evaluation settings using RMSE.
+Each cell shows the percent RMSE increase relative to the full StefaLand ResConn model (lower is better), highlighting the contributions of pretraining and task adaptation.
+Blank or hatched cells indicate ablations not evaluated for that setting.
+To isolate the contributions of pretraining and task adaptation in StefaLand, we conduct a structured ablation study across multiple tasks and spatial evaluation protocols, comparing four variants against the proposed pretrained model with task adapter.
+Direct prediction removes pretraining entirely and treats the target as an additional input channel, training the model end to end only on downstream data.
+Scratch training uses the full StefaLand architecture but trains all parameters from random initialization, isolating the effect of large-scale pretraining.
+No adapter removes the task-specific adapter and feeds pretrained embeddings directly to the task decoder, testing whether simple feature reuse is sufficient.
+Linear probing freezes the pretrained encoder and applies only a linear prediction head, representing the weakest form of adaptation.
+Across streamflow, soil moisture, and global Caravan experiments, performance degradation relative to the full model is dominated by the removal of pretraining.
+Direct prediction and scratch training consistently yield the largest errors, especially under spatial generalization.
+In contrast, removing or simplifying the adapter produces smaller but consistent performance drops, indicating that task adaptation improves results but cannot compensate for the absence of pretrained representations.
+These trends are consistent across random and regional holdouts and across tasks with differing data characteristics, highlighting pretrained land-surface representations as the primary driver of StefaLand’s performance.
+Numerical results for ablations, are provided in Appendix B.4.
 4 Discussion
 4.1 Key Findings and Contributions
-As we review the literature, the methods to improve spatial generalization to data‐scarce regions are rare and rarely effective (Beery et al., 2018, Gacu et al., 2025).
-In contrast, StefaLand, combined with lightweight fine‐tuning heads, achieves state-of-the-art or competitive performance across four broad problem classes: streamflow (both CAMELS and global), soil moisture, soil composition, and landslide susceptibility, while also strengthening the parameterization of differentiable process-based models.
-Across tasks, the strongest gains come from architectures that fuse StefaLand embeddings with explicit temporal modeling via residual connections, indicating that pretraining on attribute-based spatiotemporal structure yields problem-relevant representations while temporal heads resolve sequence dynamics.
-These outcomes support the premise that foundation models can democratize prediction quality in data-scarce regions by improving out-of-domain transfer.
-The five dynamic prediction cases, along with benchmark models that reproduce state-of-the-art results in the literature, together paint a clear picture.
-The pretraining of StefaLand on attributes builds deep landscape representations.
-These features are highly relevant to hydrologic prediction tasks and can avoid overfitting compared to those built in task-specific supervised learning (LSTM or StefaLand-direct), improving model spatial generalization.
-They are also decidedly more relevant to such tasks than those obtained from existing Earth Foundation models expensively trained on massive amounts of satellite images.
-It shows that the importance of larger pretraining data may not necessarily exceed that of problem relevance, and image-like data may not be the optimal data representation for such tasks.
-We stress that our approach is complementary to satellite-based foundation models: where they exploit large-scale visual patterns, StefaLand focuses on problem-relevant attributes, offering an efficient and domain-specific alternative.
-We have tested a large number of adapter formulations.
-While we can certainly test more finetuning options, especially internal foundation model layers, the computational and data-storage demands already start to be limiting.
-Our attribute-based approach is at least an order of magnitude more efficient than pixel-wise satellite transformers.
-Our underlying transformer has far fewer parameters (roughly 12 million) and avoids the heavy data management requirements of image-centric pretraining.
-TerraMind’s larger configuration corresponds to about 7,680 GPU hours (Jakubik et al., 2025), PrithviWxC require roughly 23,040 GPU hours (Schmude et al., 2024) and Aurora required roughly 14,592 GPU hours (Bodnar et al., 2025).
-In contrast, StefaLand’s attribute-based pretraining requires only about 720 GPU hours, making high-quality spatial generalization feasible on limited budgets.
-These figures do not include the large gaps in data-storage and transfer demands during pretraining (StefaLand, Terramind and PrithviWxC used $\approx 2$, $\approx 11$ and $\approx 27$ data as mentioned earlier).
-In fact, it is costly and resource-straining to run the satellite-imagery-focused foundation models, which impedes us from comparing with them in every case.
-Finally, to the best of our effort, we could not identify other earth foundation models that are designed for land-surface predictions.
+Methods that reliably improve spatial generalization to data-scarce regions remain rare in the literature (Beery et al., 2018; Gacu et al., 2025).
+In contrast, StefaLand combined with lightweight task-specific heads achieves state-of-the-art or competitive performance across four broad problem classes: streamflow prediction (both CAMELS and global), soil moisture, soil composition, and landslide susceptibility, while also strengthening the parameterization of differentiable process-based models.
+Across tasks, the strongest gains arise from architectures that combine pretrained StefaLand representations with explicit temporal modeling via residual connections, indicating that attribute-centric pretraining captures problem-relevant structure while downstream sequence models resolve task-specific dynamics.
+Together, these results support the premise that foundation models can improve out-of-domain transfer and help democratize prediction quality in data-scarce regions.
+Across five dynamic prediction settings, evaluated against benchmark models that reproduce state-of-the-art results reported in prior work, a consistent pattern emerges.
+Pretraining StefaLand on landscape attributes yields deep representations that are highly relevant to hydrologic and land-surface prediction tasks, enabling stronger spatial generalization than purely supervised approaches such as LSTM-based models or training from scratch.
+Comparisons with alternative pretrained representations further suggest that problem relevance of the pretraining signal is at least as important as scale alone, particularly for tasks governed by physical and environmental processes rather than visual appearance.
+We emphasize that this attribute-based approach is complementary to satellite-centric foundation models: while image-based models excel at extracting large-scale visual patterns, StefaLand focuses on structured environmental attributes that are directly aligned with land-surface processes, offering a more targeted and efficient alternative.
+A key practical advantage of StefaLand is computational efficiency.
+The attribute-based pretraining strategy avoids pixel-wise processing and large image volumes, resulting in a compact transformer with roughly 12 million parameters and substantially lower data-management requirements.
+As a result, StefaLand pretraining can be performed on modest computational budgets, making large-scale spatial generalization experiments feasible without access to specialized infrastructure.
+While additional adapter and finetuning variants could be explored, the current results already demonstrate strong performance across diverse tasks, and further expansion must balance marginal gains against growing computational and storage costs.
 4.2 Limitations and Future Work
@@
 Appendix A Detailed Model Architecture
-This appendix provides the complete mathematical formulation of the StefaLand model architecture more detailed then 2.
+This appendix provides the complete mathematical formulation of the StefaLand model architecture.
 A.1 Embedding Dynamic and Static Inputs
@@
 These differentiable versions can be trained end-to-end within neural network frameworks, combining process understanding with machine learning flexibility (Shen et al., 2023).
-For the process-based backbone, we employed the Hydrologiska Byråns Vattenbalansavdelning (HBV) model (Aghakouchak and Habib, 2010, Beck et al., 2020, Bergström, 1976; 1992, Seibert and Vis, 2012), a relatively simple bucket-type conceptual hydrologic model.
+For the process-based backbone, we employed the Hydrologiska Byråns Vattenbalansavdelning (HBV) model (Aghakouchak and Habib, 2010; Beck et al., 2020; Bergström, 1976, 1992; Seibert and Vis, 2012), a relatively simple bucket-type conceptual hydrologic model.
 HBV has state variables like snow storage, soil water, and subsurface storage, and can simulate flux variables such as evapotranspiration (ET), recharge, surface runoff, shallow subsurface flow, and groundwater flow.
@@
 These physics-based approaches incorporate hydrological process understanding while maintaining the ability to learn from data..
-Table B1: CAMELS Streamflow PUB and PUR Results (Physics-Based Models)
+Table 9: CAMELS Streamflow PUB and PUR Results (Physics-Based Models)
 Model Random holdout (ungauged basins) Regional holdout (ungauged regions)
@@
 As shown in the table below, linear regression performs poorly across all tasks by a fair margin when compared to our neural network approaches.
-Table B2: Additional experiments with linear regression baselines.
+Table 10: Additional experiments with linear regression baselines.
 Experiment Random holdout Regional holdout
@@
 Camels Streamflow Linear Regression 2.190 2.180 0.500 2.260 2.250 0.500
-Global Streamflow Linear Regression 1.823 1.746 0.252 1.816 1.721 0.248
+Caravan Streamflow Linear Regression 2.612 2.431 0.142
 Soil Moisture Linear Regression 0.120 0.101 0.188 0.121 0.103 0.187
+B.3 External Foundation Model Comparisons
+For completeness, we explored several existing foundation models developed for Earth observation and atmospheric applications, including TerraMind, PrithviWxC, and Galileo (Jakubik et al., 2025; Hsu et al., 2024; Tseng et al., 2025).
+All models were evaluated using consistent downstream protocols, with pretrained encoders either frozen or minimally adapted and paired with task-specific heads comparable to those used for StefaLand.
+These experiments were intended to probe the extent to which representations learned from large-scale EO or atmospheric data transfer to land–surface and hydrologic prediction tasks.
+These experiments are not intended as exhaustive benchmarks or as performance upper bounds for the evaluated models, but rather as feasibility probes to understand whether their pretrained representations can be directly repurposed for land–surface and geohazard tasks under lightweight adaptation
+Because these models differ substantially in their native input formats and pretraining objectives, task-specific adaptations were required.
+TerraMind and PrithviWxC were coupled with the same residual adaptation architecture used for StefaLand, with only the added adaptation units trained.
+Due to the intensive data and storage requirements of PrithviWxC, inputs were restricted to surface-level variables most directly related to land–surface interactions, together with static attributes, while multi-level atmospheric variables were excluded.
+For context, StefaLand, TerraMind, and PrithviWxC were pretrained on approximately 2, 11, and 27 terabytes of data, respectively.
+Galileo was evaluated in the landslide susceptibility setting, where inputs consist of multichannel static environmental attributes rather than multispectral time series.
+To accommodate this difference, we applied an input adaptation strategy that pooled spatial features along one dimension, projected the 17 environmental channels to 12 channels via a linear layer, and treated the remaining spatial dimension as pseudo-temporal input.
+We initialized Galileo using its pretrained Transformer encoder (768-dimensional embeddings, four attention layers with 12 heads each, and a feed-forward dimension of 3072), followed by a two-layer classification head (768→256→1) with ReLU activation and dropout (rate 0.3).
+The resulting model was fine-tuned end-to-end using AdamW (learning rate $10^{-4}$, weight decay 0.01), with a batch size of 128 for 1000 epochs on the same train–test split as other landslide experiments, using gradient clipping with a maximum norm of 1.0.
+Collectively, these experiments provide a broad exploratory comparison of how foundation models pretrained on EO imagery or atmospheric data behave when adapted to land–surface and geohazard prediction tasks.
+Table 11: Performance of external foundation model baselines across tasks. All models use frozen pretrained encoders with the same residual adaptation head.
+Task Model RMSE $\downarrow$ µbRMSE $\downarrow$ Corr $\uparrow$
+CAMELS Streamflow (PUB) TerraMind-resConn 1.332 $\pm$ 0.0410 1.301 $\pm$ 0.0375 0.777 $\pm$ 0.0071
+CAMELS Streamflow (PUR) TerraMind-resConn 1.420 $\pm$ 0.2021 1.398 $\pm$ 0.1932 0.763 $\pm$ 0.0172
+Soil Moisture (Random) TerraMind-resConn 0.083 $\pm$ 0.0021 0.062 $\pm$ 0.0007 0.694 $\pm$ 0.0289
+PrithviWxC-resConn 0.081 $\pm$ 0.0019 0.060 $\pm$ 0.0004 0.703 $\pm$ 0.0390
+Soil Moisture (Europe) TerraMind-resConn 0.101 0.080 0.519
+PrithviWxC-resConn 0.103 0.079 0.523
+Table 12: Landslide susceptibility prediction using the Galileo foundation model compared with published and proposed baselines on the Oregon SLIDO dataset.
+Model (10m) Accuracy $\uparrow$ Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ ROC AUC $\uparrow$
+Logistic Regression 2D 0.744 0.720 0.795 0.756 0.823
+Random Forest 2D 0.765 0.737 0.822 0.777 0.849
+CNN2D 0.880 0.896 0.858 0.877 0.854
+Galileo + CNN2D 0.750 0.764 0.720 0.742 0.834
+StefaLand + CNN2D 0.903 0.859 0.963 0.908 0.911
+B.4 Ablations
+Table 13: StefaLand ablations on CAMELS streamflow under random (PUB) and regional (PUR) spatial holdout.
+Variant PUB: random holdout (ungauged basins) PUR: regional holdout (ungauged regions)
+RMSE $\downarrow$ ubRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$ RMSE $\downarrow$ ubRMSE $\downarrow$ Corr $\uparrow$ NSE $\uparrow$
+StefaLand direct 1.882 $\pm$ 0.0901 1.849 $\pm$ 0.0691 0.538 $\pm$ 0.0105 0.395 $\pm$ 0.1431 1.982 $\pm$ 0.4012 1.949 $\pm$ 0.3871 0.230 $\pm$ 0.0784 0.201 $\pm$ 1.2201
+StefaLand scratch 1.355 $\pm$ 0.0394 1.332 $\pm$ 0.0368 0.801 $\pm$ 0.0031 0.661 $\pm$ 0.0372 1.516 $\pm$ 0.3723 1.378 $\pm$ 0.3496 0.771 $\pm$ 0.0211 0.560 $\pm$ 0.3120
+StefaLand noResConn 1.171 $\pm$ 0.0325 1.154 $\pm$ 0.0319 0.823 $\pm$ 0.0022 0.706 $\pm$ 0.3260 1.376 $\pm$ 0.1987 1.356 $\pm$ 0.1712 0.798 $\pm$ 0.0195 0.610 $\pm$ 0.1345
+StefaLand linear 1.452 $\pm$ 0.0542 1.366 $\pm$ 0.0533 0.751 $\pm$ 0.0127 0.661 $\pm$ 0.0430 1.484 $\pm$ 0.2588 1.453 $\pm$ 0.2402 0.672 $\pm$ 0.0391 0.542 $\pm$ 0.3100
+StefaLand resConn 1.111 $\pm$ 0.0378 1.068 $\pm$ 0.0374 0.869 $\pm$ 0.0067 0.717 $\pm$ 0.1600 1.344 $\pm$ 0.2097 1.334 $\pm$ 0.1873 0.801 $\pm$ 0.0220 0.635 $\pm$ 0.2460
+Table 14: StefaLand ablation study on soil moisture prediction under random and regional holdout.
+Variant Random location holdout Regional holdout (Europe)
+RMSE $\downarrow$ ubRMSE $\downarrow$ Corr $\uparrow$ RMSE $\downarrow$ ubRMSE $\downarrow$ Corr $\uparrow$
+StefaLand direct 0.140 $\pm$ 0.0431 0.103 $\pm$ 0.0041 0.637 $\pm$ 0.0352 0.135 0.112 0.503
+StefaLand scratch 0.074 $\pm$ 0.0011 0.058 $\pm$ 0.0003 0.749 $\pm$ 0.0172 0.108 0.064 0.528
+StefaLand noResConn 0.075 $\pm$ 0.0009 0.057 $\pm$ 0.0001 0.741 $\pm$ 0.0201 0.095 0.058 0.545
+StefaLand linear 0.084 $\pm$ 0.0010 0.061 $\pm$ 0.0002 0.720 $\pm$ 0.0192 0.100 0.063 0.393
+StefaLand resConn (proposed) 0.068 $\pm$ 0.0013 0.054 $\pm$ 0.0004 0.783 $\pm$ 0.0054 0.090 0.059 0.638
+Figure 3: Adapter ablation on streamflow random spatial splits (PUB).
+Boxplots summarize per-basin performance distributions for five adapter designs (Feedforward, Bottleneck, MoE, Gated, Residual).
+We report RMSE (lower is better), correlation (higher is better), and NSE (higher is better) across held-out basins, showing that the Residual adapter yields the most consistent gains, particularly in Corr and NSE.
 Appendix C Experimental Details
-Table C1: StefaLand Pretraining Configuration
+C.1 Model Configurations and Hyperparameters
+Table 15: StefaLand pretraining configuration.
 Parameter Value
@@
 Task pretrain
-Model MFFormer_dec_LSTM
+Model Stefaland_dec_LSTM
 Random seed 111
@@
 Static loss ratio 0.5
-Table C2: StefaLand Pretraining Variables and Sources
-Variable Type Variable Name Source
-Time Series Forcings Precipitation, Short-wave solar radiation downwards, Relative humidity, Maximum temperature, Minimum temperature, Potential evapotranspiration Multi-Source Weather (MSWX) and Multi-Source Weighted-Ensemble Precipitation (MSWEP) (Beck et al., 2022; 2019)
-Static Attributes Forest cover fraction, grassland cover fraction Climate Change Initiative (CCI) land cover dataset (ESA, 2017)
-Normalized Difference Vegetation Index (NDVI) Terra Moderate Resolution Imaging Spectroradiometer (MODIS) Vegetation Indices (MOD13A3) (Didan, 2015a)
-Sand, silt, clay fractions Harmonized World Soil Database (HWSD) (FAO et al., 2012)
-Elevation, slope, aspect Global Multi-resolution Terrain Elevation Data (GMTED) (Danielson and Gesch, 2011, Ramcharan et al., 2018)
-Soil depth Global 1-km Gridded Thickness of Soil, Regolith, and Sedimentary Deposit Layers (Pelletier et al., 2016)
-Carbonate sedimentary rock fraction Global Lithological Map (GLiM) (Hartmann and Moosdorf, 2012)
-Rock porosity, permeability GLobal HYdrogeology MaPS (GLHYMPS) (Gleeson et al., 2014)
-Population density Gridded Population of the World (GPW) v4 dataset (CIESIN, 2016)
-GDP per capita; population density Gross Domestic Product and Human Development Index over 1990-2015 (Kummu et al., 2018)
-Forest intact fraction Intact Forest Landscapes Data (Potapov et al., 2017)
-Outputs None (self-supervised pretraining)
-Table C3: Attribute Groups Used in Group Masking Pretraining
+Table 16: Attribute groups used in group masking pretraining.
 Group Variables
@@
 Climate aridity, meanP, ETPOT_Hargr, meanTa, seasonality_P, seasonality_PET, snow_fraction, snowfall_fraction
-Table C4: CAMELS Streamflow HBV Model Hyperparameters
+Table 17: CAMELS streamflow HBV model hyperparameters.
 Parameter Value
@@
 Type RmseLoss
-Table C5: CAMELS Streamflow Variables and Sources
-Variable Type Variable Name Source
-Time Series Forcings Precipitation, Temperature, Potential evapotranspiration, Solar radiation, Vapor pressure Catchment Attributes and Meteorology for Large-sample Studies (CAMELS) (Addor et al., 2017, Newman et al., 2014)
-Static Attributes Elevation, slope, catchment area, forest cover, LAI, GVF, soil depth, porosity, conductivity, sand, silt, clay fractions, carbonate fraction, permeability, aridity, snow fraction, precipitation extremes CAMELS
-Outputs Streamflow CAMELS gauge records
-Table C6: Soil Moisture Model Configuration
+Table 18: Soil moisture model configuration.
 Parameter Value
 General Settings
-Mode train_test
+Mode traintest
 Random seed 111111
-Data loader onlylstm_loader
-Data sampler finetuning_noHBV
+Data loader onlylstmloader
+Data sampler finetuningnoHBV
 Training Configuration
@@
 Type RmseLoss
-Table C7: Soil Moisture Variables and Sources
+C.2 Variables and Data Sources
+Table 19: StefaLand pretraining variables and sources.
 Variable Type Variable Name Source
-Time Series Forcings Albedo (BSA, WSA) Moderate Resolution Imaging Spectroradiometer (MODIS) MCD43A3 version 6 (Schaaf, Crystal and Wang, Zhuosen, 2021)
-LST (Day, Night) MODIS Land Surface Temperature/Emissivity Daily (MYD11A1) Version 6.1 (Wan et al., 2021)
-Precipitation Global Precipitation Measurement (GPM), & MSWEP &ERA5 precipitation (Huffman et al., 2019, Beck et al., 2019, Muñoz Sabater, 2019)
-Forecast albedo, LAI (high/low vegetation), soil temperature (layer 1), surface pressure, solar radiation, 2 m temperature, evaporation, precipitation, U/V wind (10 m) ECMWF Reanalysis v5 (ERA5) (Muñoz Sabater, 2019)
-Static Attributes elevation, slope, aspect, roughness, curvature Global 1/5/10/100-km topography derivatives (Amatulli et al., 2018)
-Sand, clay, silt, bulk density HWSD v1.2 (FAO et al., 2012)
-Land cover; urban; open water; snow/ice ESA CCI Land Cover (ESA, 2017)
-NDVI Vegetation Indices Monthly L3 Global 0.05Deg CMG (Didan et al., 2015)
-Outputs Soil moisture International Soil Moisture Network (ISMN) (Dorigo et al., 2013b; 2011)
-Table C8: Global Streamflow Variables and Sources
-Variable Type Variable Name Source
-Time Series Forcings Precipitation, Temperature, Potential evapotranspiration, Radiation, Humidity MSWX and MSWEP (Beck et al., 2022; 2019)
+Time Series Forcings Precipitation, Short-wave solar radiation downwards, Relative humidity, Maximum temperature, Minimum temperature, Potential evapotranspiration from Multi-Source Weather (MSWX) and Multi-Source Weighted-Ensemble Precipitation (MSWEP) (Beck et al., 2022, 2019)
 Static Attributes Forest cover fraction, grassland cover fraction Climate Change Initiative (CCI) land cover dataset (ESA, 2017)
@@
 Sand, silt, clay fractions Harmonized World Soil Database (HWSD) (FAO et al., 2012)
-Elevation, slope, aspect Global Multi-resolution Terrain Elevation Data (GMTED) (Danielson and Gesch, 2011, Ramcharan et al., 2018)
+Elevation, slope, aspect Global Multi-resolution Terrain Elevation Data (GMTED) (Danielson and Gesch, 2011; Ramcharan et al., 2018)
 Soil depth Global 1-km Gridded Thickness of Soil, Regolith, and Sedimentary Deposit Layers (Pelletier et al., 2016)
@@
 Forest intact fraction Intact Forest Landscapes Data (Potapov et al., 2017)
-Outputs Streamflow Global runoff data (GRDC, 2024)
-Table C9: Landslide (SLIDO, Oregon) Variables and Sources
+Outputs None (self-supervised pretraining)
+Table 20: CAMELS streamflow variables and sources.
+Variable Type Variable Name Source
+Time Series Forcings Precipitation, Temperature, Potential evapotranspiration, Solar radiation, Vapor pressure Catchment Attributes and Meteorology for Large-sample Studies (CAMELS) (Addor et al., 2017; Newman et al., 2014)
+Static Attributes Elevation, slope, catchment area, forest cover, LAI, GVF, soil depth, porosity, conductivity, sand, silt, clay fractions, carbonate fraction, permeability, aridity, snow fraction, precipitation extremes CAMELS
+Outputs Streamflow CAMELS gauge records
+Table 21: Soil moisture variables and sources.
+Variable Type Variable Name Source
+Time Series Forcings Albedo (BSA, WSA) Moderate Resolution Imaging Spectroradiometer (MODIS) MCD43A3 version 6 (Schaaf, Crystal and Wang, Zhuosen, 2021)
+LST (Day, Night) MODIS Land Surface Temperature/Emissivity Daily (MYD11A1) Version 6.1 (Wan et al., 2021)
+Precipitation Global Precipitation Measurement (GPM), MSWEP, and ERA5 precipitation (Huffman et al., 2019; Beck et al., 2019; Muñoz Sabater, 2019)
+Forecast albedo, LAI (high/low vegetation), soil temperature (layer 1), surface pressure, solar radiation, 2 m temperature, evaporation, precipitation, U/V wind (10 m) ECMWF Reanalysis v5 (ERA5) (Muñoz Sabater, 2019)
+Static Attributes elevation, slope, aspect, roughness, curvature Global 1/5/10/100-km topography derivatives (Amatulli et al., 2018)
+Sand, clay, silt, bulk density HWSD v1.2 (FAO et al., 2012)
+Land cover; urban; open water; snow/ice ESA CCI Land Cover (ESA, 2017)
+NDVI Vegetation Indices Monthly L3 Global 0.05Deg CMG (Didan et al., 2015)
+Outputs Soil moisture International Soil Moisture Network (ISMN) (Dorigo et al., 2013b, 2011)
+Table 22: Streamflow input variables and attributes used from the Caravan dataset.
+Variable Type Variable Name Source
+Time Series Forcings Precipitation (P), Air temperature (Ta), Potential evapotranspiration (PET) ERA5-Land via Caravan (Kratzert et al., 2023; Muñoz-Sabater et al., 2021)
+Surface pressure, 10 m wind components ($u$, $v$) ERA5-Land via Caravan (Kratzert et al., 2023; Muñoz-Sabater et al., 2021)
+Net solar radiation, net thermal radiation ERA5-Land via Caravan (Kratzert et al., 2023; Muñoz-Sabater et al., 2021)
+Snow water equivalent, soil moisture (4 layers) ERA5-Land via Caravan (Kratzert et al., 2023; Muñoz-Sabater et al., 2021)
+Static Basin Attributes Latitude, longitude, catchment area Caravan metadata (Kratzert et al., 2023)
+Aridity indices (ERA5-Land, FAO Penman–Monteith) Caravan derived attributes (Kratzert et al., 2023)
+Mean precipitation, precipitation seasonality Caravan derived attributes (Kratzert et al., 2023)
+Mean air temperature, PET seasonality Caravan derived attributes (Kratzert et al., 2023)
+Elevation, slope Global terrain products via Caravan (Kratzert et al., 2023)
+Soil texture fractions (sand, silt, clay) Harmonized World Soil Database (HWSD) via Caravan (FAO et al., 2012; Kratzert et al., 2023)
+Soil erosion index Global soil datasets via Caravan (Kratzert et al., 2023)
+Forest cover fraction Global land cover products via Caravan (Kratzert et al., 2023)
+Outputs Streamflow Gauge observations compiled in Caravan (Kratzert et al., 2023)
+Table 23: Landslide (SLIDO, Oregon) variables and sources.
 Variable Type Variable Name Source
@@
 Soil sand, silt, clay, bulk density, saturated hydraulic conductivity Probabilistic Remapping of SSURGO (POLARIS) (Chaney et al., 2019)
-Lithology Global Lithological Map (GLiM) (har)
+Lithology Global Lithological Map (GLiM) (Hartmann and Moosdorf, 2012)
 Rainfall PRISM (PRISM Climate Group, 2014)
@@
 Outputs Landslide occurrence (binary) Statewide Landslide Information Database for Oregon (SLIDO) (Franczyk, J. J et al., 2020)
-Table C10: Soil Composition (ISRIC) Variables and Sources
+Table 24: Soil composition (ISRIC) variables and sources.
 Variable Type Variable Name Source
@@
 Outputs Soil property (clay; sand; silt) World Soil Information Service (WoSIS) (Batjes et al., 2020)
-Table C11: PrithviWxC Surface Variables and Sources (MERRA-2)
-Variable Type Variable Name Source
-Time Series Forcings Precipitation, fluxes, winds, soil moisture/temperature, LAI, runoff, etc. NASA MERRA-2 reanalysis (Gelaro et al., 2017)
-Static Attributes Land/ocean/ice fractions, surface geopotential, subgrid orography NASA MERRA-2 constants (Gelaro et al., 2017)
-Table C12: Computation Resources for StefaLand and Comparison Models
+C.3 Computational Resources
+Table 25: Computation Resources for StefaLand and Comparison Models
 Model Seconds/Epoch #GPUs GPU Type Memory
 StefaLand (Pretraining) 16,000 6 NVIDIA V100 240 GB
-StefaLand with resConn 30 2 NVIDIA V100 80 GB
-StefaLand without resConn 26 2 NVIDIA V100 80 GB
+StefaLand ResConn 30 2 NVIDIA V100 80 GB
+StefaLand no Adapter 26 2 NVIDIA V100 80 GB
 LSTM Baseline 12 2 NVIDIA V100 80 GB
@@
 StefaLand-resConn HBV1.1 320 2 NVIDIA V100 80 GB
-StefaLand-no resConn HBV1.1 300 2 NVIDIA V100 80 GB
-  • Note: All values except pretraining are for the CAMELS benchmark experiment. The relative differences in computational requirements are consistent across other experiments.
-C.1 Pretraining Data Handling
-To handle large volumes efficiently, data were stored in shards rather than fully loaded into memory.
-At the start of each epoch, shards were randomly reindexed, providing diverse samples while avoiding costly materialization of the full dataset.
-This strategy ensures both reproducibility and wide sample coverage.
-The accompanying code release in the reproducibility statement enables reconstruction of this dataset and reproduces the full pretraining pipeline.
-C.2 Dataset Splitting
+StefaLand-no Adapter HBV1.1 300 2 NVIDIA V100 80 GB
+  • Note: All values except pretraining are for the CAMELS benchmark experiment. Relative computational differences are consistent across other experiments.
+C.4 Pretraining Data Handling
+Figure 4: Spatial distribution of the global streamflow dataset.
+Basins are categorized according to the availability of runoff observations: basins with relatively abundant runoff records (blue), basins with sparse runoff records (orange), and basins without runoff data (green).
+Marker size corresponds to basin area, classified into three categories based on the 33rd and 67th percentiles of catchment areas.
+A global dataset was constructed for model pretraining, including 8,634 catchments and designed to characterize climatic, ecological, soil, topographic, geological, and socioeconomic conditions.
+The dataset includes both daily meteorological forcings and long-term averaged static attributes.
+Daily meteorological variables comprise precipitation, downward shortwave radiation, relative humidity, maximum temperature, and minimum temperature, derived from the Multi-Source Weather (MSWX) and Multi-Source Weighted-Ensemble Precipitation (MSWEP) datasets at a spatial resolution of 0.1° (Beck et al., 2022, 2019).
+Potential evapotranspiration was estimated using the Hargreaves method.
+The study catchments are divided into three groups: 3,434 GRDC catchments with relatively abundant historical runoff records, 3,248 GRDC catchments with sparse runoff records, and 1,952 HydroBasins level-8 catchments without runoff observations (GRDC, 2024; Lehner and Grill, 2013).
+Ecosystem states are represented by forest and grassland cover fractions derived from the Climate Change Initiative (CCI) land cover dataset (ESA, 2017), along with the Normalized Difference Vegetation Index (NDVI) from MODIS (Didan, 2015a).
+Soil properties include sand, silt, and clay fractions from the Harmonized World Soil Database (HWSD) (FAO et al., 2012).
+Topographic attributes include elevation, slope, and aspect obtained from the Global Multi-resolution Terrain Elevation Data (GMTED) (Danielson and Gesch, 2011; Ramcharan et al., 2018), as well as terrain-derived soil depth from the Global 1-km Gridded Thickness of Soil, Regolith, and Sedimentary Deposit Layers dataset (Pelletier et al., 2016).
+Geological attributes comprise carbonate sedimentary rock fractions from the Global Lithological Map (GLiM) (Hartmann and Moosdorf, 2012) and rock porosity and permeability from the GLobal HYdrogeology MaPS (GLHYMPS) dataset (Gleeson et al., 2014).
+Socioeconomic conditions are characterized using population density from the Gridded Population of the World (GPW) v4 dataset (CIESIN, 2016), gross domestic product and population data from the gridded global GDP and Human Development Index datasets (Kummu et al., 2018), and forest intactness from the Intact Forest Landscapes dataset (Potapov et al., 2017).
+All static attributes were mapped to a common 0.01° grid prior to basin-scale aggregation to ensure spatial consistency across data sources and improve spatial averaging over irregular basin geometries.
+A complete list of variables is provided in Table 19.
+C.5 Dataset Splitting
 For the WoSIS soil dataset, we collected soil property data from 106,503 locations.
@@
 For the landslide dataset at 30 m resolution, we used 14,604 historical landslide points. We split the dataset into 70% for training, 20% for validation, and 10% for testing.
-Appendix D Model Architectures
-Figure D1: StefaLand model architecture with residual connections.
-The pretrained StefaLand encoder (frozen during fine-tuning) processes static attributes and generates embeddings that are combined with meteorological forcings through residual connections.
-This architecture enables iterative integration of transformer features with input data through the CNN module, allowing the model to effectively leverage pretrained representations while adapting to task-specific requirements.
-Figure D2: StefaLand model architecture without residual connections.
-In this configuration, the frozen transformer embeddings are used only once through a standard adapter that combines them with forcings and static features.
-This represents a more conventional fine-tuning approach where transformer features are not integrated iteratively throughout the decoding process.
-Figure D3: Different adapter architectures tested in our experiments.
-Left: Gated Adapter with a selector mechanism that controls information flow.
-Center: BottleNeck Adapter with compression and expansion phases.
-Right: Residual Adapter that adds transformer features through a skip connection.
+Appendix D Additional Figures
+Figure 5: Fine-tuning pipeline used in our downstream experiments.
+Static attributes and meteorological forcings are encoded by the pretrained StefaLand encoder (frozen), then passed through a task adapter and sequence model (LSTM), followed by projection layers to generate predictions optimized with a task loss against ground truth.
+Figure 6: Adapter architectures evaluated in our experiments.
+We compare a gated adapter, a bottleneck adapter with compression and expansion stages, A mixture of Experts and a basic feedforeward and a residual adapter that injects pretrained features via a skip connection.
 Appendix E Metric Calculations